### v1

# Learning Minimal NAP Specifications for Neural Network Verification

Abstract.
Specifications play a crucial role in neural network verification. They define the precise input regions we aim to verify, typically represented as L-infinity norm balls. While recent research suggests using neural activation patterns (NAPs) as specifications for verifying unseen test set data, it focuses on computing the most refined NAPs, often limited to very small regions in the input space. In this paper, we study the following problem: Given a neural network, find a minimal (coarsest) NAP that is sufficient for formal verification of the network’s robustness. Finding the minimal NAP specification not only expands verifiable bounds but also provides insights into which neurons contribute to the model’s robustness. To address this problem, we propose several exact and approximate approaches. Our exact approaches leverage the verification tool to find minimal NAP specifications in either a deterministic or statistical manner. Whereas the approximate methods efficiently estimate minimal NAPs using adversarial examples and local gradients, without making calls to the verification tool. This allows us to inspect potential causal links between neurons and the robustness of state-of-the-art neural networks, a task for which existing verification frameworks fail to scale. Our experimental results suggest that minimal NAP specifications require much smaller fractions of neurons compared to the most refined NAP specifications, yet they can significantly expand the verifiable boundaries to several orders of magnitude larger.
Do, Not, Us, This, Code, Put, the, Correct, Terms, for, Your, Paper
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
The growing prevalence of deep learning systems in decision-critical applications has elevated safety concerns regarding AI systems, such as their vulnerability to adversarial attacks  (Goodfellow et al., 2015; Dietterich and Horvitz, 2015). Therefore, the verification of AI systems has become increasingly important and attracted much attention from the research community. The field of neural network verification largely follows the paradigm of software verification – using formal methods to verify desirable properties of systems using rigorous mathematical specifications and proofs (Wing, 1990). Most existing works (Katz et al., 2017, 2019; Huang et al., 2017a, 2020; Wang et al., 2021b) utilize $L_{\infty}$ ball specifications for verifying robustness guarantees, i.e., verify if neural networks consistently predict the same label within local neighborhoods ($L_{\infty}$ balls) of reference data points.
Figure 1. The comparison between the NAP specification (depicted by the green region) and the $L_{\infty}$ ball specification (represented by the gray region) in the MNIST dataset. Image $a$ is the reference image, and image $c$ is the closest image to $a$ with an $L_{\infty}$ distance of 0.5294. However, image $c$ cannot be verified using the $L_{\infty}$ ball specification because an adversarial example $b$ can be found at an $L_{\infty}$ distance of 0.2. In contrast, the NAP specification enables verification of inputs that are semantically close despite their geometric distance being far apart, such as image $c$.
While this type of specification provides a fair and effective testbed for evaluating neural networks’ robustness, those verifiable $L_{\infty}$ balls are usually too tight, making them unlikely to cover unseen test set data (Geng et al., 2023). For instance, the $L_{\infty}$ maximum verifiable bounds in the MNIST dataset is smaller than 0.2, whereas the smallest distance among images sharing the same label is greater than 0.5. Thus, narrowing the gap between verifiable regions and real data distribution to further enable verification in real-world settings poses a challenge.
To this end,  Geng et al. (2023) proposes using neural activation patterns (NAPs) as specifications. Neural activation patterns can be viewed as an abstraction of each neuron’s state. For ReLU activation, it can be categorized into an activation state and a deactivation state. Since NAPs represent the intrinsic decision-making processes of neural networks, geologically distant data usually share the same NAPs if they belong to the same class. Compared to conventional local neighborhood specifications, NAP specifications cover much more flexible and potentially larger regions, as illustrated in Figure 1. This allows the verification of unseen real data rather than synthesized data from local neighbourhoods, paving the way for real-world application verification.
However, it is noteworthy that the current approach to computing NAPs relies on a simple statistical method that assumes each neuron contributes to certifying the robustness behaviors of neural networks. Consequently, the computed NAPs are often the most refined ones. However, this is an overly restrictive assumption, as many studies (Frankle and Carbin, 2019; Lu et al., 2019; Liang et al., 2021; Wang et al., 2021a) have revealed that a non-trivial portion of neurons in neural networks may not play a substantial role. This motivates us to address the following challenge: Given a neural network, find a minimal (coarsest) NAP that is sufficient for formal verification of the network’s robustness. This problem is important for the following reasons: i) Minimal NAP specifications correspond to potentially much larger regions in the input space compared to the finest ones, this increases the potential of verification of unseen data; ii) Minimal NAP specifications provide insight into which neurons are responsible for the model making robust predictions, helping us to reveal the black-box nature of neural networks.
To find the minimal NAP specifications, we introduce two basic algorithms: Refine and Coarsen. These algorithms exhaustively check all possible candidates using off-the-shelf verification tools, such as Marabou (Katz et al., 2019). While these approaches provide correctness guarantees, they are not efficient for verifying large neural networks as calls to the verification tool are usually very expensive. To tackle this inefficiency, we propose statistical versions of Refine and CoarsenSample_Refine and Sample_Coarsen that leverage sampling and statistical learning principles to find the minimal NAP specification in a more efficient manner. However, current verification tools cannot scale up to state-of-the-art neural network models, limiting the ability of verification-dependent approaches to compute the minimal NAP specification.
This motivates us to study estimation methods that do not rely on verification tools. To this end, we find that adversarial examples and local gradients can provide valuable information about mandatory neurons, which are the essential building blocks of minimal NAPs. We then translate these findings into two approximate approaches: Adversarial_Prune and Gradient_Search. They can be used together to estimate mandatory neurons of state-of-art neural networks such as VGG-19 (Simonyan and Zisserman, 2014) and inspect possible links between models’ robustness performance and internal neurons, a task that existing verification frameworks struggle to scale for.
Moreover, previous research has claimed that NAP specifications cover significantly larger regions compared to $L_{\infty}$ ball specifications, yet this assertion lacks empirical evidence. In this paper, we address this gap by introducing a simple approach to estimate the volume of regions corresponding to NAP specifications. Furthermore, this method provides us with a rough approximation of the volumetric change from refined NAP specifications to minimal NAP specifications. Our experimental results demonstrate significant volumetric changes, thereby underscoring the efficacy of minimal NAP specifications in real-world verification scenarios. We summarize our contributions as follows:
  • We propose the problem of learning the minimal NAP specification and illustrate its importance in neural network verification. We present two simple approaches that leverage the verification tool with correctness guarantees.
  • We introduce several key concepts related to the problem of learning minimal NAP specifications, including the abstraction function, NAP specification, and mandatory neurons. We demonstrate that the problem can be easily solved with access to mandatory neurons and propose two efficient estimation approaches to learn them.
  • We propose two statistical learning approaches to the minimal NAP specification problem. They leverage sampling and estimation techniques as heuristics to achieve better efficiency.
  • We present a simple approach for estimating the volume of regions corresponding to NAPs. This helps us understand the volumetric change in verifiable space between the minimal NAP specifications and the refined NAP specifications.
  • Our experiments indicate that the minimal NAP specifications involve much smaller fractions of neurons compared to the most refined NAPs. And they significantly expand the verifiable bound by several orders of magnitude.
2. Background
In this section, we introduce basic knowledge and notations of adversarial attacks and neural network verification, with an emphasis on verification using NAP specifications. This may help readers better understand the importance of learning minimal NAP specifications.
2.1. Neural Networks for Classification Tasks
In this paper, we focus on feed-forward ReLU neural networks. Generally speaking, a feed-forward network $N$ is comprised of $L$ layers, where each layer performs a linear transformation followed by a ReLU activation. We denote the pre-activation value and post-activation value at the $l$-th layer as $z^{(l)}(x)$ and $\hat{z}^{(l)}(x)$, respectively. The $l$-th layer computation is expressed as follows: $z^{(l)}(x)=\mathbf{W}^{(l)}\hat{z}^{(l-1)}(x)+\mathbf{b}^{(l)}$, $\hat{z}^{(l)}(x)=\mathbf{ReLu}(z^{(l)}(x))$, with $\mathbf{W}^{(l)}$ being the weight matrix and $\mathbf{b}^{(l)}$ representing the bias for the $l$-th layer. We denote the number of neurons in the $l$-th layer as $d_{l}$, and the $i$-th neuron in layer $l$ as $N_{i,l}$. The pre-activation value and post-activation value of $N_{i,l}$ at input $x$ is computed by $z_{i}^{(l)}(x)$ and $\hat{z}_{i}^{(l)}(x)$. The network can also be viewed as a function $\mathbf{F}^{<N>}:=\mathbb{R}^{d_{0}}\rightarrow\mathbb{R}^{d_{L}}$ such that $\mathbf{F}^{<N>}(x):=z^{(L)}(x)$, where $\mathbf{F}^{<N>}_{i}(x):=z_{i}^{(L)}(x)$ represents the output of $i$-th neuron in the last layer. We will omit $N$ when the context is clear. In a multi-class classification task with $C$ as the set of classes, when provided with an input $x$, the neural network predicts class/label $c\in C$ if the output of the $c$-th neuron $\textbf{F}_{c}(x)$ is the highest among the neurons in the $L$-th layer.
2.2. Adversarial Attacks and Neural Networks Verification Problem
Given a neural network $\mathbf{F}$ and a reference point $x$, adversarial attacks aim to search for a point $x^{\prime}$ that is geometrically close to the reference point $x$ such that $x^{\prime}$ and $x$ belong to different classes. Here, we use the canonical specification, that is, we want to search in the local neighborhoods ($L_{\infty}$ norm balls) of $x$, formally denoted as $B(x,\epsilon):=\{x^{\prime}\mid||x-x^{\prime}||_{\infty}\leq\epsilon\}$, where $\epsilon$ is the radius. For certain $\epsilon$, given we know $x$ belongs to class $j$, we say an adversarial point is found if:
(1) $\exists x^{\prime}\in B(x,\epsilon)\quad\exists i\in C\quad\mathbf{F}_{i}(x^{\prime})-\mathbf{F}_{j}(x)>0$
In practice, the change from the original data $x$ to adversarial data $x^{\prime}$ should be imperceptible, so they are more likely to be recognized as the same class/label from a human perspective. There are also metrics other than the $L_{\infty}$ norm to represent the ”similarity” between $x$ and $x^{\prime}$, such as the $L_{0}$ and $L_{2}$ norms (Xu et al., 2020). However, almost all of them fall into the local neighborhood specification paradigm. This is different from the NAP specification, as we will discuss later.
Neural networks have been shown to be vulnerable to adversarial attacks. Even imperceptible changes could easily alter the predicted outcome. This highlights the importance of neural network verification. The neural network verification problem can be seen as the flip side of adversarial attacks. When it can be formally proven that no adversarial point exists in $B(x,\epsilon)$, the neural network verification problem is considered solved. Formally, we want to verify:
(2) $\forall x^{\prime}\in B(x,\epsilon)\quad\forall i\neq j\quad\mathbf{F}_{j}(x^{\prime})-\mathbf{F}_{i}(x^{\prime})>0$
For a simpler presentation, we assume that $\mathbf{F}(x)$ is a binary classifier. For any given specification, $\mathbf{F}(x)\geq 0$ indicates that the model is verified; otherwise, we can find an adversarial example. Solving such a problem is known to be NP-hard (Katz et al., 2017), and current state-of-the-art approaches rely on specialized solvers. The difficulty escalates significantly as the number of neurons in neural networks increases, limiting successful verifications to only very small neural networks. Therefore, achieving scalability in verification remains an ongoing challenge.
2.3. Neuron Abstractions and Neural Activation Patterns
To better discuss robustness verification using NAP specifications, we first introduce the relevant concepts of neuron abstractions, neuron abstraction functions, and neural activation patterns.
Neuron Abstractions
[…]
Definition 2.1 (Neuron Abstraction Function).
Given a neural network $N$ and the abstraction state set $\mathbb{S}$. A neuron abstraction function is the mapping $\mathcal{A}:N\rightarrow\mathbb{S}$. Formally, for an arbitrary neuron $N_{i,l}$ where $0\leq i\leq d_{l}$ and $11\leq l\leq L-$ 1, the function abstracts $N_{i,l}$ to some state $s_{k}\in\mathbb{S}$, i.e., $\mathcal{A}(N_{i,l})=s_{k}$.
The above characterization of neuron abstraction , it does not instruct us on how to perform binary abstraction in the absence of neuron values. Thus, we include the input(s) $x$ or $X$ as a new parameter in $\mathcal{A}$, and it is omitted when the context is clear. Here, we discuss three types of $\mathcal{A}$: the unary abstraction function function $\dot{\mathcal{A}}$, the binary abstraction function $\ddot{\mathcal{A}}$, and the statistical abstraction function $\widetilde{\mathcal{A}}$.
The unary $\dot{\mathcal{A}}$ function
[…]
(3) $\dot{\mathcal{A}}(N_{i,l},x)=\mathbf{*}$
The binary $\ddot{\mathcal{A}}$ function
When an input $x$ is passed through an internal neuron $N_{i,l}$, we can easily determine the binary abstraction state of $N_{i,l}$ based on its post-activation value $\hat{z}_{i}^{(l)}(x)$. This motivates us to define the binary abstraction function as follows:
(4) $\ddot{\mathcal{A}}(N_{i,l},x)=\begin{cases}\mathbf{0}&\text{if }\hat{z}_{i}^{(l)}(x)=0\\ \mathbf{1}&\text{if }\hat{z}_{i}^{(l)}(x)>0\\ \end{cases}$
The statistical $\widetilde{\mathcal{A}}$ function
The binary $\ddot{\mathcal{A}}$ function is limited in calculating a single input, and meets a challenge when multiple inputs come into play, as two distinct inputs may lead to disagreements in a neuron’s abstraction state. We then approach this problem statistically by introducing the $\widetilde{\mathcal{A}}$ function defined as follows:
(5) $\widetilde{\mathcal{A}}(N_{i,l},X)=\begin{cases}\mathbf{0}&\text{if }\frac{|\{x_{j}|\ddot{\mathcal{A}}(N_{i,l},x_{j})=\mathbf{0},\text{ }x_{j}\in X\}|}{|X|}\geq\delta\\ \mathbf{1}&\text{if }\frac{|\{x_{j}|\ddot{\mathcal{A}}(N_{i,l},x_{j})=\mathbf{1},\text{ }x_{j}\in X\}|}{|X|}\geq\delta\\ \mathbf{*}&\text{otherwise}\\ \end{cases}$
where $\delta$ is a real number from $[0,1]$, and $X$ represents a set of inputs, i.e., $X:=\{x_{1},x_{2},...,x_{n}\}$. The parameter $\delta$ represents an error ratio to accommodate possible disagreements on neuron states between multiple inputs. Intuitively, the introduction of $\delta$ allows multiple inputs $x_{1},...,x_{n}$ to vote on a neuron’s state. For instance, when $\delta$ is set to 0.99, then 99% or more of the inputs must agree that a neuron is activated for the neuron to be in the $\mathbf{1}$ state.
(a) A simple 2x2 fully connected neuron network.
(b) A subset of the NAP $\mathcal{P}$ family forms a binary tree. Each child subsumes its parent, and the root is the most abstracted NAP. The leaf node represents the most refined NAPs, where each neuron is either $\mathbf{1}$ or $\mathbf{0}$. From the root, the tree is spanned by abstracting each neuron in an order of $N_{0,1},N_{1,1},N_{0,2},N_{1,2}$. Setting a different order will create a different tree of NAPs.
Figure 2. A simple 2x2 fully connected neural network and a subset of its NAP in a binary tree structure.
Definition 2.2 (Neural Activation Pattern).
Given a neural network $N$ and a neuron abstraction function $\mathcal{A}$. A neural activation pattern (NAP) $P$ is a tuple that consists of the abstraction state of all neurons in $N$. Formally, $P:=\langle\mathcal{A}(N_{i,l})\text{ }|\text{ }N_{i,l}\in N,\mathcal{A}\in\{\ddot{\mathcal{A}},\dot{\mathcal{A}}\}\rangle$, also denoted as $\mathcal{A}(N)$. The neuron $N_{i,l}$’s abstraction state specified by NAP $P$ is represented as $P_{i,l}$, i.e., $P_{i,l}=\mathcal{A}(N_{i,l})$.
Partially ordered NAP
We denote the power set of NAPs in $N$ as $\mathcal{P}$. The number of all possible NAPs in $N$ scales exponentially as the number of neurons increases. For such a large set, if we aim to find the minimal NAP — the central problem in this work, we first have to establish an order so that NAPs can be compared. To this end, we define the following partial order:
Definition 2.3 (Partially ordered NAP).
For any given two NAPs $P,P^{\prime}\in\mathcal{P}$, we say $P^{\prime}$ subsumes $P$ if, for each neuron $N_{i,l}$, its state in $P$ is an abstraction of that in $P^{\prime}$. Formally, this can be defined as:
(6) $P^{\prime}\preccurlyeq P\text{ }\iff\text{ }P^{\prime}_{i,l}\preceq P_{i,l}\text{ }\text{ }\forall N_{i,l}\in N$
Moreover, two NAPs $P,P^{\prime}$ are equivalent if $P\preccurlyeq P^{\prime}$ and $P^{\prime}\preccurlyeq P$.
To give a concrete example, Figure 1(b) depicts a subset of NAPs of a simple neural network consisting of 2 hidden layers and 4 neurons, as presented in Figure 1(a).
[…]
precedes-or-equals
[…]
1011
[…]
. This is because these NAPs sit on different branches formed by splitting certain ReLUs.
Region Outlined by NAPs
One key requirement for verification specification is its ability to represent a specific region within the input space.
[…]
Figure 3. Linear regions are shattered by the simple 2X2 neural network. Each linear region corresponds to a most refined NAP, but not necessarily vice versa. More abstracted NAPs are formed by ignoring lines/hyperplanes created by neurons.
2.4. NAP Specifications for Robustness Verification
This subsection illustrates how NAP specifications can be utilized for robustness verification. We discuss how the abstraction states of neurons can serve as signatures of distinct classes with the introduction of the class NAP definition.
Definition 2.4 (Class NAP).
In a classification task with the class set $C$, for any class $c\in C$, a class NAP $P^{c}$ is a NAP comprising abstract states outputted by an abstraction function given $X_{c}$, where $X_{c}$ denotes the set of inputs belonging to class $c$. Formally, $P^{c}:=\langle\mathcal{A}(N_{i,l},X_{c})\mid N_{i,l}\in N,\mathcal{A}\in\{\widetilde{\mathcal{A}},\dot{\mathcal{A}}\}\rangle$. The power set of $P^{c}$ is denoted as $\mathcal{P}^{c}$.
Recall that robustness verification can be thought of as proving that no adversarial examples exist in the local neighbourhoods of some reference point $x$. Intuitively, for NAP specifications, it is equivalent to show that no adversarial examples exist in the class NAP $P^{c}$; in other words, inputs exhibiting the NAP $P^{c}$ must be predicted as class $c$. Geng et al. (2023) argues that class NAPs must satisfy several essential requirements to qualify as NAP specifications. And finding qualified NAP specifications implies the underlying robustness problem is verified. We formally frame these requirements into the following properties:
The non-ambiguity property
Since we want class NAPs to serve as certificates for a certain class, they must be distinct from each other. Otherwise, there could exist an input that exhibits two class NAPs, which leads to conflicting predictions. Formally, we aim to verify the following:
[…]
Pc2
[…]
From a geological perspective, there must be no overlaps between class NAPs. In other words, it is also equivalent to verifying:
$\forall c_{1},c_{2}\in C\text{ s.t. }c_{1}\neq c_{2}\quad R_{P^{c_{1}}}\bigcap R_{P^{c_{2}}}=\emptyset$
We empirically observe that successful models, particularly neural networks that have achieved high accuracy in classification tasks, naturally exhibit this property. Because successful models tend to avoid confusion in predictions, their class NAPs are usually very distinct from each other.
The NAP robustness property
To serve as NAP specifications, class NAPs $P^{c}$ ensure that if an input exhibits it, i.e., $\mathcal{A}(N,x)\preccurlyeq P,\mathcal{A}\in\{\ddot{\mathcal{A}},\dot{\mathcal{A}}\}$, the input must be predicted as the corresponding class $c$. Formally, we have:
(10) $\forall x\in R_{P^{c}}\quad\forall k\in C\text{ s.t. }k\neq c\quad\mathbf{F}_{c}(x)-\mathbf{F}_{k}(x)>0$
in which
(11) $R_{P^{c}}=\{x\mid{\mathcal{A}(N,x)\preccurlyeq P^{c}}\}$
In contrast to canonical $L_{\infty}$ norm balls, class NAPs are more flexible in terms of size and shape. Additionally, there is no need to specify a reference point, since the locations of potential reference points are also encoded by class NAPs. However, it is possible that no class NAP $P^{c}$ in $\mathcal{P}^{c}$ satisfies this property. This can be mitigated by meeting the subsequent weaker property.
The NAP-augmented robustness property
Instead of relying solely on class NAPs as specifications, local neighbours can still be employed in conjunction for verification. This hybrid form of specification has several advantages: 1) It narrows down the scope of verifiable regions when no class NAPs can meet the NAP robustness property; 2) NAP constraints essentially fix ReLU states, refining the search space for verification tools; 3) It focuses on the verification of valid test inputs rather than adversarial examples. Formally, this property can be stated as:
(12) $\forall x^{\prime}\in B(x,\epsilon)\bigcap R_{P^{c}}\quad\forall k\in C\text{ s.t. }k\neq c\quad\mathbf{F}_{c}(x)-\mathbf{F}_{k}(x)>0$
in which
(13) $B(x,\epsilon)=\{x^{\prime}\mid||x-x^{\prime}||_{\infty}\leq\epsilon\}\quad R_{P^{c}}=\{x^{\prime}\mid{\mathcal{A}(N,x^{\prime})\preccurlyeq P^{c}}\}$
To summarize, we state that a class NAP can serve as a NAP specification if it satisfies either the NAP robustness property or the NAP-augmented robustness property. Clearly, the former property is stronger, and it is possible that we can’t find a class NAP $P^{c}$ in $\mathcal{P}^{c}$ to satisfy this property. Fortunately, we can always find NAPs that satisfy the latter property by narrowing the verifiable region using additional $L_{\infty}$ norm ball specifications. Moreover, NAP specifications can be easily encoded with most of the state-of-the-art verification tools such as Marabou. In this paper, we will not dive into NAP encodings. Interested readers can refer to Geng et al. (2023) for more details.
3. The minimal NAP specification problem
In this section, we formulate the problem of learning minimal NAP specifications and present two naive approaches for solving this problem. Since our approaches require interactions with verification tools, we begin by introducing relevant notations describing the relationships between NAPs and verification tools.
Figure 4. A DAG representing all possible NAPs, with green nodes signifying they can serve as a specification, while red nodes indicate they cannot. Within each node, a black circle means the corresponding neuron is abstracted by $\widetilde{\mathcal{A}}$, whereas the white circle means it is abstracted by $\dot{\mathcal{A}}$. Thus, the top node represents the coarsest NAP, while the bottom node entails the most refined one. Our goal is to find the NAP that resides in the green nodes while its size is minimal.
3.1. Problem Formulation
Let $(\mathcal{P}^{c},\preccurlyeq)$ be a partially ordered set corresponding to a family of class NAPs regarding some class $c\in C$. For simplicity, we omit the superscript $c$ and refer to class NAPs simply as NAPs when the context is clear. We assume we have access to a verification tool, $\mathcal{V}:\mathcal{P}\to\{0,1\}$, which maps a class NAP $P\in\mathcal{P}$ to a binary set. Here, $\mathcal{V}(P)=1$ denotes a successful verification of the underlying robustness query, while 0 indicates the presence of an adversarial example. From an alternative perspective, $\mathcal{V}(P)=1$ also signifies that $P$ is a NAP specification, i.e., it satisfies NAP(-augmented) robustness properties; whereas $\mathcal{V}(P)=0$ implies the opposite.
It is not hard to see that $\mathcal{V}$ is monotone with respect to the NAP family $(\mathcal{P},\preccurlyeq)$. Given $P\preccurlyeq P^{\prime}$ and $\mathcal{V}(P^{\prime})=1$, it follows that $\mathcal{V}(P)=1$. However, given $P\preccurlyeq P^{\prime}$ and $\mathcal{V}(P)=1$, we cannot determine $\mathcal{V}(P^{\prime})$. In other words, refining a NAP (by increasing the number of neurons abstracted to $\mathbf{0}$ or $\mathbf{1}$) can only enhance the likelihood of successful verification of the underlying robustness query.
Definition 3.1 (The minimal NAP specification problem).
Given a family of NAPs $\mathcal{P}$ and a verification tool $\mathcal{V}$, the minimal NAP specification problem is to find a NAP $P$ such that
$\underset{P\in\mathcal{P},\mathcal{V}(P)=1}{\arg\min}|P|$
where $|P|$ is defined as the size of $P$, representing the number of neurons that are abstracted to the $\mathbf{0}$ or $\mathbf{1}$, i.e., |$|\{N_{i,l}\mid P_{i,l}=\mathbf{0}\text{ or }\mathbf{1}\}|$. The (largest) minimal NAP specification’s size is denoted as $s$.
When $P$ is minimal, it implies that for any NAP $P^{\prime}$ that is strictly more coarse than $P$, $\mathcal{V}(P^{\prime})=0$. Formally, this can be expressed as: $\forall P\preccurlyeq P^{\prime},P^{\prime}\neq P:\mathcal{V}(P^{\prime})=0$. Thus, there could exist multiple minimal NAP specifications. In such cases, we only need to choose one of them. On the other hand, it is possible that even the most refined NAPs cannot verify the robustness query. In such cases, we claim that no minimal NAP specifications exist. Additionally, since the computation using verification tools is usually expensive, we are interested in methods that efficiently find a minimal NAP specification, i.e., minimizing the number of calls to $\mathcal{V}$.
We present two naive approaches to solving the problem. Since these approaches involve refining and coarsening the NAP, we will first formally define these two actions.
Refine and Coarsen Action of NAP
Recall our definition of NAP, the coarsest NAP is the one using $\dot{\mathcal{A}}$ to abstract each neuron in $N$. We denote this NAP as $\dot{P}:=\langle\dot{\mathcal{A}}(N_{i,l})\text{ }|\text{ }N_{i,l}\in N\rangle$. $\dot{P}$ is the smallest NAP with $|\dot{P}|=0$. In addition, we define the most refined NAP as the one that applies $\widetilde{\mathcal{A}}$ to abstract each neuron in $N$. We denote it as $\widetilde{P}:=\langle\widetilde{\mathcal{A}}(N_{i,l})\text{ }|\text{ }N_{i,l}\in N\rangle$. $\widetilde{P}$ is the largest NAP with size with $|\widetilde{P}|\leq|N|$. Clearly, $\widetilde{P}\preccurlyeq\dot{P}$. Given any NAP $P$, if we want to refine $P$ through a specific neuron $N_{i,l}$, we apply the $\widetilde{\mathcal{A}}$ function to $N_{i,l}$. We denote this refinement action as $\widetilde{\Delta}(N_{i,l})$. This action will either increase or leave unchanged $|P|$. Similarly, we denote the coarsen action as $\dot{\Delta}(N_{i,l})$. This action will either decrease or leave unchanged $|P|$. We present the semantics of these actions as follows:
               $P\vdash\cdot\Downarrow P$
  $P\vdash\Delta_{1}\Downarrow P^{\prime}$         $P^{\prime}\vdash\Delta_{2}\Downarrow P^{\prime\prime}$              $P\vdash\Delta_{1};\Delta_{2}\Downarrow P^{\prime\prime}$
                           $P_{i,l}=\mathbf{*}$ Refine              $P\vdash\widetilde{\Delta}(N_{i,l})\Downarrow\widetilde{\mathcal{A}}(N_{i,l})\cup P\setminus P_{i,l}$                                 $P_{i,l}\in\{\mathbf{0},\mathbf{1}\}$ Coarsen                $P\vdash\dot{\Delta}(N_{i,l})\Downarrow\dot{\mathcal{A}}(N_{i,l})\cup P\setminus P_{i,l}$
where $P,P^{\prime},P^{\prime\prime}$ are from the NAP family $\mathcal{P}$, and $N$ is the underlying neural network. $\Delta_{1}$ and $\Delta_{2}$ are any two sequences of refine and coarsen actions.
3.2. The Refine Approach
Conceptually, the Refine approach iteratively increases the number of refined neurons in NAP $P$ until $\mathcal{V}(P)=1$, i.e., $P$ is able to prove the underlying robustness query. In other words, we gradually increase the size parameter $k$ and iterate over each NAP $P$ of size $k$ to check if $\mathcal{V}(P)=1$, as illustrated in Algorithm 1. To determine if a solution to the problem exists, we first check if the most refined NAP can succeed in verification. We proceed to iterative refinement only if $\mathcal{V}(\widetilde{P})=1$. However, the algorithm is not efficient and requires $12^{|N|}-$ 1 calls to $\mathcal{V}$ in the worst case, as proven in Theorem 3.2. Please refer to the proof in Appendix A. Therefore Refine is only practical when the search space of the NAP family $\mathcal{P}$ is small.
Theorem 3.2.
The algorithm Refine returns a minimal NAP specification with $\mathcal{O}(2^{|N|})$ calls to $\mathcal{V}$.
3.3. The Coarsen Approach
In contrast to the Refine approach, the Coarsen approach starts from the most refined NAP and then gradually coarsens it. We first check if the problem is well-defined by verifying if the most refined NAP $\widetilde{P}$ succeeds in verification. Then, for each neuron, we attempt to coarsen it using $\dot{\mathcal{A}}$; if the resulting NAP no longer verifies the query, we refine it back using $\widetilde{\mathcal{A}}$; otherwise, we keep the coarsened NAP. We describe the above procedure in Algorithm 2. The algorithm could require $|N|$ calls to $\mathcal{V}$ in the worst case, as proven in Theorem 3.3. Please refer to the proof in Appendix A.
Theorem 3.3.
The algorithm Coarsen returns a minimal NAP specification with $\mathcal{O}(|N|)$ calls to $\mathcal{V}$.
Input: The neural network $N$
Output: A minimal NAP specification $P$
1 Function Refine(N)
2       $P\leftarrow\widetilde{\mathcal{A}}(N)$
3       if $\mathcal{V}(P)==0$ then
             return $None$ ;   /* Return None if even the most refined NAP fails verification */
4
5       else
6             for $k$ from 1 to $|N|$ do
                   $P\leftarrow\dot{\mathcal{A}}(N)$   /* Refinement starts from the coarsest NAP $\dot{P}$/
                   $K\_comb\leftarrow$ $Pick\_K\_Neurons(k,N)$   /* Returns all combinations of size k */
7                   for $comb\textbf{ in }K\_comb$ do
8                         for $N_{i,l}\textbf{ in }comb$ do
                               $P\leftarrow\widetilde{\Delta}(N_{i,l})$  /* Refine neurons in the chosen combination */
9
10                        if $|\mathcal{V}(P)|==1$ then
                               return $P$   /* Found the minimal NAP */
11
12
13
14
15
Algorithm 1 refine
Input: The neural network $N$
Output: A minimal NAP specification $P$
1 Function Coarsen(N)
2       $P\leftarrow\widetilde{\mathcal{A}}(N)$
3       if $\mathcal{V}(P)==0$ then
             return $None$ ;   /* Return None if the most refined NAP fails verification */
4
5       else
6             for $N_{i,l}$ in $N$ do
                   $P\leftarrow\dot{\Delta}(N_{i,l})$ ;   /* Try to abstract this neuron using $\dot{\Delta}$/
7                   if $\mathcal{V}(P)==0$ then
                         $P\leftarrow\widetilde{\Delta}(N_{i,l})$ ;   /* Refine the neuron back if the verification fails */
8
9
10            return $P$;
11
Algorithm 2 Coarsen
4. Verification Free Approaches
While the Refine and Coarsen algorithms can find minimal NAP specifications with correctness guarantees, their inefficiency poses challenges with the verification of large neural networks. To address this issue, we introduce two efficient approaches for estimating minimal NAP specifications without requiring expensive calls to the verification tool. Unlike verification-based approaches, these estimation methods are deeply linked to mandatory neurons, a key concept in dissecting the minimal NAP specification problem, as discussed below.
Definition 4.1 (Mandatory neuron).
[…]
$M=\{N_{i,l}\mid P_{i,l}\in\{\mathbf{0},\mathbf{1}\},P\textit{ is minimal}\}$
Note that $M$ is the union of the set of mandatory neurons from all minimal NAP specifications. It follows that $|M|\geq s$, where $s$ denotes the size of the largest minimal NAP specification.
The minimal NAP specification problem can be solved trivially if we gain access to $M$. Thus, our verification-free approaches are designed to determine mandatory neurons and estimate $M$. To better understand these our approaches, we first discuss the properties of mandatory neurons. Recall that verifying a robustness query given a NAP specification $P$ is equivalent to showing that $\mathbf{F}(x)\geq 0$ for input $x$ in region $R_{P}$. Thus, the necessary conditions of a mandatory neuron $N_{i,l}$ can be written as follows:
  1. (1)
    If $N_{i,l}$ is in state $\mathbf{0}$, it implies when $\hat{z}_{i}^{(l)}(x)=0,\mathbf{F}(x)\geq 0$. In addition, $\exists x\text{ s.t. }\hat{z}_{i}^{(l)}(x)>0,\mathbf{F}(x)<0$.
  2. (2)
    If $N_{i,l}$ is in state $\mathbf{1}$, it implies $\forall x\text{ s.t. }\hat{z}_{i}^{(l)}(x)>0,\mathbf{F}(x)\geq 0$. In addition, when $\hat{z}_{i}^{(l)}(x)=0,\mathbf{F}(x)<0$.
[…]
4.1. Adversarial-guided Prune
We first introduce Adversarial_Prune to identify mandatory neurons. Intuitively, it attempts to show a neuron $N_{i,l}$ is mandatory by providing an adversarial example $x^{\prime}$ exists.
When an adversarial example $x^{\prime}$ is found, it immediately indicates that the NAP $\ddot{\mathcal{A}}(N,x^{\prime})$ fails the verification, i.e., $\mathcal{V}(\ddot{\mathcal{A}}(N,x^{\prime}))=0$. Moreover, it also implies that any NAP subsumed by $\ddot{\mathcal{A}}(N,x^{\prime})$ fails verification. For instance, suppose an adversarial example $x^{\prime}$ is found for a simple one-layer four-neuron neural network and $\ddot{\mathcal{A}}(N,x^{\prime})$ is $\langle\mathbf{1},\mathbf{0},\mathbf{1},\mathbf{0}\rangle$. We can infer that NAPs like $\langle\mathbf{1},\mathbf{0},\mathbf{1},\mathbf{*}\rangle$, $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{*}\rangle$, $\langle\mathbf{1},\mathbf{*},\mathbf{*},\mathbf{0}\rangle$, and $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle$ fail the verification. This information is particularly useful when determining if a neuron is mandatory. For example, if we know that NAP $P:=\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{1}\rangle$ is a specification, i.e., $\mathcal{V}(P)=1$, then we can easily deduce that the fourth neuron $N_{4,1}$ is mandatory. This is because that coarsening the fourth neuron would expand $P$ to $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{*}\rangle$, which would include the adversarial example $x^{\prime}$ and thus fail the verification, as illustrated in Figure 4(a). It is evident that the neuron where $P$ and $\ddot{\mathcal{A}}(N,x^{\prime})$ disagrees must be mandatory.
However, when the two NAPs disagree on multiple neurons, things become a little bit different. Suppose the NAP specification $P$ is $\langle\mathbf{1},\mathbf{1},\mathbf{*},\mathbf{1}\rangle$, i.e., $\mathcal{V}(\langle\mathbf{1},\mathbf{1},\mathbf{*},\mathbf{1}\rangle)=1$. We know $\mathcal{V}(\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle)=0$ by the adversarial example $x^{\prime}$. In this case, if we coarsen the second and fourth neurons, $N_{2,1}$ and $N_{4,1}$, the NAP specification will expand to $\langle\mathbf{1},\mathbf{*},\mathbf{*},\mathbf{*}\rangle$, which will cover the $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle$, thus failing the verification. In this case, $N_{2,1}$ and $N_{4,1}$ could both be mandatory neurons or either one of them is mandatory, as illustrated in Figures 4(b), 4(c), 4(d). So, we simply let $\{N_{2,1},N_{4,1}\}$ be the upper bound of mandatory neurons (learned from $x^{\prime}$). Formerly, given a NAP $P$, we say a neuron $N_{i,l}$ is in the upper bound of mandatory neurons $M$ if satisfies the following condition:
  1. (1)
    $N_{i,l}$ must be in the binary states, i.e., $P_{i,l}\in\{\mathbf{0},\mathbf{1}\}$
  2. (2)
    There exists $x^{\prime}$ such that $\ddot{\mathcal{A}}(N_{i,l},x^{\prime})$ XORs with $P_{i,l}$, i.e., $\exists x^{\prime}\text{ such that }\ddot{\mathcal{A}}(N_{i,l},x^{\prime})\oplus P_{i,l}=1$
(a) $N_{4,1}$ is mandatory.
(b) $N_{4,1},N_{2,1}$ are mandatory.
(c) $N_{4,1}$ is mandatory.
(d) $N_{2,1}$ is mandatory.
Figure 5. Geometric interpretation of NAPs on mandatory neurons. The first subfigure represents the case when the two NAPs disagree on $N_{4,1}$. The other three subfigures represent the three cases when the two NAPs disagree on $N_{4,1},N_{2,1}$. Regions colored green pass verification, whereas red indicates an adversarial example exists. Here, we omit states for $N_{1,1},N_{3,1}$ in those NAPs for simplicity.
The field of adversarial attacks has been extensively studied, offering a wealth of methods that we can leverage. These approaches are usually computationally efficient, making it easy to access a large collection of adversarial examples. Therefore, we can compute the upper bound of mandatory neurons efficiently by simply taking the union of upper bounds learned for each adversarial example, as illustrated in Algorithm 3.
Input: The neural network $N$, the input dataset $X$
Output: A collection of mandatory neurons
1 Function Adversarial_Prune($N$, $X$)
2       $Mand\leftarrow\emptyset$ ; $P\leftarrow\widetilde{\mathcal{A}}(N)$ for $x_{j}$ in $X$ do
3             $x^{\prime}_{j}\leftarrow Adversarial\_Attack(x_{j})$ for $N_{i,l}$ in $N$ do
4                   if $P_{i,l}\in\{\mathbf{0},\mathbf{1}\}$ and $\ddot{\mathcal{A}}(N_{i,l},x^{\prime}_{j})\oplus P_{i,l}$ then
                         $Mand\leftarrow Mand\cup\{N_{i,l}\}$   /* Create mandatory neurons for $x^{\prime}_{j}$/
5
6
7      return $Mand$
Algorithm 3 Adversarial_Prune
4.2. Gradient-guided Search
We introduce another approach called Gradient_Search to identify mandatory neurons. Similar to Adversarial_Prune, Gradient_Search avoids costly interactions with the verification tool. Instead, it leverages gradient estimations to analyze the structure of $\mathbf{F}$.
Recall our definition of a neural network in Section 2.1, where the output $\mathbf{F}(x)$ can be seen as a function of the post-activation value of any internal neuron $\hat{z}_{i}^{(l)}$. We denote this function as $\mathbf{F}(\hat{z}(x))$, omitting $i$ and $l$ for simplicity. The function $\mathbf{F}(\hat{z})$ is non-linear and operates within an input range of $\hat{z}\in[0,+\infty)$. It is worth noting that the gradient $\frac{\partial\mathbf{F}}{\partial\hat{z}}$ can provide valuable insights into the local structure of $\mathbf{F}$. This insight may help determine the necessity condition of mandatory neurons by checking if there exists a $\hat{z}$ such that $\mathbf{F}(\hat{z})<0$ for the corresponding neuron. To be more specific, given some sampled data $\hat{z}_{1},\ldots,\hat{z}_{n}$, we can compute their corresponding function values $\mathbf{F}(\hat{z}_{1}),\ldots,\mathbf{F}(\hat{z}_{n})$ along with their gradients $\frac{\partial\mathbf{F}}{\partial\hat{z}}|_{\hat{z}_{1}},\ldots,\frac{\partial\mathbf{F}}{\partial\hat{z}}|_{\hat{z}_{n}}$. Then if there exists a $\hat{z}_{j}$ such that $|\mathbf{F}(\hat{z}_{j})|$ is sufficiently small, yet the norm of its gradient $|\frac{\partial\mathbf{F}}{\partial\hat{z}}|_{\hat{z}_{j}}|$ is significantly large, then it is very likely that $\mathbf{F}$ will go below zero nearby, as illustrated in Figure 6. Then, if we happen to know that the corresponding neuron $N_{i,l}$ is in state $\mathbf{0}$, it is highly likely that $N_{i,l}$ is mandatory. Conversely, if a neuron $N_{i,l}$ is in state $\mathbf{1}$, and we sample an $x$ such that $\hat{z}_{i}^{(l)}=0$ and $\mathbf{F}(x)<0$, we immediately recognize that the neuron is mandatory. Algorithm 4 presents pseudocode for estimating whether a neuron is mandatory based on the discussions above.
Figure 6. The gradient provides insights into the local structure of $\mathbf{F}$. When the gradient is large and the function value is small, it strongly suggests that $\mathbf{F}$ will go below zero nearby.
Input: The neural network $N$, the input dataset $X$, and thresholds $\beta$, $\gamma$
Output: A collection of mandatory neurons
1 Function Gradient_Search($N$, $X$, $\beta$, $\gamma$)
2       $Mand\leftarrow\emptyset$ ; $P\leftarrow\widetilde{\mathcal{A}}(N)$ for $N_{i,l}$ in $N$ do
3             for $x_{j}$ in $X$ do
4                   $\hat{z}_{j}\leftarrow\hat{z}_{i}^{(l)}(x_{j})$ if $\hat{z}_{j}>0$  then
5                         if  $|\mathbf{F}(\hat{z}_{j})|<\beta$ and $|\frac{\partial\mathbf{F}}{\partial\hat{z}_{i}^{(l)}}|_{\hat{z}_{j}}|>\gamma$ and $P_{i,l}==\mathbf{0}$ then
                               $Mand\leftarrow Mand\cup\{N_{i,l}\}$   /* Neuron is in state $\mathbf{0}$, can’t coarsen */
6
7                   else
8                         if $\mathbf{F}(\hat{z}_{j})<0$ and $P_{i,l}==\mathbf{1}$ then
                               $Mand\leftarrow Mand\cup\{N_{i,l}\}$   /* Neuron is in state $\mathbf{1}$, can’t coarsen */
9
10
11
12
13      return $Mand$
Algorithm 4 Gradient_Search
Both the estimation method Gradient_Search and Adversarial_Prune can be used together. Taking the intersection of their estimated results may provide us with a better overall estimate. In addition, the estimated mandatory neurons can serve as an initial starting point in our statistical versions of the Refine and Coarsen approaches introduced in Section 5.
Our experimental results regarding these two estimation approaches indicate a notable trend: neurons from deeper layers are more likely to be mandatory. This observation aligns with commonly-held beliefs that deeper neurons are generally linked to high-level feature representation, and thus play a more important role than shallow neurons in final classification decisions. This observation leads us to a valuable heuristic: prioritizing deeper neurons over shallower ones as we iterate through neurons in our implementations of the Refine and Coarsen algorithms.
5. Statistical Learning Approaches
In this section, we introduce two new algorithms that address the minimal NAP specification problem. While they share the same overarching concept as refine and coarsen, these methods employ sampling and statistical learning principles to efficiently learn a minimal NAP specification. Recall that the set of mandatory neurons $M$ is the union of all neurons that appears in any minimal specification and is an upper bound of the largest minimal specification.
5.1. The Statistical Refine Approach
Mandatory neurons are crucial for forming NAP specifications, as their binary states play a critical role in determining the neural network’s robustness performance. We leverages this property to find mandatory neurons statistically. To be more specific, suppose we sample some NAPs $P_{1},P_{2},\ldots,P_{n}$ from the NAP family $\mathcal{P}$. For those NAPs that qualify as specifications (i.e., $\mathcal{V}(P)=1$), mandatory neurons should appear more frequently in them than in those NAPs that fail the verification tool (i.e., $\mathcal{V}(P)=0$).
Based on this insight, we propose an approach called Sample_Refine that relies on non-repetitive sampling to identify mandatory neurons for solving the minimal NAP problem. We start with the coarsest NAP and iteratively collect the most probable mandatory neurons. In every iteration, we sample $k$ NAPs by refining unvisited neurons with some probability $\theta$. The sampled NAPs are then fed to the verification tool, and the neuron that appears most frequently in verifiable NAPs is the most probable mandatory neuron in this iteration. The neuron is marked as visited, and the process stops when either we collect $s$ neurons (assuming that $s$ is known), or the current mandatory neurons form a NAP specification $P$, i.e., $\mathcal{V}(P)=1$. Finally, we return the learned NAP, obtained by applying $\widetilde{\mathcal{A}}$ to the collected neurons. The algorithm 5 provides an overview of the above procedure.
It is worth noting that Sample_Refine doesn’t guarantee correctness, as we may end up collecting only the $s$ most probable mandatory neurons, which may not be sufficient to form a specification. Another concern regarding this algorithm is sampling efficiency, specifically the potential for the number of samples required to grow exponentially with the size of the minimal NAP specification $s$. To understand why this is problematic, consider a scenario where the only minimal NAP specification $P$ consists of all mandatory neurons; in this case, all $|M|$ neurons must be selected for $P$ to be learned. If $\theta$ is set to a constant value, then the expected number of samples needed to obtain the NAP specification is $(\frac{1}{\theta})^{|M|}$. To address this, we set $\theta=\left(\frac{|M|}{|M|+1}\right)^{|M|}$. This choice ensures that the sampling efficiency is polynomial in both $|M|$ and $s$, as proven in Theorem 5.1. Please refer to the proof in the Appendix B. In addition, Theorem 5.1 also shows that with high probability, a mandatory neuron will be found with $O(log|N|)$ calls to $\mathcal{V}$.
Theorem 5.1.
With probability $\theta=|(\frac{|M|}{|M|+1})|^{|M|}$, Sample_Refine has $1-\delta$ probability of outputting a minimal NAP specification with $\Theta\left(|M|^{2}(\log|N|+\log(s/\delta))\right)$ examples each iteration and $O(s|M|^{2}(\log|N|+\log(s/\delta)))$ total calls to $\mathcal{V}$.
5.2. The Statistical Coarsen Approach
The Coarsen algorithm begins with the most refined NAP and progressively coarsens each neuron until the verification process fails. Enhancing the algorithm’s performance is possible by coarsening multiple neurons during each iteration. However, a fundamental question emerges: How do we determine which set of neurons to coarsen in each round?
We present Sample_Coarsen to answer this question. In this approach, we assume that each neuron is independent of the others and select neurons to coarsen in a statistical manner. Specifically, in each iteration, we randomly coarsen a subset of refined neurons in the current NAP simultaneously to see if the new NAP can pass verification. We repeat this process until the NAP size reaches $s$. Algorithm 6 provides the pseudocode for Sample_Coarsen.
Input: The neural network $N$, the probability $\theta$, sample size $k$, and the size $s$
Output: A minimal NAP specification $P$
1 Function Sample_NAPs($unvisited$, $\theta$)
       $P\leftarrow\dot{\mathcal{A}}(N)$   /* Use the coarsest NAP as a blank template */
2       for $N_{i,l}$ in $unvisited$ do
3             $rand\leftarrow\text{random}(0,1)$ if $rand\leq\theta$ then
                   $P\leftarrow\widetilde{\Delta}(N_{i,l})$   /* Refine unvisited neurons using $\widetilde{\Delta}$ with probability $\theta$/
4
5
6      return $P$
7Function Sample_Refine($visited$, $\theta$, $k$, $s$)
8       while $|visited|<s\text{ and }\mathcal{V}(P)==0$  do
             $unvisited\leftarrow N\setminus visited$ $ctr=dict()$   /* Create a counter for each neuron in $unvisited$/
9             for _normal-_\__ in $range(k)$ do
                   $P\leftarrow Sample\_NAPs(unvisited,\theta)$   /* Sample $k$ NAPs */
10                   for $N_{i,l}$ in $unvisited$ do
11                         if $P_{i,l}==\widetilde{\mathcal{A}}(N_{i,l})$ and $\mathcal{V}(P)==1$ then
                               $ctr[N_{i,l}]$ += 1   /* Count successful NAPs for each neuron */
12
13
14
            $N_{mad}\leftarrow\operatorname*{argmax}_{N_{i,l}\in unvisited}\{ctr[N_{i,l}]\}$   /* Pick potential mandatory neuron */
15             $visited\leftarrow visited\cup\{N_{mad}\}$
16      return $\widetilde{\mathcal{A}}(visited)$
[…]
       return $None$   /* Return None if the most refined NAP fails verification */
18
19 else
20       if heuristics then
21             $visited\leftarrow Gradient\_Search(N)\cap Adversarial\_Prune(N)$
22       else
             $visited\leftarrow\emptyset$   /* Start from the coarsest NAP */
23
24      $Sample\_Refine(visited,\theta,k,s)$
Algorithm 5 Sample_Refine
Similar to Sample_Refine, Sample_Coarsen also faces the same challenge related to sample efficiency. To better illustrate this, suppose a minimal NAP specification $P$ with size $s$ can be found after one iteration of refinement. Then, the probability of selecting the exact number of $s$ mandatory neurons in $P$ is $\theta^{s}$. Consequently, if $\theta$ is set to a constant value, the expected number of samples needed to find the NAP is $(\frac{1}{\theta})^{s}$. Unlike Sample_Refine, we know the refinement will always ensure the resulting NAP can pass verification. So, once such a NAP is learned, we can narrow down the estimated mandatory neurons by an expected factor of $\theta$. In this way, the expected number of samples and the expected number of iterations are inversely related, yet their product is the total number of calls to $\mathcal{V}$. We demonstrate that by setting $\theta=e^{-\frac{1}{s}}$, we not only make the expected number of samples polynomial in $s$, but also minimize the total number of calls to $\mathcal{V}$, as proven in Theorem 5.2. Please refer to the proof in Appendix B.
Theorem 5.2.
With probability $\theta$=$e^{-\frac{1}{s}}$, Sample_Coarsen learns a minimal NAP specification with $\mathcal{O}(slog|N|)$ calls to $\mathcal{V}$.
Input: The neural network $N$, the probability $\theta$, and the size $s$
Output: A minimal NAP specification $P$
1 Function Sample_Coarsen($mand\_neurons$, $\theta$, $s$)
2       $P\leftarrow\widetilde{\mathcal{A}}(N)$; $mand\_neurons\leftarrow N$ if $\mathcal{V}(P)==0$ then
             return $None$   /* Return None if the most refined NAP fails verification */
3
4       else
5             while $|P|>s$ do
6                   $P\leftarrow Sample\_NAPs(mand\_neurons,\theta)$ if $\mathcal{V}(P)==1$  then
7                         $found\_neurons\leftarrow\emptyset$ for $N_{i,l}$ in $N$ do
8                               if $P_{i,l}==\widetilde{\mathcal{A}}(N_{i,l})$ then
9                                     $found\_neurons\leftarrow found\_neurons\cup\{N_{i,l}\}$
10
                        $mand\_neurons\leftarrow found\_neurons$   /* Reduce search space */
11
12                   else
                         $P\leftarrow Sample\_Naps(mand\_neurons,\theta)$   /* Sample a new NAP */
13
14
            return $P$   /* Return the minimal NAP of size $s$/
15
16
Algorithm 6 Sample_Coarsen
Setting the sample probability $\theta$
Setting $s$ poses a challenge in practice, as we assume that $s$ is always provided in Sample_Refine and Sample_Coarsen. However, this can be addressed by dynamically updating $\theta$ based on the result of $\mathcal{V}(P)$ (Liang et al., 2011). With $\theta$ from theorem 5.2, Sample_Coarsen finds a NAP specification with probability $\left(e^{-1/s}\right)^{s}=e^{-1}$. Recall that theorem 5.1 state that Sample_Refine finds a NAP specification with probability $\left(\frac{|M|}{|M|+1}\right)^{|M|}$. The probability approaches $e^{-1}$ asymptotically. Thus, we aim to set $\theta$ such that the $Pr(\mathcal{V}(P)=1)=e^{-1}$. Intuitively, if a sampled NAP $P$ is a specification, we decrease $\theta$ so less neurons will be refined or more neurons will be coarsened. Similarly, if $P$ is not a specification, $\theta$ needs to be increased.
Given that $\theta\in[0,1]$, we can parameterize it using the Sigmoid function $\sigma(\lambda)=\left(1+e^{-\lambda}\right)^{-1}$, where $\lambda\in(-\infty,\infty)$. Since $Pr(\mathcal{V}(P)=1)$ depends on $\theta$ as well, we express it as a function of $\lambda$, $g(\lambda)=Pr(\mathcal{V}(P)=1)$. Then, setting $Pr(\mathcal{V}(P)=1)=e^{-1}$ can be achieved through the following minimization problem:
(14) $L(\lambda)=\frac{1}{2}(g(\lambda)-e^{-1/s})^{2}$
The loss function $L(\lambda)$ can be minimized by statistical learning using stochastic gradient descent. With a step size $\eta$, update $\lambda$ using $\lambda\leftarrow\lambda-\eta\frac{dL}{d\lambda}$. Note that $\frac{dL}{d\lambda}$ can be expressed as:
(15) ${\frac{dL}{d\lambda}}=(g(\lambda)-e^{-1/s})\frac{dg(\lambda)}{d\lambda}$
Given $g(\lambda)=Pr(\mathcal{V}(P)=1)$, we can replace $g(\lambda)$ with $\mathcal{V}(P)$ for stochastic gradient update. Additionally, since $\frac{dg(\lambda)}{d\lambda}>0$, we simply ignore it as its multiplication effect can be represented by $\eta$. Therefore, the final update rule is given by:
(16) $\lambda\leftarrow\lambda-\eta(\mathcal{V}(P)-e^{-1})$
6. Volume Estimation of $R_{P}$
Figure 7. Volume Estimation of $R_{P}$ using an orthotope in a 2-dimensional case. The gray rectangle represents the input space, with the training set depicted by a collection of data points $+$. The polygon corresponds to some NAP $P$. Initially, we identify an anchor point, denoted by $+$. Then, we construct the orthotope, represented by the green rectangle, by extending upper and lower bounds starting from the anchor point until it extends beyond $P$.
Conceptually, NAP specifications typically correspond to significantly larger input regions compared to local neighbourhood specifications. This serves as the primary motivation for utilizing NAPs as specifications. However, previous work lacks sufficient justification or evidence to support this claim. In this section, we propose a simple method for approximating the volume of $R_{P}$, i.e., the region corresponding to a NAP $P$. This allows us to: 1) quantify the size difference between $R_{P}$ and $L_{\infty}$ ball specifications; 2) gain insights into the volumetric change from the most refined NAP specification to the minimal NAP specification.
Computing the exact volume of $R_{P}$ is at least NP-hard, as determining the exact volume of a polygon is known to be NP-hard (Dyer and Frieze, 1988). Moreover, computing the exact volume of $R_{P}$ can be even more challenging due to its potential concavity. To this end, our method estimates the volume of $R_{P}$ by efficient computation of an orthotope that closely aligns with $R_{P}$, as illustrated in Figure 7. We briefly describe it as follows:
Finding an anchor point
The first step is to find an anchor point to serve as the center of the orthotope. Ideally, this anchor point should be positioned close to the center of $R_{P}$ to ensure a significant overlap between the orthotope and $R_{P}$. However, computing the actual center of $R_{P}$ is costly. Thus, we look for a pseudo-center from the training set $X$ that resides in $R_{P}$. This pseudo-center can be computed by finding the point that uses the smallest $L_{\infty}$ ball to cover other data points, solved as the following optimization problem:
$c_{\text{pseudo}}=\underset{x\in R_{P}}{\arg\min}\max_{x^{\prime}\in R_{P}}\|x-x^{\prime}\|_{\infty}$
where $R_{P}=\{x\mid\mathcal{A}(N,x)\preccurlyeq P,x\in X\}$. When $|X|$ is small, $c_{\text{pseudo}}$ can be computed directly; for larger $|X|$, a statistical computation strategy is required.
Constructing the orthotope
Once the pseudo-center $c_{\text{pseudo}}$ is determined, we want to create an orthotope around $c_{\text{pseudo}}$ to closely align with $R_{P}$. The orthotope is constructed by determining pairs of upper and lower bounds $U^{(i)}$ and $L^{(i)}$ for each dimension $i$. Specifically, $U^{(i)}$ and $L^{(i)}$ are computed through expansion in two opposite directions from $c_{\text{pseudo}}$ along dimension $i$ until they extend beyond $R_{P}$. This expansion can be expressed as:
$\max_{U^{(i)}}\{x^{\prime}\in R_{P}\,|\,x^{\prime}:=c_{\text{pseudo}}+U^{(i)}\}\text{ ; }\max_{L^{(i)}}\{x^{\prime}\in R_{P}\,|\,x^{\prime}:=c_{\text{pseudo}}-L^{(i)}\}$
Here, $U^{(i)}$ and $L^{(i)}$ represent the upper and lower bounds in dimension $i$ respectively, originating from $c_{\text{pseudo}}$. These bounds can be efficiently calculated with binary search.
The choice of the archer point is crucial in our approach. If it is located at a corner of $R_{p}$, the volume calculation will be highly biased. This can pose a problem when we seek to understand the volumetric change from the most refined NAP specification to the minimal NAP specification. Additionally, using the orthotope as an estimator provides convenience in understanding the volumetric change simply by examining differences in each input dimension.
7. Evaluation
In this section, we conduct a comprehensive evaluation of our methods for learning minimal NAP specifications across a range of benchmarks, spanning from a simple binary classifier to the state-of-the-art image classification model. We illustrate that minimal NAP specifications usually involve only a fraction of neurons compared to the most refined NAPs, yet they dramatically extend the verifiable bounds to several magnitudes larger.
Experiment Setup
All experiments in this section were conducted on an Ubuntu 20.04 LTS machine with 172 GB of RAM and an Intel(R) Xeon(R) Silver Processor. For verification, we utilized Marabou (Katz et al., 2019), a dedicated state-of-the-art neural network verifier. We configured a timeout of 10 minutes for each call to the verification tool. If the timeout is exceeded, the current neuron is retained in the minimal NAP specification even if its status cannot be determined.
[…]
0 1
$P$$\mathcal{V}$ train test $P$$\mathcal{V}$ train test
The $\widetilde{\mathcal{A}}$ function 102 - 78.11 81.40 93 - 87.06 80.28
Coarsen 31 102 98.22 95.35 32 93 99.65 94.37
Adversarial_Prune 61 - 81.06 88.37 54 - 87.06 80.28
Gradient_Search 43 - 82.84 88.37 39 - 93.37 87.32
Sample_Refine 53 187 92.43 89.37 56 264 89.86 88.51
Sample_Coarsen 42 47 94.69 92.34 45 41 94.15 91.52
Wisconsin Breast Cancer with Binary Classifier
We conduct our first experiment using a four-layer neural network as a binary classifier, where each layer consists of 32 neurons. This classifier is trained on the Wisconsin Breast Cancer (WBC) dataset (William et al., 1995), representing a decision-critical task where robustness is essential. Our trained model achieves a test set accuracy of 95.61%. We calculate the most refined NAP specifications $\widetilde{P}^{0}$ and $\widetilde{P}^{1}$ for labels 0 and 1 using the statistical abstraction function $\widetilde{\mathcal{A}}$ with a confidence ratio of $\delta=0.95$. The size of $\widetilde{P}^{0}$ and $\widetilde{P}^{1}$ are 102 and 93, respectively. In contrast, the sizes of the minimal NAP specifications learned by the Coarsen algorithm for labels 0 and 1 are significantly reduced to 31 and 32, respectively. It is worth mentioning that our estimation approaches provide a fairly accurate estimate of the mandatory neurons, despite computing rather loose upper bounds. To be more specific, Adversarial_Prune and Gradient_Search compute 61 and 43 mandatory neurons for label 0, respectively. Together, they cover 25 out of the 31 mandatory neurons appearing in the minimal NAP specification for label 0. For label 1, Adversarial_Prune and Gradient_Search compute 54 and 39 mandatory neurons, respectively, covering 25 out of the 32 mandatory neurons appearing in the minimal NAP specification for label 1.
Regarding the statistical approaches, Sample_Refine computes NAP specifications of size 53 and 56 for label 0 and label 1, respectively. This is achieved while making 187 and 264 calls, respectively. Although these numbers are quite large, it is expected since it has to sample multiple trials in each iteration. In contrast, Sample_Coarsen is significantly more efficient. It can learn NAP specifications of size 42 and 45 for label 0 and label 1 using only 47 and 41 calls, respectively.
Recall that one of the main motivations for learning the minimal NAP specifications is their potential to verify larger input regions compared to the refined NAP specifications. To support this claim, we compute the percentile of unseen test data that can be verified using these NAPs. We choose test data as they are sampled from the input space and can be viewed as a proxy to understand the verifiable bounds of different NAPs. We observe that the most refined NAP specifications $\widetilde{P}^{0}$ and $\widetilde{P}^{1}$ can cover 81.40% and 80.28% of the test data for labels 0 and 1, respectively. In contrast, this coverage significantly increases to 95.35% and 94.37% of the test data for the minimal NAP specifications of labels 0 and 1, respectively. To intuitively understand the extent of change in verifiable regions from the refined NAP specifications to the minimal NAP specifications, we also compare their estimated volume of the corresponding region $R_{P}$. The estimated volumetric changes are substantial, with label 0 experiencing an increase 58324.35 times larger and label 1 showing an increase 3053.53 times larger, respectively. To comprehend this change in each dimension, Figure 8 illustrates the comparison of verifiable input ranges between the refined NAP specifications and the minimal NAP specifications for labels 0 and 1, referencing the anchor point.
(a) Comparing verifiable input ranges for label 0
(b) Comparing verifiable input ranges for label 1
Figure 8. Comparison of verifiable input ranges between the refined NAP specifications and the minimal NAP specifications for labels 0 and 1 on the Wisconsin Breast Cancer dataset, with reference to the anchor point. The minimal NAP specifications enable verification over a larger input range for all dimensions.
Table 2. Overview of the size of learned minimal NAP specifications on the MNIST benchmark.
0 1 4
$P$$\mathcal{V}$ train test $P$$\mathcal{V}$ train test $P$$\mathcal{V}$ train test
The $\widetilde{\mathcal{A}}$ function 751 - 79.50 80.51 745 - 86.01 85.11 712 - 77.54 80.24
Coarsen 480 751 98.68 98.78 491 745 98.90 98.59 506 712 98.51 97.45
Adversarial_Prune 618 - 83.13 71.32 630 - 79.02 76.51 699 - 82.66 84.12
Gradient_Search 195 - 92.1 91.80 237 - 89.73 88.39 169 - 87.59 88.51
Sample_Refine 564 4059 90.03 87.91 570 3663 88.36 86.84 581 4532 82.90 83.46
Sample_Coarsen 532 33 93.19 93.01 559 27 94.12 93.68 562 25 93.89 93.64
MNIST with Fully Connected Network
To show that our insights and approaches can be applied to more complicated datasets and networks, we conduct the second set of experiments using the mnistfc_256x4 model (VNNCOMP, 2021), a 4-layer fully connected network with 256 neurons per layer trained on the MNIST dataset. We specifically focus on classes 0, 1, and 4, as the most refined NAPs for other labels either encounter timeouts or fail the verification. We compute the most refined NAP specifications $\widetilde{P}^{0}$, $\widetilde{P}^{1}$, and $\widetilde{P}^{4}$ for labels 0, 1, and 4 using the $\widetilde{\mathcal{A}}$ function with a confidence ratio of $\delta=0.99$. The size of $\widetilde{P}^{0}$, $\widetilde{P}^{1}$, and $\widetilde{P}^{4}$ are 751, 745 and 712, respectively, which aligns with the results from (Geng et al., 2023).
In contrast, the sizes of the minimal NAP specifications learned by the Coarsen algorithm for labels 0, 1, and 4 are significantly reduced to 480, 491, and 506, respectively. Moreover, our estimation approaches provide highly accurate estimates of the mandatory neurons present in minimal NAP specifications. For instance, considering label 0, Adversarial_Prune and Gradient_Search identify 618 and 195 mandatory neurons, respectively. And 445 and 160 out of the 480 mandatory neurons presented in the minimal NAP specification are discovered by Adversarial_Prune and Gradient_Search, respectively.
[…]
Moreover, the learned minimal NAP specifications correspond to significantly larger verifiable regions compared to the refined NAP specifications. Using the percentile of test data as a metric, the minimal NAP specifications increase the coverage ratio from 80.51% to 98.78%, 85.11% to 98.59%, and 80.24% to 97.45% for labels 0, 1, and 4, respectively. From the perspective of estimated volume, the volumetric changes are on the order of $10^{8}$ times larger.
ImageNet with Deep Convolutional Neural Networks
Current verification approaches struggle to scale to state-of-the-art neural networks such as deep convolutional neural networks, making them inapplicable for verifying certain critical systems like deep learning-based autonomous driving software. Consequently, our verification-dependent approaches also face scalability challenges. However, we can still assess the network’s robustness from a neuron-centric perspective using our estimation approaches, an aspect not well-studied in the current literature.
Thus, for the third experiment, we choose a deep convolutional network, specifically VGG-19 trained on the ImageNet dataset, as the benchmark for estimating mandatory neurons. We select neurons from fully connected layers [^1] We leave NAP computation for other layers such as convolutional layers for future work. of VGG-19, which contains a total of 8192 neurons, to compute NAPs. The original ImageNet includes 1000 classes with millions of images. We also narrow our focus to the top five largest classes in ImageNet. Each class consists of around 1000 training images and 300 test images. Table 3 presents the results of the learned NAP using verification-free approaches.
To show that these estimated mandatory neurons are indeed meaningful, we employ several approaches from the model interpretability domain such as Grad-CAM (Selvaraju et al., 2017) and saliency maps (Simonyan et al., 2014a). We observe a strong correlation between neurons exhibiting high interpretability under certain metrics and our estimated mandatory neurons computed by Adversarial_Prune and Gradient_Search. However, we also find that some mandatory neurons do not seem to participate significantly in interpretability, yet they are indeed necessary for the robustness of models. We find this interesting and plan to investigate those neurons in future work.
Although we cannot definitively determine if these NAPs can serve as specifications, these NAPs do cover significant portions of unseen test data. For example, the NAP formed by mandatory neurons learned through Gradient_Search can cover $87.01\%$ of unseen test data. This aspect makes them a promising basis for computing NAP specifications, especially if verification tools become accessible for such large models in the future.
Table 3. Overview of the size of learned minimal NAP specifications on the ImageNet benchmark.
box_turtle spiny_lobster acorn_squash confectionery stone_wall
$P$ test $P$ test $P$ test $P$ test $P$ test
The $\widetilde{\mathcal{A}}$ function 1978 39.42 1006 39.61 1003 29.23 878 33.84 971 31.54
Adversarial_Prune 1863 39.42 955 39.61 823 25.68 845 23.07 865 39.61
Gradient_Search 611 53.30 572 87.01 301 52.40 256 44.23 260 54.90
8. Related Work
8.1. Neural Network Verification
Neural network verification has attracted much attention due to the increasingly widespread applications of neural networks in safety-critical systems. it’s NP-hard nature resulting from the non-convexity introduced by activation functions (Katz et al., 2017) makes it a challenging task. Thus most of the existing work on neural network verification focuses on designing scalable verification algorithms. For instance, while initially proposed solver-based approaches (Huang et al., 2017b; Ehlers, 2017; Cheng et al., 2017; Tjeng et al., 2019) were limited to verifying small-scale neural networks with fewer than 100 neurons, state-of-the-art approaches (Xu et al., 2021; Wang et al., 2021c; Lu and Kumar, 2019) can verify more complex neural networks. It is worth mentioning that most existing work adopts local neighborhood specifications to verify the robustness properties of neural networks (Shapira et al., 2023). Despite being a reliable measure, using specifications that define local neighborhood of reference data points may not cover any test data, let alone generalizing to the verification of unseen test set data. Yet evidently, being able to verify the neural network on test data is an important task in real-world applications. The development of NAP specification (Geng et al., 2023) sheds light on this challenge. While previous work on NAP specifications proves its effectiveness, it assumes that all neurons contribute equally to the model’s robustness. In this paper, we challenge this assumption by demonstrating that only a small fraction of neurons are crucially related to robustness. This not only allow us study robustness from the neurons’ perspective but also significantly pushes the known verifiable boundaries.
8.2. Abstract Interpretation
Abstract interpretation (Cousot and Cousot, 1977) is a fundamental concept in software analysis and verification, particularly for approximating the semantics of discrete program states. By sacrificing precision, abstract interpretation typically enables scalable and faster proof finding during verification (Cousot and Cousot, 2014). Although abstract interpretation for neural network verification has been proposed and studied in previous literature (Gehr et al., 2018; Mirman et al., 2018), abstract interpretation of neural activation patterns for verification is a relatively new field. Perhaps the most related work from the perspective of abstract interpretation is learning minimal abstractions (Liang et al., 2011). While our work shares similarities in problem formulation and statistical approaches, we address fundamentally different problems. One limitation in our work is that our abstraction states may be too coarse: value in range $(0,+\infty)$ is abstracted into one state. This approach could over-approximate neuron behavior and thus fail to prove certain properties. We observe that neuron values exhibit different patterns in range for different input classes, suggesting the potential existence of more abstraction states. We plan to explore this direction in our future work.
8.3. Neural Activation Patterns
Neural activation patterns have commonly been used to understand the internal decision-making process of neural networks. One popular line of research is feature visualization (Yosinski et al., 2015; Bäuerle et al., 2022), which investigates which neurons are activated or deactivated given different inputs. This is also naturally related to the field of activation maximization (Bau et al., 2017; Simonyan et al., 2014b; Erhan et al., 2009), which studies what kind of inputs could mostly activate certain neurons in the neural network. In this way, certain prediction outcomes may be attributed to the behavior of specific neurons, thereby increasing the interpretability of the underlying models. In summary, most of existing works focus on learning statistical correlations between either NAPs and inputs or NAPs and prediction outcomes. However, whether such correlations can be trusted or even verified is not fully studied. To this end, we propose the concept of mandatory neurons and highlight their importance in the robustness of model predictions. Such causal links between neurons and prediction outcomes are not only identified but also verified. We believe this ”identify then verify” paradigm can be extended to existing research on NAPs to certify our understanding of neural networks. And we leave the exploration of this direction for our future work.
9. Conclusion
We introduce a new problem — learning the minimal NAP specification, and discuss its importance in neural network verification. Finding minimal NAP specification not only enables the verification of larger input regions compared to existing methods but also provides a means of inspecting when and how neural networks can make reliable and robust predictions. To solve this problem, we first propose two simple approaches Refine and Coarsen, which leverage off-the-shelf verification tools that find the minimal NAP specification with correctness guarantees. We also propose the statistical version of Refine and Coarsen, that combine sampling and statistical learning principles to achieve correctness in an more efficient manner. However, these approaches depend on underlying verification tools that are computationally expensive. To this end, we propose two approximate approaches, Adversarial_Prune and Gradient_Search. These methods utilize adversarial attacks and local gradient computation to efficiently investigate potential causal relationships between specific neurons and the model’s robustness. This is a task that current verification frameworks struggle to accomplish. Finally, to appreciate the volumetric change from the most refined NAP specification to the minimal NAP specification, we propose a simple method for estimating the volume of the region corresponding to a NAP. Our experiments indicate that minimal NAP specifications utilize much smaller fractions of neurons compared to the most refined NAPs. Nevertheless, they enable a substantial expansion of the verifiable boundaries by several orders of magnitude.
Appendix A Proofs of simple approaches
Theorem A.1 (Property of Refine).
The algorithm Refine returns a minimal NAP specification with $\mathcal{O}(2^{|N|})$ calls to $\mathcal{V}$.
Proof.
Let $P$ be the returned NAP. We prove this by contradiction. Suppose we can further refine $P$, meaning there exists a NAP $P^{\prime}$ such that $|P^{\prime}|\leq|P|$ and $\mathcal{V}(P^{\prime})=1$. However, the algorithm states that any $P^{\prime}$ smaller than $|P|$ fails verification, which contradicts $\mathcal{V}(P^{\prime})=1$.
In the worst case, the NAP size $k$ runs up to $|N|$. For each $k$, we need to check $\binom{|N|}{k}$ number of NAPs. In total, this number of NAPs we need to check is $\binom{|N|}{1}+\binom{|N|}{2}+\cdots+\binom{|N|}{|N|}=2^{|N|}-1$ according to the binomial theorem, resulting in a runtime complexity of $\mathcal{O}(2^{|N|})$.
Theorem A.2 (Property of Coarsen).
The algorithm Coarsen returns a minimal NAP specification with $\mathcal{O}(|N|)$ calls to $\mathcal{V}$.
Proof.
Let $P$ be the NAP returned by Coarsen. Our goal is to show that any $P^{\prime}$ smaller than $P$ results in $\mathcal{V}(P^{\prime})=0$. To construct such a smaller $P^{\prime}$, we need to apply the refine action $\widetilde{\Delta}$ on $P$ through some neuron $N_{i,l}$, i.e., $P^{\prime}:=\widetilde{\Delta}(N_{i,l})=\widetilde{\mathcal{A}}(N_{i,l})\cup P\setminus P_{i,l}$. According to the algorithm, $\mathcal{V}(P^{\prime})=0$. In the worst case, the algorithm needs to iterate through each neuron in $N$, resulting in a runtime complexity of $\mathcal{O}(|N|)$. ∎
Appendix B Proofs of statistical approaches
Our proofs of properties of Sample_Refine and Sample_Coarsen mainly follow those in (Liang et al., 2011). Interested readers may refer to it for detailed proofs.
Theorem B.1 (Property of Sample_Refine).
With probability $\theta=|(\frac{|M|}{|M|+1})|^{|M|}$, Sample_Refine has $1-\delta$ probability of outputting a minimal NAP specification with $\Theta\left(|M|^{2}(\log|N|+\log(s/\delta))\right)$ examples each iteration and $O(s|M|^{2}(\log|N|+\log(s/\delta)))$ total calls to $\mathcal{V}$.
Proof.
Considering the size of the minimal specification $s$, Sample_Refine will execute $s$ iterations. If we sample $k$ times in each iteration, then the probability of selecting a mandatory neuron is at least $1-\frac{\delta}{s}$. Consequently, by applying a union bound, the algorithm will identify a NAP specification with a probability of at least $1-\delta$.
Now, let’s delve deeper into one iteration of the process. The fundamental concept is that a mandatory neuron $m$ demonstrates a stronger correlation with proving the robustness query ($\mathcal{V}(P)=1$) compared to a non-mandatory one. This enhanced correlation increases the probability of its selection significantly when $k$ is sufficiently large.
Let’s revisit the notion of $M$, representing the set of mandatory neurons with a size of $|M|$. Now, let’s focus on a specific mandatory neuron $n^{+}\in M$. We define $B_{j-}$ as the event indicating $k_{n^{-}}>k_{n^{+}}$, and $B$ as the event where $B_{j-}$ holds for any non-mandatory neuron $n^{-}\in N\setminus M$. Importantly, if $B$ fails to occur, then the algorithm will correctly identify a mandatory neuron. Hence, our primary objective is to establish that $Pr(B)\leq\frac{\delta}{s}$. Initially, employing a union bound provides:
(17) $Pr(B)\leq\sum_{n^{-}}Pr(B_{n^{-}})\leq|N|\max_{n^{-}}Pr(B_{n^{-}})$
Let’s delve into each training example $P^{(i)}$ and introduce the notation $X_{i}=(1-\mathcal{V}(P^{(i)}))(P_{n^{-}}^{(i)}-P_{n^{+}}^{(i)})$. It’s worth emphasizing that $B_{n^{-}}$ manifests precisely when $\frac{1}{n}(k_{n^{-}}-k_{n^{+}})=\frac{1}{n}\sum_{i=1}^{n}X_{i}>0$. Our objective now is to bound this quantity utilizing Hoeffding’s inequality, considering the mean as:
(18) $\mathbb{E}[X_{i}]=Pr(\mathcal{V}(P)=1,P_{n^{-}}\in\{\mathbf{1},\mathbf{0}\})-Pr(\mathcal{V}(P)=1,P_{n^{+}}\in\{\mathbf{1},\mathbf{0}\})$
and the bounds are $-1\leq\mathbb{E}[X_{i}]\leq 1$ Setting $\epsilon=-\mathbb{E}[X_{i}]$, we get:
(19) $Pr(B_{n^{-}})\leq e^{\frac{-k\epsilon^{2}}{2}},n^{+}\in M,n^{-}\notin M.$
Substituting (16) into (15) and rearranging terms, we can solve for $k$:
(20) $\frac{\delta}{s}\leq|N|e^{\frac{-k\epsilon^{2}}{2}}\text{ implies}k\geq\frac{2(log|N|+log(\frac{s}{\delta}))}{\epsilon^{2}}$
Our attention now shifts towards deriving a lower bound for $\epsilon$, which intuitively reflects the discrepancy (in terms of correlation with proving the robustness query) between a mandatory neuron and a non-mandatory one. It’s noteworthy that $Pr(P_{n}\in{\{\mathbf{1},\mathbf{0}}\})=\theta$ for any $n\in N$. Furthermore, since $n^{-}$ is non-mandatory, we can assert that $Pr(\mathcal{V}(P)=1|P_{n^{-}}\in\{\mathbf{1},\mathbf{0}\})=Pr(\mathcal{V}(P)=1)$. By leveraging these observations, we can express:
(21) $\epsilon=\theta(Pr(\mathcal{V}(P)=1|P_{n^{+}}\in\{\mathbf{1},\mathbf{0}\})-Pr(\mathcal{V}(P)=1))$
Let’s view $C$ as the collection of minimal NAP specifications. We can treat $C$ as a set of clauses in a Disjunctive Normal Form (DNF) formula: $\mathcal{V}(P;C)=\neg\bigvee_{c\in C}\bigwedge_{n\in c}P_{n}$, where we explicitly specify the dependence of $\mathcal{V}$ on the clauses $C$. For example, if $C={{1,2},{3}}$, it corresponds to $\mathcal{V}(P)=\neg[(P_{1}\land P_{2})\lor P_{3}]$. Now, let $C_{j}={c\in C:n\in c}$ denote the clauses containing $n$. We reformulate $Pr(\mathcal{V}(P)=1)$ as the sum of two components: one originating from the mandatory neuron $n^{+}$ and the other from the non-mandatory neuron:
(22) $Pr(\mathcal{V}(P)=1)$ $=Pr(\mathcal{V}(P;C_{n^{+}})=1,\mathcal{V}(P;C\setminus C_{n^{+}})=0)$
(23) $\quad+Pr(\mathcal{V}(P;C\setminus C_{n^{+}})=1)$
Calculating $Pr(\mathcal{V}(P)=1|P_{n^{+}}\in\{\mathbf{1},\mathbf{0}\})$ follows a similar process. The only distinction arises from conditioning on $P_{n^{+}}\in\{\mathbf{1},\mathbf{0}\}$, which introduces an extra factor of $\frac{1}{\theta}$ in the first term because conditioning divides by $Pr(P_{n^{+}})=\theta$. The second term remains unchanged since no $c\notin C_{n^{+}}$ is mandatory in $P_{n^{+}}$. Substituting these two outcomes back into the equation yields:
(24) $\epsilon=(1-\theta)Pr(\mathcal{V}(P;C_{n^{+}})=1,\mathcal{V}(P;C\setminus C_{n^{+}})=0)$
Now, our objective is to establish a lower bound for (22) across all possible $\mathcal{V}$ (equivalently, $C$), where $n^{+}$ is permitted to be mandatory in $C$. Interestingly, the worst possible $C$ can be obtained by either having $|M|$ disjoint clauses ($C={{n}:n\in M}$) or a single clause ($C={M}$ if $s=|M|$). The intuition behind this is that when $C$ consists of $|M|$ clauses, there are numerous possibilities ($|M|-1$ of them) for some $c\notin C_{n^{+}}$ to be true, making it challenging to determine that $n^{+}$ is a mandatory neuron; in such cases, $\epsilon=(1-\theta)\theta(1-\theta)^{|M|-1}$. Conversely, if $C$ comprises a single clause, then letting this clause be true becomes exceedingly challenging; in this scenario, $\epsilon=(1-\theta)\theta^{|M|}$.
Let’s consider the scenario where $C$ comprises $|M|$ clauses. We aim to maximize $\epsilon$ concerning $\theta$ by setting the derivative $\frac{d\epsilon}{d\theta}=0$ and solving for $\theta$. This optimization yields $\theta=\frac{1}{|M|+1}$ as the optimal value. Substituting this value into the formula of $\epsilon$, we obtain $\epsilon=\frac{1}{|M|+1}(\frac{|M|}{|M|+1})^{|M|}$. Note that $(\frac{|M|}{|M|+1})^{|M|}$ can be lower bounded by $e^{-1}$, implying that $\epsilon^{-2}=\mathcal{O}(|M|^{2})$. Substituting this result to equation (18) will conclude the proof. ∎
Theorem B.2 (Property of Sample_Coarsen).
With probability $\theta$=$e^{-\frac{1}{s}}$, Sample_Coarsen learns a minimal NAP specification with $\mathcal{O}(slog|N|)$ calls to $\mathcal{V}$.
Proof.
Let’s first estimate the number of calls that Sample_Coarsen makes to $\mathcal{V}$. We denote the number of calls as $\mathcal{C}(P^{L})$, where $P^{L}$ is the most refined NAP. Then $\mathcal{C}(P^{L})$ can be computed recursively using the following rule:
(25) $\mathcal{C}(P^{L})=\begin{cases}|P|&\text{ if }|P^{L}|\leq s+1\\ 1+\mathbb{E}[(1-\mathcal{V}(P))\mathcal{C}(P)+\mathcal{V}(P)\mathcal{C}(P^{L})]&\text{ otherwise }\\ \end{cases}$
where $P$ is the sampled NAP. By assumption, there exists a NAP $P^{S}\preccurlyeq P$ of size $s$ that passes the verification. Define $G(P)=\neg(P^{S}\preccurlyeq P)$, which is 0 when $P^{S}$ subsumes $P$, i.e., all mandatory neurons in $P^{S}$ shows up in the sampled $P$. This follows that $Pr(G(P)=0)=Pr(P^{S}\preccurlyeq P)=\theta^{s}$. Note that $G(P)\geq\mathcal{V}(P)$, as the NAP $P^{S}$ suffices to prove the robustness query. We can estimate the upper bound of $\mathcal{C}(P^{L})$ by replacing $\mathcal{V}$ with $G$:
(26) $\mathcal{C}(P^{L})$ $\leq 1+\mathbb{E}[(1-G(P))\mathcal{C}(P)+G(P)\mathcal{C}(P^{L})]$
(27) $\leq 1+\theta^{s}\mathbb{E}[\mathcal{C}(P)|P^{S}\preccurlyeq P]+(1-\theta^{s})\mathcal{C}(P^{L})$
(28) $\leq\mathbb{E}[\mathcal{C}(P)|P^{S}\preccurlyeq P]+\theta^{-s}$
We now denote $\mathcal{C}(n)=max_{|P|=n}\mathcal{C}(P)$ as the maximum over NAP of size $n$. Note that given $P^{S}\preccurlyeq P$, $|P|=s+N$ where $N$ is a binomial random variable with $\mathbb{E}(N)=\theta(n-s)$.
Using the bound $\mathcal{C}(n)\leq(1-\theta^{n})\mathcal{C}(n-1)+\theta^{n}\mathcal{C}(n)\theta^{-s}$, we can observe that $\mathcal{C}(n)\leq\frac{\theta^{-s}}{1-\theta^{n}}\cdot n$. In addition, when $n$ is large enough, $\mathcal{C}(n)$ is concave, then by use Jensen’s inequality:
(29) $\mathcal{C}(n)\leq\mathcal{C}(\mathbb{E}[s+N])+\theta^{-s}=\mathcal{C}(s+\theta(n-s))+\theta^{-s}$
To solve the recurrence, this gives us:
(30) $\mathcal{C}(n)\leq\frac{\theta^{-s}logn}{log\theta^{-1}}+s+1$
The equation above illustrates a tradeoff between reducing the number of iterations (by increasing $\log\theta^{-1}$) and reducing the number of samples (by decreasing $\theta^{-s}$). To minimize $\mathcal{C}(n)$, we need to set $\theta$ so that the gradient of $\mathcal{C}(n)$ w.r.t $\theta^{-1}$ is 00. This gives us: $\frac{sx^{s-1}}{\log x}-\frac{x^{s-1}}{\log^{2}x}=0$, solving this gives us $\theta=e^{-\frac{1}{s}}$. Consequently, the upper bound becomes $\mathcal{C}(n)=es\log n+s+1=O(s\log n)$. ∎

### v2

@@
 In contrast, the NAP specification enables verification of inputs that are semantically close despite their geometric distance being far apart, such as image $c$.
-While this type of specification provides a fair and effective testbed for evaluating neural networks’ robustness, those verifiable $L_{\infty}$ balls are usually too tight, making them unlikely to cover unseen test set data (Geng et al., 2023).
+While this type of specification provides a fair and effective testbed for evaluating neural networks’ robustness, those verifiable $L_{\infty}$ balls are usually too tight, making them unlikely to cover unseen test set data (Geng et al., 2023a).
 For instance, the $L_{\infty}$ maximum verifiable bounds in the MNIST dataset is smaller than 0.2, whereas the smallest distance among images sharing the same label is greater than 0.5.
@@
 To this end,  Geng et al.
-(2023) proposes using neural activation patterns (NAPs) as specifications.
+(2023a) proposes using neural activation patterns (NAPs) as specifications.
 Neural activation patterns can be viewed as an abstraction of each neuron’s state.
@@
 […]
+0.55-
 […]
@@
 […]
+0.55-
 1011
@@
 Geng et al.
-(2023) argues that class NAPs must satisfy several essential requirements to qualify as NAP specifications.
+(2023a) argues that class NAPs must satisfy several essential requirements to qualify as NAP specifications.
 And finding qualified NAP specifications implies the underlying robustness problem is verified.
@@
 […]
+0.55-
 Pc2
@@
 […]
+0.55-
 […]
@@
 Interested readers can refer to Geng et al.
-(2023) for more details.
+(2023a) for more details.
 3. The minimal NAP specification problem
@@
 We present the semantics of these actions as follows:
+[…]
 $P\vdash\cdot\Downarrow P$
-  $P\vdash\Delta_{1}\Downarrow P^{\prime}$         $P^{\prime}\vdash\Delta_{2}\Downarrow P^{\prime\prime}$              $P\vdash\Delta_{1};\Delta_{2}\Downarrow P^{\prime\prime}$
-                           $P_{i,l}=\mathbf{*}$ Refine              $P\vdash\widetilde{\Delta}(N_{i,l})\Downarrow\widetilde{\mathcal{A}}(N_{i,l})\cup P\setminus P_{i,l}$                                 $P_{i,l}\in\{\mathbf{0},\mathbf{1}\}$ Coarsen                $P\vdash\dot{\Delta}(N_{i,l})\Downarrow\dot{\mathcal{A}}(N_{i,l})\cup P\setminus P_{i,l}$
+[…]
+[…]
+[…]
+[…]
 where $P,P^{\prime},P^{\prime\prime}$ are from the NAP family $\mathcal{P}$, and $N$ is the underlying neural network.
@@
 The algorithm Coarsen returns a minimal NAP specification with $\mathcal{O}(|N|)$ calls to $\mathcal{V}$.
+1
 Input: The neural network $N$
 Output: A minimal NAP specification $P$
-1 Function Refine(N)
+2
+3Function Refine(N)
 4       $P\leftarrow\widetilde{\mathcal{A}}(N)$
 5       if $\mathcal{V}(P)==0$ then
-             return $None$ ;   /* Return None if even the most refined NAP fails verification */
-4
+             return $None$ ;
+              /* Return None if even the most refined NAP fails verification */
+6
 7       else
 8             for $k$ from 1 to $|N|$ do
-                   $P\leftarrow\dot{\mathcal{A}}(N)$   /* Refinement starts from the coarsest NAP $\dot{P}$/
-                   $K\_comb\leftarrow$ $Pick\_K\_Neurons(k,N)$   /* Returns all combinations of size k */
+                   $P\leftarrow\dot{\mathcal{A}}(N)$
+                    /* Refinement starts from the coarsest NAP $\dot{P}$/
+                   $K\_comb\leftarrow$ $Pick\_K\_Neurons(k,N)$
+                    /* Returns all combinations of size k */
 9                   for $comb\textbf{ in }K\_comb$ do
 10                         for $N_{i,l}\textbf{ in }comb$ do
-                               $P\leftarrow\widetilde{\Delta}(N_{i,l})$  /* Refine neurons in the chosen combination */
-9
+                               $P\leftarrow\widetilde{\Delta}(N_{i,l})$
+                                /* Refine neurons in the chosen combination */
+11
 12                        if $|\mathcal{V}(P)|==1$ then
-                               return $P$   /* Found the minimal NAP */
-11
-12
+                               return $P$
+                                /* Found the minimal NAP */
 13
@@
 15
+16
+17
 Algorithm 1 refine
@@
 3       if $\mathcal{V}(P)==0$ then
-             return $None$ ;   /* Return None if the most refined NAP fails verification */
+             return $None$ ;
+              /* Return None if the most refined NAP fails verification */
 4
@@
 6             for $N_{i,l}$ in $N$ do
-                   $P\leftarrow\dot{\Delta}(N_{i,l})$ ;   /* Try to abstract this neuron using $\dot{\Delta}$/
+                   $P\leftarrow\dot{\Delta}(N_{i,l})$ ;
+                    /* Try to abstract this neuron using $\dot{\Delta}$/
 7                   if $\mathcal{V}(P)==0$ then
-                         $P\leftarrow\widetilde{\Delta}(N_{i,l})$ ;   /* Refine the neuron back if the verification fails */
+                         $P\leftarrow\widetilde{\Delta}(N_{i,l})$ ;
+                          /* Refine the neuron back if the verification fails */
 8
@@
 In this case, if we coarsen the second and fourth neurons, $N_{2,1}$ and $N_{4,1}$, the NAP specification will expand to $\langle\mathbf{1},\mathbf{*},\mathbf{*},\mathbf{*}\rangle$, which will cover the $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle$, thus failing the verification.
-In this case, $N_{2,1}$ and $N_{4,1}$ could both be mandatory neurons or either one of them is mandatory, as illustrated in Figures 4(b), 4(c), 4(d).
+In this case, $N_{2,1}$ and $N_{4,1}$ could both be mandatory neurons or either one of them is mandatory, as illustrated in Figures 5(b), 5(c), 5(d).
 So, we simply let $\{N_{2,1},N_{4,1}\}$ be the upper bound of mandatory neurons (learned from $x^{\prime}$).
@@
 Output: A collection of mandatory neurons
-1 Function Adversarial_Prune($N$, $X$)
-2       $Mand\leftarrow\emptyset$ ; $P\leftarrow\widetilde{\mathcal{A}}(N)$ for $x_{j}$ in $X$ do
-3             $x^{\prime}_{j}\leftarrow Adversarial\_Attack(x_{j})$ for $N_{i,l}$ in $N$ do
+1
+2Function Adversarial_Prune($N$, $X$)
+3       $Mand\leftarrow\emptyset$ ; $P\leftarrow\widetilde{\mathcal{A}}(N)$
+4       for $x_{j}$ in $X$ do
+5             $x^{\prime}_{j}\leftarrow Adversarial\_Attack(x_{j})$
+6             for $N_{i,l}$ in $N$ do
 7                   if $P_{i,l}\in\{\mathbf{0},\mathbf{1}\}$ and $\ddot{\mathcal{A}}(N_{i,l},x^{\prime}_{j})\oplus P_{i,l}$ then
-                         $Mand\leftarrow Mand\cup\{N_{i,l}\}$   /* Create mandatory neurons for $x^{\prime}_{j}$/
-5
-6
+                         $Mand\leftarrow Mand\cup\{N_{i,l}\}$
+                          /* Create mandatory neurons for $x^{\prime}_{j}$/
+8
+9
 10      return $Mand$
+11
 Algorithm 3 Adversarial_Prune
@@
 Output: A collection of mandatory neurons
-1 Function Gradient_Search($N$, $X$, $\beta$, $\gamma$)
-2       $Mand\leftarrow\emptyset$ ; $P\leftarrow\widetilde{\mathcal{A}}(N)$ for $N_{i,l}$ in $N$ do
+1
+2Function Gradient_Search($N$, $X$, $\beta$, $\gamma$)
+3       $Mand\leftarrow\emptyset$ ; $P\leftarrow\widetilde{\mathcal{A}}(N)$
+4       for $N_{i,l}$ in $N$ do
 5             for $x_{j}$ in $X$ do
-4                   $\hat{z}_{j}\leftarrow\hat{z}_{i}^{(l)}(x_{j})$ if $\hat{z}_{j}>0$  then
+6                   $\hat{z}_{j}\leftarrow\hat{z}_{i}^{(l)}(x_{j})$
+7                   if $\hat{z}_{j}>0$  then
 8                         if  $|\mathbf{F}(\hat{z}_{j})|<\beta$ and $|\frac{\partial\mathbf{F}}{\partial\hat{z}_{i}^{(l)}}|_{\hat{z}_{j}}|>\gamma$ and $P_{i,l}==\mathbf{0}$ then
-                               $Mand\leftarrow Mand\cup\{N_{i,l}\}$   /* Neuron is in state $\mathbf{0}$, can’t coarsen */
-6
+                               $Mand\leftarrow Mand\cup\{N_{i,l}\}$
+                                /* Neuron is in state $\mathbf{0}$, can’t coarsen */
+9
+10
 11                  else
 12                         if $\mathbf{F}(\hat{z}_{j})<0$ and $P_{i,l}==\mathbf{1}$ then
-                               $Mand\leftarrow Mand\cup\{N_{i,l}\}$   /* Neuron is in state $\mathbf{1}$, can’t coarsen */
-9
-10
-11
-12
+                               $Mand\leftarrow Mand\cup\{N_{i,l}\}$
+                                /* Neuron is in state $\mathbf{1}$, can’t coarsen */
+13
+14
+15
+16
 17      return $Mand$
+18
 Algorithm 4 Gradient_Search
@@
 Output: A minimal NAP specification $P$
-1 Function Sample_NAPs($unvisited$, $\theta$)
-       $P\leftarrow\dot{\mathcal{A}}(N)$   /* Use the coarsest NAP as a blank template */
+1
+2Function Sample_NAPs($unvisited$, $\theta$)
+       $P\leftarrow\dot{\mathcal{A}}(N)$
+        /* Use the coarsest NAP as a blank template */
 3       for $N_{i,l}$ in $unvisited$ do
-3             $rand\leftarrow\text{random}(0,1)$ if $rand\leq\theta$ then
-                   $P\leftarrow\widetilde{\Delta}(N_{i,l})$   /* Refine unvisited neurons using $\widetilde{\Delta}$ with probability $\theta$/
-4
-5
+4             $rand\leftarrow\text{random}(0,1)$
+5             if $rand\leq\theta$ then
+                   $P\leftarrow\widetilde{\Delta}(N_{i,l})$
+                    /* Refine unvisited neurons using $\widetilde{\Delta}$ with probability $\theta$/
+6
+7
 8      return $P$
-7Function Sample_Refine($visited$, $\theta$, $k$, $s$)
+9
+10
+11Function Sample_Refine($visited$, $\theta$, $k$, $s$)
+12
 13      while $|visited|<s\text{ and }\mathcal{V}(P)==0$  do
-             $unvisited\leftarrow N\setminus visited$ $ctr=dict()$   /* Create a counter for each neuron in $unvisited$/
-9             for _normal-_\__ in $range(k)$ do
-                   $P\leftarrow Sample\_NAPs(unvisited,\theta)$   /* Sample $k$ NAPs */
+14             $unvisited\leftarrow N\setminus visited$
+             $ctr=dict()$
+              /* Create a counter for each neuron in $unvisited$/
+15             for __\__ in $range(k)$ do
+                   $P\leftarrow Sample\_NAPs(unvisited,\theta)$
+                    /* Sample $k$ NAPs */
 16                   for $N_{i,l}$ in $unvisited$ do
 17                         if $P_{i,l}==\widetilde{\mathcal{A}}(N_{i,l})$ and $\mathcal{V}(P)==1$ then
-                               $ctr[N_{i,l}]$ += 1   /* Count successful NAPs for each neuron */
-12
-13
-14
-            $N_{mad}\leftarrow\operatorname*{argmax}_{N_{i,l}\in unvisited}\{ctr[N_{i,l}]\}$   /* Pick potential mandatory neuron */
+                               $ctr[N_{i,l}]$ += 1
+                                /* Count successful NAPs for each neuron */
+18
+19
+20
+            $N_{mad}\leftarrow\operatorname*{argmax}_{N_{i,l}\in unvisited}\{ctr[N_{i,l}]\}$
+              /* Pick potential mandatory neuron */
 21             $visited\leftarrow visited\cup\{N_{mad}\}$
+22
 23      return $\widetilde{\mathcal{A}}(visited)$
+24
+25
 […]
-       return $None$   /* Return None if the most refined NAP fails verification */
-18
+27 if $\mathcal{V}(P)==0$ then
+       return $None$
+        /* Return None if the most refined NAP fails verification */
+28
 29 else
@@
 32       else
-             $visited\leftarrow\emptyset$   /* Start from the coarsest NAP */
-23
+             $visited\leftarrow\emptyset$
+              /* Start from the coarsest NAP */
+33
 34      $Sample\_Refine(visited,\theta,k,s)$
+35
+36
 Algorithm 5 Sample_Refine
@@
 Output: A minimal NAP specification $P$
-1 Function Sample_Coarsen($mand\_neurons$, $\theta$, $s$)
-2       $P\leftarrow\widetilde{\mathcal{A}}(N)$; $mand\_neurons\leftarrow N$ if $\mathcal{V}(P)==0$ then
-             return $None$   /* Return None if the most refined NAP fails verification */
-3
+1
+2Function Sample_Coarsen($mand\_neurons$, $\theta$, $s$)
+3       $P\leftarrow\widetilde{\mathcal{A}}(N)$; $mand\_neurons\leftarrow N$
+4       if $\mathcal{V}(P)==0$ then
+             return $None$
+              /* Return None if the most refined NAP fails verification */
+5
 6       else
 7             while $|P|>s$ do
-6                   $P\leftarrow Sample\_NAPs(mand\_neurons,\theta)$ if $\mathcal{V}(P)==1$  then
-7                         $found\_neurons\leftarrow\emptyset$ for $N_{i,l}$ in $N$ do
+8                   $P\leftarrow Sample\_NAPs(mand\_neurons,\theta)$
+9                   if $\mathcal{V}(P)==1$  then
+10                         $found\_neurons\leftarrow\emptyset$
+11                         for $N_{i,l}$ in $N$ do
 12                               if $P_{i,l}==\widetilde{\mathcal{A}}(N_{i,l})$ then
 13                                     $found\_neurons\leftarrow found\_neurons\cup\{N_{i,l}\}$
-10
-                        $mand\_neurons\leftarrow found\_neurons$   /* Reduce search space */
-11
+14
+                        $mand\_neurons\leftarrow found\_neurons$
+                          /* Reduce search space */
+15
 16                   else
-                         $P\leftarrow Sample\_Naps(mand\_neurons,\theta)$   /* Sample a new NAP */
-13
-14
-            return $P$   /* Return the minimal NAP of size $s$/
-15
-16
+                         $P\leftarrow Sample\_Naps(mand\_neurons,\theta)$
+                          /* Sample a new NAP */
+17
+18
+            return $P$
+              /* Return the minimal NAP of size $s$/
+19
+20
+21
 Algorithm 6 Sample_Coarsen
@@
 We compute the most refined NAP specifications $\widetilde{P}^{0}$, $\widetilde{P}^{1}$, and $\widetilde{P}^{4}$ for labels 0, 1, and 4 using the $\widetilde{\mathcal{A}}$ function with a confidence ratio of $\delta=0.99$.
-The size of $\widetilde{P}^{0}$, $\widetilde{P}^{1}$, and $\widetilde{P}^{4}$ are 751, 745 and 712, respectively, which aligns with the results from (Geng et al., 2023).
+The size of $\widetilde{P}^{0}$, $\widetilde{P}^{1}$, and $\widetilde{P}^{4}$ are 751, 745 and 712, respectively, which aligns with the results from (Geng et al., 2023a).
 In contrast, the sizes of the minimal NAP specifications learned by the Coarsen algorithm for labels 0, 1, and 4 are significantly reduced to 480, 491, and 506, respectively.
@@
 Yet evidently, being able to verify the neural network on test data is an important task in real-world applications.
-The development of NAP specification (Geng et al., 2023) sheds light on this challenge.
+The development of NAP specification (Geng et al., 2023a) sheds light on this challenge.
 While previous work on NAP specifications proves its effectiveness, it assumes that all neurons contribute equally to the model’s robustness.

### v4

@@
-Learning Minimal NAP Specifications for Neural Network Verification
+Learning Minimal Neural Specifications
 Abstract.
-Specifications play a crucial role in neural network verification.
-They define the precise input regions we aim to verify, typically represented as L-infinity norm balls.
-While recent research suggests using neural activation patterns (NAPs) as specifications for verifying unseen test set data, it focuses on computing the most refined NAPs, often limited to very small regions in the input space.
-In this paper, we study the following problem: Given a neural network, find a minimal (coarsest) NAP that is sufficient for formal verification of the network’s robustness.
-Finding the minimal NAP specification not only expands verifiable bounds but also provides insights into which neurons contribute to the model’s robustness.
-To address this problem, we propose several exact and approximate approaches.
-Our exact approaches leverage the verification tool to find minimal NAP specifications in either a deterministic or statistical manner.
-Whereas the approximate methods efficiently estimate minimal NAPs using adversarial examples and local gradients, without making calls to the verification tool.
-This allows us to inspect potential causal links between neurons and the robustness of state-of-the-art neural networks, a task for which existing verification frameworks fail to scale.
-Our experimental results suggest that minimal NAP specifications require much smaller fractions of neurons compared to the most refined NAP specifications, yet they can significantly expand the verifiable boundaries to several orders of magnitude larger.
-Do, Not, Us, This, Code, Put, the, Correct, Terms, for, Your, Paper
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
+Formal verification is only as good as the specification of a system, which is also true for neural network verification.
+Existing specifications follow the paradigm of data as specification, where the local neighborhood around a reference data point is considered correct or robust.
+While these specifications provide a fair testbed for assessing model robustness, they are too restrictive for verifying any unseen test data points – a challenging task with significant real-world implications.
+Recent work shows great promise through a new paradigm, neural representation as specification, which uses neural activation patterns (NAPs) for this purpose.
+However, it computes the most refined NAPs, which include many redundant neurons.
+In this paper, we study the following problem: Given a neural network, find a minimal (general) NAP specification that is sufficient for formal verification of the network’s global robustness.
+Finding the minimal NAP specification not only expands verifiable bounds but also provides insights into which set of neurons contributes more to the model’s robustness.
+To address this problem, we propose three approaches—conservative, statistical, and optimistic—each offering distinct trade-offs between efficiency and performance.
+The first two rely on the verification tool to find minimal NAP specifications.
+The optimistic method efficiently estimates minimal NAPs using adversarial examples, without making calls to the verification tool until the very end.
+Each of these methods offers distinct strengths and trade-offs in terms of minimality and computational speed, making each approach suitable for scenarios with different priorities.
+The learnt minimal NAP specification allows us to inspect potential causal links between neurons and the robustness of state-of-the-art neural networks, a task for which existing work fails to scale.
+Our experimental results suggest that minimal NAP specifications require much smaller fractions of neurons compared to the NAP specifications computed by previous work, yet they can significantly expand the verifiable boundaries to several orders of magnitude larger.
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
 1. Introduction
@@
 Therefore, the verification of AI systems has become increasingly important and attracted much attention from the research community.
-The field of neural network verification largely follows the paradigm of software verification – using formal methods to verify desirable properties of systems using rigorous mathematical specifications and proofs (Wing, 1990).
-Most existing works (Katz et al., 2017, 2019; Huang et al., 2017a, 2020; Wang et al., 2021b) utilize $L_{\infty}$ ball specifications for verifying robustness guarantees, i.e., verify if neural networks consistently predict the same label within local neighborhoods ($L_{\infty}$ balls) of reference data points.
+The field of neural network verification largely follows the paradigm of software verification – using formal methods to verify desirable properties of systems through rigorous mathematical specifications and proofs (Wing, 1990).
+A notable trend in existing works (Katz et al., 2017b, 2019; Huang et al., 2017a, 2020; Wang et al., 2021b) focuses on scaling verification to larger and more complex neural networks.
+While scalability is undeniably important and requires the collective effort of the research community, this paper explores an orthogonal angle that has been largely overlooked: defining meaningful specifications.
+(a) Local neighborhood as specifications of the 3 reference points compared to using NAP as specification, NAP learnt from the 3 reference points and their local neighborhoods.
+Covers 4.23% and 7.64% of the WL region respectively.
+(b) NAP as specification, NAP learnt from the 3 reference points. Covers 10.61% of the WL region.
+(c) NAP as specification, NAP learnt from data points in the entire WL region. Covers 26.36% of the WL region.
+(d) NAP as specification, with NAP learnt from the entire WL region and coarsened. Covers 32.43% of the WL region.
 Figure 1.
-The comparison between the NAP specification (depicted by the green region) and the $L_{\infty}$ ball specification (represented by the gray region) in the MNIST dataset.
-Image $a$ is the reference image, and image $c$ is the closest image to $a$ with an $L_{\infty}$ distance of 0.5294.
-However, image $c$ cannot be verified using the $L_{\infty}$ ball specification because an adversarial example $b$ can be found at an $L_{\infty}$ distance of 0.2.
-In contrast, the NAP specification enables verification of inputs that are semantically close despite their geometric distance being far apart, such as image $c$.
-While this type of specification provides a fair and effective testbed for evaluating neural networks’ robustness, those verifiable $L_{\infty}$ balls are usually too tight, making them unlikely to cover unseen test set data (Geng et al., 2023a).
-For instance, the $L_{\infty}$ maximum verifiable bounds in the MNIST dataset is smaller than 0.2, whereas the smallest distance among images sharing the same label is greater than 0.5.
-Thus, narrowing the gap between verifiable regions and real data distribution to further enable verification in real-world settings poses a challenge.
+Illustration of verifiable regions of the Weak Left (WL) region of the ACAS_Xu (Katz et al., 2017a) advisories for a head-on encounter with $a_{\text{prev}}=\text{Clear of Conflict (COC)},\,\tau=0\text{s}$.
+The red bounding boxes represent the verifiable regions when using data as specifications, constrained by the reference points and their adversarial examples.
+The green hatched areas denote verifiable regions when using Neural Activation Patterns (NAP) as specifications.
+To illustrate, nearly all existing works follow a ”data as specification” paradigm, where the specification is defined by the consistency of local neighborhoods—often $L_{\infty}$ balls—around reference data points.
+While local neighborhood specification provides a fair and effective testbed for evaluating neural network robustness, it primarily covers a tiny convex region of input data that can mathematically be described by adding noise to the reference point, as illustrated by the red bounding box in Figure 1(a).
+This small region is constrained by it’s adversarial examples.
+It is too restrictive to adequately address a broader, unseen test set, which are real data sampled from the underlying distribution.
+Ideally, the specification should produce a verifiable region that includes all data points from the same class, where the data points may be distributed in a non-linear and non-convex manner within the input space.
+For instance, consider the safety-critical Airborne Collision Avoidance System for Unmanned Aircraft (ACAS Xu), the neural network processes an input representing the state of the aircraft and outputs one of five advisories: Clear of Conflict (COC), Weak Left (WL), Strong Left (SL), Weak Right (WR), or Strong Right (SR).
+Observe in 1 the the inputs belonging to Weak Left (WL) depicted in grey forms a irregular shape.
+This complexity underscores the need for flexible and robust specifications capable of accurately capturing such intricate distributions.
+Unlike local neighborhood specifications, manually defining such a specification is impractical.
+This poses a tricky chicken-egg problem: machine learning (ML) is necessary because it’s challenging to formally write down a precise definition (aka specification); but to be able to verify machine learning models, a formal specification would be needed.
+We argue that, in order to tackle this challenge, a separate learning algorithm for specifications is necessary.
+In this view, the “data as specification” paradigm is a simple but extremely overfitted algorithm for specification learning, which simply picks a small neighborhood of a reference data point in the input space.
 To this end, Geng et al.
-(2023a) proposes using neural activation patterns (NAPs) as specifications.
-Neural activation patterns can be viewed as an abstraction of each neuron’s state.
-For ReLU activation, it can be categorized into an activation state and a deactivation state.
-Since NAPs represent the intrinsic decision-making processes of neural networks, geologically distant data usually share the same NAPs if they belong to the same class.
-Compared to conventional local neighborhood specifications, NAP specifications cover much more flexible and potentially larger regions, as illustrated in Figure 1.
-This allows the verification of unseen real data rather than synthesized data from local neighbourhoods, paving the way for real-world application verification.
-However, it is noteworthy that the current approach to computing NAPs relies on a simple statistical method that assumes each neuron contributes to certifying the robustness behaviors of neural networks.
-Consequently, the computed NAPs are often the most refined ones.
-However, this is an overly restrictive assumption, as many studies (Frankle and Carbin, 2019; Lu et al., 2019; Liang et al., 2021; Wang et al., 2021a) have revealed that a non-trivial portion of neurons in neural networks may not play a substantial role.
-This motivates us to address the following challenge: Given a neural network, find a minimal (coarsest) NAP that is sufficient for formal verification of the network’s robustness.
-This problem is important for the following reasons: i) Minimal NAP specifications correspond to potentially much larger regions in the input space compared to the finest ones, this increases the potential of verification of unseen data; ii) Minimal NAP specifications provide insight into which neurons are responsible for the model making robust predictions, helping us to reveal the black-box nature of neural networks.
-To find the minimal NAP specifications, we introduce two basic algorithms: Refine and Coarsen.
-These algorithms exhaustively check all possible candidates using off-the-shelf verification tools, such as Marabou (Katz et al., 2019).
-While these approaches provide correctness guarantees, they are not efficient for verifying large neural networks as calls to the verification tool are usually very expensive.
-To tackle this inefficiency, we propose statistical versions of Refine and CoarsenSample_Refine and Sample_Coarsen that leverage sampling and statistical learning principles to find the minimal NAP specification in a more efficient manner.
-However, current verification tools cannot scale up to state-of-the-art neural network models, limiting the ability of verification-dependent approaches to compute the minimal NAP specification.
-This motivates us to study estimation methods that do not rely on verification tools.
-To this end, we find that adversarial examples and local gradients can provide valuable information about mandatory neurons, which are the essential building blocks of minimal NAPs.
-We then translate these findings into two approximate approaches: Adversarial_Prune and Gradient_Search.
-They can be used together to estimate mandatory neurons of state-of-art neural networks such as VGG-19 (Simonyan and Zisserman, 2014) and inspect possible links between models’ robustness performance and internal neurons, a task that existing verification frameworks struggle to scale for.
-Moreover, previous research has claimed that NAP specifications cover significantly larger regions compared to $L_{\infty}$ ball specifications, yet this assertion lacks empirical evidence.
-In this paper, we address this gap by introducing a simple approach to estimate the volume of regions corresponding to NAP specifications.
-Furthermore, this method provides us with a rough approximation of the volumetric change from refined NAP specifications to minimal NAP specifications.
-Our experimental results demonstrate significant volumetric changes, thereby underscoring the efficacy of minimal NAP specifications in real-world verification scenarios.
-We summarize our contributions as follows:
-  • We propose the problem of learning the minimal NAP specification and illustrate its importance in neural network verification.
-We present two simple approaches that leverage the verification tool with correctness guarantees.
-  • We introduce several key concepts related to the problem of learning minimal NAP specifications, including the abstraction function, NAP specification, and mandatory neurons.
-We demonstrate that the problem can be easily solved with access to mandatory neurons and propose two efficient estimation approaches to learn them.
-  • We propose two statistical learning approaches to the minimal NAP specification problem. They leverage sampling and estimation techniques as heuristics to achieve better efficiency.
-  • We present a simple approach for estimating the volume of regions corresponding to NAPs.
-This helps us understand the volumetric change in verifiable space between the minimal NAP specifications and the refined NAP specifications.
-  • Our experiments indicate that the minimal NAP specifications involve much smaller fractions of neurons compared to the most refined NAPs.
-And they significantly expand the verifiable bound by several orders of magnitude.
+(2023) proposes using a new and more promising paradigm - ”neural representation as specification”, which learns a specification in the representation space of the trained machine learning model in the form of neural activation patterns (NAPs).
+NAPs are value abstractions of hidden neurons which have been shown useful for understanding the inference of a model (Gopinath et al., 2018).
+Most importantly, a well-trained neural network would exhibit similar activation patterns for input data from the same class, regardless of their actual distance in the input space (Bengio et al., 2013; Tishby and Zaslavsky, 2015; Geng et al., 2023).
+This key observation suggests that if we learn a NAP—a common activation pattern shared by a certain class of data—it can serve as a candidate specification for verifying data from that class.
+Once such a NAP specification is successfully verified, we say any data covered by this NAP (exhibiting this pattern) provably belongs to the corresponding class.
+Ideally, if that NAP covers all data from a class, it can be considered a machine-checkable definition of that class.
+Geometrically speaking, compared to the $L_{\infty}$ ball specifications, NAP specifications provide greater flexibility, enabling coverage of larger and more irregular regions, This advantage is particularly notable when reference points are near the boundaries of the WL region as illustrated in Figure 1(a).
+In 1(b) we observe that the common NAP learned from the three reference points covers a significantly larger region compared to the combined coverage of the individual NAPs learned from each reference point independently.
+1(c) demonstrates even greater coverage when the NAP specification is learned from all inputs belonging to the WL class.
+However, it is noteworthy that the current approach to computing NAPs relies on a straightforward statistical method that assumes each neuron contributes to certifying the robustness of neural networks.
+Consequently, the computed NAPs are often overly refined.
+This is a restrictive assumption, as many studies (Frankle and Carbin, 2019; Lu et al., 2019; Liang et al., 2021; Wang et al., 2021a) have shown that a significant portion of neurons in neural networks may not play a substantial role.
+In the spirit of Occam’s Razor, we aim to systematically remove these redundant neurons that do not impact robustness.
+This motivates us to address the following challenge: given a neural network, how can we identify a minimal NAP (i.e., the coarsest abstraction) that is sufficient for verifying the network’s robustness?
+This problem is important for the following reasons: i) Minimal NAP specifications cover potentially much larger regions in the input space compared to the most refined ones, enhancing the ability to verify more unseen data; ii) Minimal NAP specifications provide insight into which neurons are responsible for the model’s robust predictions, helping to uncover the black-box nature of neural networks.
+For instance, if we aim to decode NAPs into human-understandable programs or rules (Li et al., 2023), minimal NAPs will always be easier to interpret than larger NAPs.
+We leave the interpretation of individual neurons as future work.
+To find the minimal NAP specifications, we first introduce a basic algorithm, Coarsen, which exhaustively checks all possible candidates using off-the-shelf verification tools, such as Marabou (Katz et al., 2019).
+Specifically, Coarsen gradually coarsens each neuron in the most refined NAPs, retaining only the coarsened neurons when Marabou confirms verification success.
+While this approach provides correctness guarantees, it is not efficient for verifying large neural networks, as calls to verification tools are typically expensive.
+To improve efficiency, we further propose statistical variants of Coarsen—namely, StochCoarsen, which leverages sampling and statistical learning principles to find minimal NAP specifications.
+1(d) highlights that coarsened NAP specifications, which involve fewer neurons, can cover an even larger portion of the WL region, highlighting the effectiveness of coarsened specifications in achieving broader coverage of the complex input space.
+However, verification-dependent approaches face challenges in scaling up to state-of-the-art neural network models due to the limitations of current verification tools.
+To address this, we develop an optimistic method, OptAdvPrune, which leverages adversarial examples to identify essential neurons—the fundamental building blocks of minimal NAPs.
+Our experimental results show it can efficiently estimate an initial starting point for minimal NAP specification and only making calls to the verification tool until the very end.
+We further apply these methods to state-of-the-art neural networks, such as VGG-19 (Simonyan and Zisserman, 2014).
+While formal verification of these estimated minimal NAP specifications remains impractical due to current scalability constraints in the underlying verification engine, we demonstrate that these NAPs capture significant hidden features and concepts learned by the model.
+As many studies suggest, visual interpretability and robustness are inherently related, emerging in learned features and representations (Alvarez Melis and Jaakkola, 2018; Boopathy et al., 2020).
+Consequently, we believe that the identified essential neurons contribute to the model’s robustness, and their activation states can serve as empirical indicators of confident predictions.
+Our contributions can be summarized as follows:
+  • We introduce the problem of learning minimal NAP specifications, emphasizing the need for a new paradigm in neural network specification.
+We present three approaches, each offering distinct trade-offs between efficiency and performance.
+  • We define key concepts such as the abstraction function, NAP specification, and essential neurons.
+We demonstrate that the problem reduces to identifying essential neurons and propose an optimistic approaches for estimating them.
+  • We propose a simple yet effective method for estimating the volume of the region mapped by the NAP.
+Our experiments show that minimal NAP specifications extend the verifiable bound by several orders of magnitude.
+  • We estimate essential neurons in the state-of-the-art VGG-19 network.
+Using a modified Grad-CAM map, we show that these essential neurons contribute to visual interpretability, providing strong evidence that they may also explain the model’s robustness.
 2. Background
@@
 2.1. Neural Networks for Classification Tasks
-In this paper, we focus on feed-forward ReLU neural networks.
-Generally speaking, a feed-forward network $N$ is comprised of $L$ layers, where each layer performs a linear transformation followed by a ReLU activation.
-We denote the pre-activation value and post-activation value at the $l$-th layer as $z^{(l)}(x)$ and $\hat{z}^{(l)}(x)$, respectively.
-The $l$-th layer computation is expressed as follows: $z^{(l)}(x)=\mathbf{W}^{(l)}\hat{z}^{(l-1)}(x)+\mathbf{b}^{(l)}$, $\hat{z}^{(l)}(x)=\mathbf{ReLu}(z^{(l)}(x))$, with $\mathbf{W}^{(l)}$ being the weight matrix and $\mathbf{b}^{(l)}$ representing the bias for the $l$-th layer.
-We denote the number of neurons in the $l$-th layer as $d_{l}$, and the $i$-th neuron in layer $l$ as $N_{i,l}$.
-The pre-activation value and post-activation value of $N_{i,l}$ at input $x$ is computed by $z_{i}^{(l)}(x)$ and $\hat{z}_{i}^{(l)}(x)$.
-The network can also be viewed as a function $\mathbf{F}^{<N>}:=\mathbb{R}^{d_{0}}\rightarrow\mathbb{R}^{d_{L}}$ such that $\mathbf{F}^{<N>}(x):=z^{(L)}(x)$, where $\mathbf{F}^{<N>}_{i}(x):=z_{i}^{(L)}(x)$ represents the output of $i$-th neuron in the last layer.
-We will omit $N$ when the context is clear.
-In a multi-class classification task with $C$ as the set of classes, when provided with an input $x$, the neural network predicts class/label $c\in C$ if the output of the $c$-th neuron $\textbf{F}_{c}(x)$ is the highest among the neurons in the $L$-th layer.
+In this paper, we focus on feed-forward neural networks equipped with ReLU activation functions.
+A feed-forward network $N$ consists of $L$ layers, where each layer alternates between a linear transformation and a ReLU activation.
+For the $l$-th layer, we denote the pre-activation and post-activation values as $z^{(l)}(x)$ and $\hat{z}^{(l)}(x)$, respectively.
+These are computed as follows: $z^{(l)}(x)=W^{(l)}\hat{z}^{(l-1)}(x)+b^{(l)},\quad\hat{z}^{(l)}(x)=ReLU(z^{(l)}(x))$, where $W^{(l)}$ and $b^{(l)}$ represent the weight matrix and bias vector of the $l$-th layer, respectively.
+Let $d_{l}$ denote the number of neurons in the $l$-th layer, and let $N_{i,l}$ refer to the $i$-th neuron in that layer.
+The pre-activation and post-activation values of $N_{i,l}$ for a given input $x$ are $z_{i}^{(l)}(x)$ and $\hat{z}_{i}^{(l)}(x)$, respectively.
+The network as a whole can be viewed as a function $F^{<N>}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{d_{L}}$, where $F^{<N>}(x)=z^{(L)}(x)$ represents the output of the final layer.
+For the $i$-th neuron in the last layer, its output is $F^{<N>}_{i}(x)=z_{i}^{(L)}(x)$.
+When the context is clear, we omit $N$ for simplicity and refer to the network as $F$.
+In multi-class classification tasks with $C$ classes, the network predicts the class $c\in C$ for a given input $x$ if the output $F_{c}(x)$ of the $c$-th neuron in the $L$-th layer is the largest among all outputs.
 2.2. Adversarial Attacks and Neural Networks Verification Problem
-Given a neural network $\mathbf{F}$ and a reference point $x$, adversarial attacks aim to search for a point $x^{\prime}$ that is geometrically close to the reference point $x$ such that $x^{\prime}$ and $x$ belong to different classes.
+Given a neural network $F$ and a reference point $x$, adversarial attacks aim to search for a point $x^{\prime}$ that is geometrically close to the reference point $x$ such that $x^{\prime}$ and $x$ belong to different classes.
 Here, we use the canonical specification, that is, we want to search in the local neighborhoods ($L_{\infty}$ norm balls) of $x$, formally denoted as $B(x,\epsilon):=\{x^{\prime}\mid||x-x^{\prime}||_{\infty}\leq\epsilon\}$, where $\epsilon$ is the radius.
 For certain $\epsilon$, given we know $x$ belongs to class $j$, we say an adversarial point is found if:
-(1) $\exists x^{\prime}\in B(x,\epsilon)\quad\exists i\in C\quad\mathbf{F}_{i}(x^{\prime})-\mathbf{F}_{j}(x)>0$
+(1) $\exists x^{\prime}\in B(x,\epsilon)\quad\exists i\in C\quad F_{i}(x^{\prime})-F_{j}(x^{\prime})>0$
 In practice, the change from the original data $x$ to adversarial data $x^{\prime}$ should be imperceptible, so they are more likely to be recognized as the same class/label from a human perspective.
@@
 This is different from the NAP specification, as we will discuss later.
-Neural networks have been shown to be vulnerable to adversarial attacks.
-Even imperceptible changes could easily alter the predicted outcome.
-This highlights the importance of neural network verification.
-The neural network verification problem can be seen as the flip side of adversarial attacks.
-When it can be formally proven that no adversarial point exists in $B(x,\epsilon)$, the neural network verification problem is considered solved.
-Formally, we want to verify:
-(2) $\forall x^{\prime}\in B(x,\epsilon)\quad\forall i\neq j\quad\mathbf{F}_{j}(x^{\prime})-\mathbf{F}_{i}(x^{\prime})>0$
-For a simpler presentation, we assume that $\mathbf{F}(x)$ is a binary classifier.
-For any given specification, $\mathbf{F}(x)\geq 0$ indicates that the model is verified; otherwise, we can find an adversarial example.
-Solving such a problem is known to be NP-hard (Katz et al., 2017), and current state-of-the-art approaches rely on specialized solvers.
-The difficulty escalates significantly as the number of neurons in neural networks increases, limiting successful verifications to only very small neural networks.
-Therefore, achieving scalability in verification remains an ongoing challenge.
+Neural networks are vulnerable to adversarial attacks, where even imperceptible changes can alter predictions significantly.
+This underscores the critical need for neural network verification, which can be viewed as highlighting safe regions that excludes adversarial regions.
+Solving the verification problem involves formally proving the absence of adversarial points in $B(x,\epsilon)$.
+Formally, the verification problem seeks to prove:
+(2) $\forall x^{\prime}\in B(x,\epsilon)\quad\forall i\neq j\quad F_{j}(x^{\prime})-F_{i}(x^{\prime})>0$
+For a simpler presentation, we assume that $F(x)$ is a binary classifier.
+For any given specification $B(x,\epsilon)$, $F(x)\geq 0$ indicates that the model is verified; otherwise, we can find an adversarial example.
+Solving such a verification problem is known to be NP-hard (Katz et al., 2017b), and achieving scalability in verification remains an ongoing challenge.
 2.3. Neuron Abstractions and Neural Activation Patterns
-To better discuss robustness verification using NAP specifications, we first introduce the relevant concepts of neuron abstractions, neuron abstraction functions, and neural activation patterns.
+To better discuss NAP specification and robustness verification, we first introduce the relevant concepts of neuron abstractions, neuron abstraction functions, and neural activation patterns.
 Neuron Abstractions
-[…]
+For a neural network $N$, an internal neuron $N_{i,l}$ (where $0\leq i\leq d_{l}$ and $11\leq l\leq L-$ 1) can have its post-activation value $\hat{z}_{i}^{(l)}(x)$ abstracted into finite states.
+This abstraction is a mapping from $\mathbb{R}$ to a set $\mathbb{S}=\{s_{1},s_{2},\ldots,s_{n}\}$, representing neuron states.
+A simple binary abstraction for ReLU activation defines two states: $s_{0}:=0$ (deactivation) and $s_{1}:=(0,\infty)$ (activation).
+Further, these states can be abstracted into a unary state $s_{*}=[0,\infty)$, covering the entire range of post-activation.
+This leads to a partial order: $s_{*}\preceq s_{0}$ and $s_{*}\preceq s_{1}$, where $s_{0}$ and $s_{1}$ refine $s_{*}$.
+For simplicity, we use $\mathbf{0},\mathbf{1},\mathbf{*}$ to represent these states.
+Our specification learning approach applies to any activation function definable within the abstraction domain, though this study focuses on ReLU.
 Definition 2.1 (Neuron Abstraction Function).
 Given a neural network $N$ and the abstraction state set $\mathbb{S}$.
-A neuron abstraction function is the mapping $\mathcal{A}:N\rightarrow\mathbb{S}$.
-Formally, for an arbitrary neuron $N_{i,l}$ where $0\leq i\leq d_{l}$ and $11\leq l\leq L-$ 1, the function abstracts $N_{i,l}$ to some state $s_{k}\in\mathbb{S}$, i.e., $\mathcal{A}(N_{i,l})=s_{k}$.
-The above characterization of neuron abstraction , it does not instruct us on how to perform binary abstraction in the absence of neuron values.
-Thus, we include the input(s) $x$ or $X$ as a new parameter in $\mathcal{A}$, and it is omitted when the context is clear.
-Here, we discuss three types of $\mathcal{A}$: the unary abstraction function function $\dot{\mathcal{A}}$, the binary abstraction function $\ddot{\mathcal{A}}$, and the statistical abstraction function $\widetilde{\mathcal{A}}$.
+A neuron abstraction function is a mapping $\mathcal{A}:N\rightarrow\mathbb{S}$.
+Formally, for an arbitrary neuron $N_{i,l}$, the function abstracts $N_{i,l}$ to some state $s_{k}\in\mathbb{S}$, i.e., $\mathcal{A}(N_{i,l})=s_{k}$.
+The above characterization of neuron abstraction does not instruct us on how to perform binary abstraction in the absence of neuron values.
+Therefore, we introduce two types of $\mathcal{A}$: unary $\dot{\mathcal{A}}$ and binary $\ddot{\mathcal{A}}$, both of which take a single input $x$ as a parameter, omitting it when the context is clear.
 The unary $\dot{\mathcal{A}}$ function
@@
 (4) $\ddot{\mathcal{A}}(N_{i,l},x)=\begin{cases}\mathbf{0}&\text{if }\hat{z}_{i}^{(l)}(x)=0\\ \mathbf{1}&\text{if }\hat{z}_{i}^{(l)}(x)>0\\ \end{cases}$
-The statistical $\widetilde{\mathcal{A}}$ function
-The binary $\ddot{\mathcal{A}}$ function is limited in calculating a single input, and meets a challenge when multiple inputs come into play, as two distinct inputs may lead to disagreements in a neuron’s abstraction state.
-We then approach this problem statistically by introducing the $\widetilde{\mathcal{A}}$ function defined as follows:
-(5) $\widetilde{\mathcal{A}}(N_{i,l},X)=\begin{cases}\mathbf{0}&\text{if }\frac{|\{x_{j}|\ddot{\mathcal{A}}(N_{i,l},x_{j})=\mathbf{0},\text{ }x_{j}\in X\}|}{|X|}\geq\delta\\ \mathbf{1}&\text{if }\frac{|\{x_{j}|\ddot{\mathcal{A}}(N_{i,l},x_{j})=\mathbf{1},\text{ }x_{j}\in X\}|}{|X|}\geq\delta\\ \mathbf{*}&\text{otherwise}\\ \end{cases}$
-where $\delta$ is a real number from $[0,1]$, and $X$ represents a set of inputs, i.e., $X:=\{x_{1},x_{2},...,x_{n}\}$.
-The parameter $\delta$ represents an error ratio to accommodate possible disagreements on neuron states between multiple inputs.
-Intuitively, the introduction of $\delta$ allows multiple inputs $x_{1},...,x_{n}$ to vote on a neuron’s state.
-For instance, when $\delta$ is set to 0.99, then 99% or more of the inputs must agree that a neuron is activated for the neuron to be in the $\mathbf{1}$ state.
-(a) A simple 2x2 fully connected neuron network.
+(a) A simple 2x2 fully connected neural network.
 (b) A subset of the NAP $\mathcal{P}$ family forms a binary tree.
@@
 Definition 2.2 (Neural Activation Pattern).
-Given a neural network $N$ and a neuron abstraction function $\mathcal{A}$.
-A neural activation pattern (NAP) $P$ is a tuple that consists of the abstraction state of all neurons in $N$.
-Formally, $P:=\langle\mathcal{A}(N_{i,l})\text{ }|\text{ }N_{i,l}\in N,\mathcal{A}\in\{\ddot{\mathcal{A}},\dot{\mathcal{A}}\}\rangle$, also denoted as $\mathcal{A}(N)$.
-The neuron $N_{i,l}$’s abstraction state specified by NAP $P$ is represented as $P_{i,l}$, i.e., $P_{i,l}=\mathcal{A}(N_{i,l})$.
-Partially ordered NAP
+Given a neural network $N$, an input $x$, and neuron abstraction functions $\ddot{\mathcal{A}}$ and $\dot{\mathcal{A}}$, a neural activation pattern (NAP) $P$ is a tuple representing the abstraction state of all neurons in $N$.
+Formally, $P:=\langle\mathcal{A}(N_{i,l},x)\mid N_{i,l}\in N\rangle$, where $\mathcal{A}$ is either $\ddot{\mathcal{A}}$ or $\dot{\mathcal{A}}$, chosen for each neuron.
+We also denote $P$ as $\mathcal{A}(N,x)$.
+The abstraction state of neuron $N_{i,l}$ is represented as $P_{i,l}$, specifically $P_{i,l}=\mathcal{A}(N_{i,l},x)$.
+When representing a refined activation pattern triggered by $x$ where $\ddot{\mathcal{A}}$ is applied to each neuron, it is expressed as $\ddot{\mathcal{A}}(N,x)$.
 We denote the power set of NAPs in $N$ as $\mathcal{P}$.
@@
 Moreover, two NAPs $P,P^{\prime}$ are equivalent if $P\preccurlyeq P^{\prime}$ and $P^{\prime}\preccurlyeq P$.
-To give a concrete example, Figure 2(b) depicts a subset of NAPs of a simple neural network consisting of 2 hidden layers and 4 neurons, as presented in Figure 2(a).
+[…]
+$\langle\mathbf{*},\mathbf{*},\mathbf{*},\mathbf{*}\rangle\preccurlyeq\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{*}\rangle\preccurlyeq\langle\mathbf{1},\mathbf{0},\mathbf{1},\mathbf{*}\rangle\preccurlyeq\langle\mathbf{1},\mathbf{0},\mathbf{1},\mathbf{0}\rangle.$
 […]
@@
 […]
-. This is because these NAPs sit on different branches formed by splitting certain ReLUs.
-Region Outlined by NAPs
-One key requirement for verification specification is its ability to represent a specific region within the input space.
-[…]
+. This occurs because these NAPs reside on different branches formed by splitting certain ReLUs.
 Figure 3.
@@
 More abstracted NAPs are formed by ignoring lines/hyperplanes created by neurons.
+Regions Outlined by NAPs
+[…]
 2.4. NAP Specifications for Robustness Verification
-This subsection illustrates how NAP specifications can be utilized for robustness verification.
-We discuss how the abstraction states of neurons can serve as signatures of distinct classes with the introduction of the class NAP definition.
+The previously introduced NAP definition primarily serves as a conceptual tool.
+To function as a specification for robustness verification, the regions outlined by NAPs ($R_{P}$) should cover a significant amount of data from a specific class.
+We demonstrate how to achieve this using the $\widetilde{\mathcal{A}}$ function.
+The statistical $\widetilde{\mathcal{A}}$ function
+The binary $\ddot{\mathcal{A}}$ function is limited to processing a single input and encounters challenges when multiple inputs are involved, as two distinct inputs may lead to conflicting abstraction states for a neuron; for instance, a neuron may be deactivated for input $x_{1}$ but activated for another input $x_{2}$.
+This can be problematic when attempting to learn a general NAP for an entire class of inputs.
+To address this issue, we approach it statistically by introducing the $\widetilde{\mathcal{A}}$ function, defined as follows:
+(6) $\widetilde{\mathcal{A}}(N_{i,l},X)=\begin{cases}\mathbf{0}&\text{if }\frac{|\{x_{j}|\ddot{\mathcal{A}}(N_{i,l},x_{j})=\mathbf{0},\text{ }x_{j}\in X\}|}{|X|}\geq\delta\\ \mathbf{1}&\text{if }\frac{|\{x_{j}|\ddot{\mathcal{A}}(N_{i,l},x_{j})=\mathbf{1},\text{ }x_{j}\in X\}|}{|X|}\geq\delta\\ \mathbf{*}&\text{otherwise}\\ \end{cases}$
+where $\delta$ is a real number from $[0,1]$, and $X$ represents a set of inputs, i.e., $X:=\{x_{1},x_{2},\ldots,x_{n}\}$.
+Since datasets often contain noisy data or challenging instances that the model cannot predict accurately, we introduce the parameter $\delta$ to accommodate standard classification settings in which Type I and Type II errors are non-negligible.
+Intuitively, the introduction of $\delta$ allows multiple inputs $x_{1},...,x_{n}$ to vote on a neuron’s state.
+For instance, when $\delta$ is set to 0.99, then 99% or more of the inputs must agree that a neuron is activated for the neuron to be in the $\mathbf{1}$ state.
+It is worth mentioning that the statistical $\widetilde{\mathcal{A}}$ function is equivalent to the method described in Geng et al.
+(2023).
 Definition 2.4 (Class NAP).
 In a classification task with the class set $C$, for any class $c\in C$, a class NAP $P^{c}$ is a NAP comprising abstract states outputted by an abstraction function given $X_{c}$, where $X_{c}$ denotes the set of inputs belonging to class $c$.
-Formally, $P^{c}:=\langle\mathcal{A}(N_{i,l},X_{c})\mid N_{i,l}\in N,\mathcal{A}\in\{\widetilde{\mathcal{A}},\dot{\mathcal{A}}\}\rangle$.
+Formally, $P^{c}:=\langle\mathcal{A}(N_{i,l},X_{c})\mid N_{i,l}\in N\rangle$, where $\mathcal{A}$ is either $\widetilde{\mathcal{A}}$ or $\dot{\mathcal{A}}$, chosen for each neuron.
 The power set of $P^{c}$ is denoted as $\mathcal{P}^{c}$.
-Recall that robustness verification can be thought of as proving that no adversarial examples exist in the local neighbourhoods of some reference point $x$.
-Intuitively, for NAP specifications, it is equivalent to show that no adversarial examples exist in the class NAP $P^{c}$; in other words, inputs exhibiting the NAP $P^{c}$ must be predicted as class $c$.
+Robustness verification involves proving that no adversarial examples exist in the local neighborhood of a reference point $x$.
+For NAP specifications, this translates to showing that no adversarial examples exist in class NAP $P^{c}$; in other words, inputs exhibiting $P^{c}$ must be classified as $c$.
 Geng et al.
-(2023a) argues that class NAPs must satisfy several essential requirements to qualify as NAP specifications.
-And finding qualified NAP specifications implies the underlying robustness problem is verified.
-We formally frame these requirements into the following properties:
+(2023) argues that class NAPs must meet several key requirements to qualify as NAP specifications, and finding such NAPs effectively verifies the underlying robustness problem.
+We formalize these requirements into the following properties:
 The non-ambiguity property
@@
 0.55-
-Pc2
+[…]
 […]
@@
 […]
-From a geological perspective, there must be no overlaps between class NAPs. In other words, it is also equivalent to verifying:
+From a geometric perspective, there must be no overlaps between class NAPs. In other words, it is also equivalent to verifying:
 $\forall c_{1},c_{2}\in C\text{ s.t. }c_{1}\neq c_{2}\quad R_{P^{c_{1}}}\bigcap R_{P^{c_{2}}}=\emptyset$
-We empirically observe that successful models, particularly neural networks that have achieved high accuracy in classification tasks, naturally exhibit this property.
-Because successful models tend to avoid confusion in predictions, their class NAPs are usually very distinct from each other.
 The NAP robustness property
-To serve as NAP specifications, class NAPs $P^{c}$ ensure that if an input exhibits it, i.e., $\mathcal{A}(N,x)\preccurlyeq P,\mathcal{A}\in\{\ddot{\mathcal{A}},\dot{\mathcal{A}}\}$, the input must be predicted as the corresponding class $c$.
+To serve as NAP specifications, class NAPs $P^{c}$ ensure that if an input exhibits it, i.e., $P\preccurlyeq\ddot{\mathcal{A}}(N,x)$, the input must be predicted as the corresponding class $c$.
 Formally, we have:
-(10) $\forall x\in R_{P^{c}}\quad\forall k\in C\text{ s.t. }k\neq c\quad\mathbf{F}_{c}(x)-\mathbf{F}_{k}(x)>0$
+(10) $\forall x\in R_{P^{c}}\quad\forall k\in C\text{ s.t. }k\neq c\quad F_{c}(x)-F_{k}(x)>0$
 in which
-(11) $R_{P^{c}}=\{x\mid{\mathcal{A}(N,x)\preccurlyeq P^{c}}\}$
-In contrast to canonical $L_{\infty}$ norm balls, class NAPs are more flexible in terms of size and shape.
-Additionally, there is no need to specify a reference point, since the locations of potential reference points are also encoded by class NAPs.
-However, it is possible that no class NAP $P^{c}$ in $\mathcal{P}^{c}$ satisfies this property.
-This can be mitigated by meeting the subsequent weaker property.
+(11) $R_{P^{c}}=\{x\mid{P^{c}\preccurlyeq\ddot{\mathcal{A}}(N,x)}\}$
 The NAP-augmented robustness property
-Instead of relying solely on class NAPs as specifications, local neighbours can still be employed in conjunction for verification.
-This hybrid form of specification has several advantages: 1) It narrows down the scope of verifiable regions when no class NAPs can meet the NAP robustness property; 2) NAP constraints essentially fix ReLU states, refining the search space for verification tools; 3) It focuses on the verification of valid test inputs rather than adversarial examples.
-Formally, this property can be stated as:
-(12) $\forall x^{\prime}\in B(x,\epsilon)\bigcap R_{P^{c}}\quad\forall k\in C\text{ s.t. }k\neq c\quad\mathbf{F}_{c}(x)-\mathbf{F}_{k}(x)>0$
+Rather than relying exclusively on class NAPs as specifications, local neighbors can also be used in combination for verification.
+This hybrid approach offers several advantages: 1) It narrows the scope of verifiable regions when class NAPs alone cannot satisfy the NAP robustness property; 2) NAP constraints essentially lock ReLU states, thereby refining the search space for verification tools; 3) It emphasizes verifying valid test inputs, while still covering broader and more flexible verifiable regions than those defined by $L_{\infty}$ norm ball specifications alone.
+Formally, this property can be expressed as:
+(12) $\forall x^{\prime}\in B(x,\epsilon)\bigcap R_{P^{c}}\quad\forall k\in C\text{ s.t. }k\neq c\quad F_{c}(x)-F_{k}(x)>0$
 in which
-(13) $B(x,\epsilon)=\{x^{\prime}\mid||x-x^{\prime}||_{\infty}\leq\epsilon\}\quad R_{P^{c}}=\{x^{\prime}\mid{\mathcal{A}(N,x^{\prime})\preccurlyeq P^{c}}\}$
+(13) $B(x,\epsilon)=\{x^{\prime}\mid||x-x^{\prime}||_{\infty}\leq\epsilon\}\quad R_{P^{c}}=\{x^{\prime}\mid{P^{c}\preccurlyeq\ddot{\mathcal{A}}(N,x^{\prime})}\}$
 To summarize, we state that a class NAP can serve as a NAP specification if it satisfies either the NAP robustness property or the NAP-augmented robustness property.
 Clearly, the former property is stronger, and it is possible that we can’t find a class NAP $P^{c}$ in $\mathcal{P}^{c}$ to satisfy this property.
-Fortunately, we can always find NAPs that satisfy the latter property by narrowing the verifiable region using additional $L_{\infty}$ norm ball specifications.
-Moreover, NAP specifications can be easily encoded with most of the state-of-the-art verification tools such as Marabou.
-In this paper, we will not dive into NAP encodings.
-Interested readers can refer to Geng et al.
-(2023a) for more details.
+Fortunately, we can always find NAPs that satisfy the latter property by narrowing the verifiable region with additional $L_{\infty}$ norm ball specifications.
 3. The minimal NAP specification problem
-In this section, we formulate the problem of learning minimal NAP specifications and present two naive approaches for solving this problem.
-Since our approaches require interactions with verification tools, we begin by introducing relevant notations describing the relationships between NAPs and verification tools.
-Figure 4.
-A DAG representing all possible NAPs, with green nodes signifying they can serve as a specification, while red nodes indicate they cannot.
-Within each node, a black circle means the corresponding neuron is abstracted by $\widetilde{\mathcal{A}}$, whereas the white circle means it is abstracted by $\dot{\mathcal{A}}$.
-Thus, the top node represents the coarsest NAP, while the bottom node entails the most refined one.
-Our goal is to find the NAP that resides in the green nodes while its size is minimal.
+In this section, we formulate the problem of learning minimal NAP specifications and present both a deterministic and statistical approach to address it.
+As our methods involve interactions with verification tools, we first introduce the relevant notations.
 3.1. Problem Formulation
@@
 For simplicity, we omit the superscript $c$ and refer to class NAPs simply as NAPs when the context is clear.
-We assume we have access to a verification tool, $\mathcal{V}:\mathcal{P}\to\{0,1\}$, which maps a class NAP $P\in\mathcal{P}$ to a binary set.
+We assume access to a verification tool, $\mathcal{V}:\mathcal{P}\to\{0,1\}$, which maps a class NAP $P\in\mathcal{P}$ to a binary set.
 Here, $\mathcal{V}(P)=1$ denotes a successful verification of the underlying robustness query, while 0 indicates the presence of an adversarial example.
@@
 It is not hard to see that $\mathcal{V}$ is monotone with respect to the NAP family $(\mathcal{P},\preccurlyeq)$.
-Given $P\preccurlyeq P^{\prime}$ and $\mathcal{V}(P^{\prime})=1$, it follows that $\mathcal{V}(P)=1$.
-However, given $P\preccurlyeq P^{\prime}$ and $\mathcal{V}(P)=1$, we cannot determine $\mathcal{V}(P^{\prime})$.
+Given $P^{\prime}\preccurlyeq P$ and $\mathcal{V}(P^{\prime})=1$, it follows that $\mathcal{V}(P)=1$.
+However, given $P^{\prime}\preccurlyeq P$ and $\mathcal{V}(P)=1$, we cannot determine $\mathcal{V}(P^{\prime})$.
 In other words, refining a NAP (by increasing the number of neurons abstracted to $\mathbf{0}$ or $\mathbf{1}$) can only enhance the likelihood of successful verification of the underlying robustness query.
 Definition 3.1 (The minimal NAP specification problem).
-Given a family of NAPs $\mathcal{P}$ and a verification tool $\mathcal{V}$, the minimal NAP specification problem is to find a NAP $P$ such that
-$\underset{P\in\mathcal{P},\mathcal{V}(P)=1}{\arg\min}|P|$
-where $|P|$ is defined as the size of $P$, representing the number of neurons that are abstracted to the $\mathbf{0}$ or $\mathbf{1}$, i.e., |$|\{N_{i,l}\mid P_{i,l}=\mathbf{0}\text{ or }\mathbf{1}\}|$.
-The (largest) minimal NAP specification’s size is denoted as $s$.
-When $P$ is minimal, it implies that for any NAP $P^{\prime}$ that is strictly more coarse than $P$, $\mathcal{V}(P^{\prime})=0$.
-Formally, this can be expressed as: $\forall P\preccurlyeq P^{\prime},P^{\prime}\neq P:\mathcal{V}(P^{\prime})=0$.
-Thus, there could exist multiple minimal NAP specifications.
-In such cases, we only need to choose one of them.
-On the other hand, it is possible that even the most refined NAPs cannot verify the robustness query.
-In such cases, we claim that no minimal NAP specifications exist.
-Additionally, since the computation using verification tools is usually expensive, we are interested in methods that efficiently find a minimal NAP specification, i.e., minimizing the number of calls to $\mathcal{V}$.
-We present two naive approaches to solving the problem. Since these approaches involve refining and coarsening the NAP, we will first formally define these two actions.
-Refine and Coarsen Action of NAP
-Recall our definition of NAP, the coarsest NAP is the one using $\dot{\mathcal{A}}$ to abstract each neuron in $N$.
-We denote this NAP as $\dot{P}:=\langle\dot{\mathcal{A}}(N_{i,l})\text{ }|\text{ }N_{i,l}\in N\rangle$.
-$\dot{P}$ is the smallest NAP with $|\dot{P}|=0$.
-In addition, we define the most refined NAP as the one that applies $\widetilde{\mathcal{A}}$ to abstract each neuron in $N$.
-We denote it as $\widetilde{P}:=\langle\widetilde{\mathcal{A}}(N_{i,l})\text{ }|\text{ }N_{i,l}\in N\rangle$.
-$\widetilde{P}$ is the largest NAP with size with $|\widetilde{P}|\leq|N|$.
-Clearly, $\widetilde{P}\preccurlyeq\dot{P}$.
-Given any NAP $P$, if we want to refine $P$ through a specific neuron $N_{i,l}$, we apply the $\widetilde{\mathcal{A}}$ function to $N_{i,l}$.
-We denote this refinement action as $\widetilde{\Delta}(N_{i,l})$.
-This action will either increase or leave unchanged $|P|$.
-Similarly, we denote the coarsen action as $\dot{\Delta}(N_{i,l})$.
-This action will either decrease or leave unchanged $|P|$.
-We present the semantics of these actions as follows:
-[…]
-$P\vdash\cdot\Downarrow P$
-[…]
-[…]
-[…]
-[…]
-where $P,P^{\prime},P^{\prime\prime}$ are from the NAP family $\mathcal{P}$, and $N$ is the underlying neural network.
-$\Delta_{1}$ and $\Delta_{2}$ are any two sequences of refine and coarsen actions.
-3.2. The Refine Approach
-Conceptually, the Refine approach iteratively increases the number of refined neurons in NAP $P$ until $\mathcal{V}(P)=1$, i.e., $P$ is able to prove the underlying robustness query.
-In other words, we gradually increase the size parameter $k$ and iterate over each NAP $P$ of size $k$ to check if $\mathcal{V}(P)=1$, as illustrated in Algorithm 1.
-To determine if a solution to the problem exists, we first check if the most refined NAP can succeed in verification.
-We proceed to iterative refinement only if $\mathcal{V}(\widetilde{P})=1$.
-However, the algorithm is not efficient and requires $12^{|N|}-$ 1 calls to $\mathcal{V}$ in the worst case, as proven in Theorem 3.2.
-Please refer to the proof in Appendix A.
-Therefore Refine is only practical when the search space of the NAP family $\mathcal{P}$ is small.
-Theorem 3.2.
-The algorithm Refine returns a minimal NAP specification with $\mathcal{O}(2^{|N|})$ calls to $\mathcal{V}$.
-3.3. The Coarsen Approach
-In contrast to the Refine approach, the Coarsen approach starts from the most refined NAP and then gradually coarsens it.
-We first check if the problem is well-defined by verifying if the most refined NAP $\widetilde{P}$ succeeds in verification.
-Then, for each neuron, we attempt to coarsen it using $\dot{\mathcal{A}}$; if the resulting NAP no longer verifies the query, we refine it back using $\widetilde{\mathcal{A}}$; otherwise, we keep the coarsened NAP.
-We describe the above procedure in Algorithm 2.
-The algorithm could require $|N|$ calls to $\mathcal{V}$ in the worst case, as proven in Theorem 3.3.
-Please refer to the proof in Appendix A.
-Theorem 3.3.
-The algorithm Coarsen returns a minimal NAP specification with $\mathcal{O}(|N|)$ calls to $\mathcal{V}$.
-1
-Input: The neural network $N$
-Output: A minimal NAP specification $P$
-2
-3Function Refine(N)
-4       $P\leftarrow\widetilde{\mathcal{A}}(N)$
-5       if $\mathcal{V}(P)==0$ then
-             return $None$ ;
-              /* Return None if even the most refined NAP fails verification */
-6
-7       else
-8             for $k$ from 1 to $|N|$ do
-                   $P\leftarrow\dot{\mathcal{A}}(N)$
-                    /* Refinement starts from the coarsest NAP $\dot{P}$/
-                   $K\_comb\leftarrow$ $Pick\_K\_Neurons(k,N)$
-                    /* Returns all combinations of size k */
-9                   for $comb\textbf{ in }K\_comb$ do
-10                         for $N_{i,l}\textbf{ in }comb$ do
-                               $P\leftarrow\widetilde{\Delta}(N_{i,l})$
-                                /* Refine neurons in the chosen combination */
-11
-12                        if $|\mathcal{V}(P)|==1$ then
-                               return $P$
-                                /* Found the minimal NAP */
-13
-14
-15
-16
-17
-Algorithm 1 refine
+Given a family of NAPs $\mathcal{P}$ and a verification tool $\mathcal{V}$, the minimal NAP specification problem is to find a NAP $P$ such that:
+$\forall P^{\prime}\preccurlyeq P,\,P^{\prime}\neq P\implies\mathcal{V}(P^{\prime})=0$
+That is, when $P$ is minimal, any NAP $P^{\prime}$ that is strictly coarser than $P$ will result in $\mathcal{V}(P^{\prime})=0$.
+The size of $P$, denoted by $|P|$ or $s$, represents the number of neurons abstracted to $\mathbf{0}$ or $\mathbf{1}$, i.e., |$|\{N_{i,l}\mid P_{i,l}=\mathbf{0}\text{ or }\mathbf{1}\}|$.
+It is important to note that ”minimal” refers to the level of abstraction, not the size of the NAP.
+Since $(\mathcal{P}^{c},\preccurlyeq)$ is a partially ordered set, multiple minimal NAP specifications may exist, rather than a single global minimum.
+In such cases, selecting any one of the minimal NAPs is sufficient.
+However, it is possible that even the most refined NAPs cannot verify the robustness query, in which case no minimal NAP specification exists.
+Given the computational cost of using verification tools, we are particularly interested in methods that efficiently find a minimal NAP specification, minimizing the number of calls to $\mathcal{V}$.
+3.2. Conservative Bottom Up Approach
+We introduce Coarsen, a method that starts with the most refined NAP and gradually coarsens it to learn minimal NAP specifications. Before proceeding, we define the relevant notations.
+Given a class input $X_{c}$, the coarsest NAP is defined as the one that uses $\dot{\mathcal{A}}$ to abstract each neuron in $N$, denoted as $\dot{P}:=\langle\dot{\mathcal{A}}(N_{i,l},X_{c})\mid N_{i,l}\in N\rangle$.
+This NAP is the smallest possible in size, with $|\dot{P}|=0$.
+Conversely, the most refined NAP applies $\widetilde{\mathcal{A}}$ to abstract each neuron, denoted as $\widetilde{P}:=\langle\widetilde{\mathcal{A}}(N_{i,l},X_{c})\mid N_{i,l}\in N\rangle$ or $\widetilde{\mathcal{A}}(N)$.
+This NAP is the largest in size, with $|\widetilde{P}|\leq|N|$.
+Clearly, we have $\dot{P}\preccurlyeq\widetilde{P}$.
+To refine any NAP $P$ through a specific neuron $N_{i,l}$, we apply the $\widetilde{\mathcal{A}}$ function, denoting this refinement as $\widetilde{\Delta}(N_{i,l})$.
+This action may either increase or leave $|P|$ unchanged.
+Conversely, we denote the coarsen action as $\dot{\Delta}(N_{i,l})$, which will either decrease or leave $|P|$ unchanged.
+The Coarsen approach starts by verifying if the most refined NAP $\widetilde{P}$ successfully passes verification.
+For each neuron, we attempt to coarsen it using $\dot{\mathcal{A}}$.
+If the resulting NAP no longer verifies the query, we revert to refining it back using $\widetilde{\mathcal{A}}$; otherwise, we retain the coarsened NAP.
+The procedure is detailed in Algorithm 1, which may require up to $|N|$ calls to $\mathcal{V}$ in the worst case, as shown in Theorem 3.2.
+Please refer to the proof in Appendix D.
 Input: The neural network $N$
@@
                    $P\leftarrow\dot{\Delta}(N_{i,l})$ ;
-                    /* Try to abstract this neuron using $\dot{\Delta}$/
+                    /* Try to coarsen(abstract) this neuron using $\dot{\Delta}$/
 7                   if $\mathcal{V}(P)==0$ then
@@
 Algorithm 1 Coarsen
-4. Verification Free Approaches
-While the Refine and Coarsen algorithms can find minimal NAP specifications with correctness guarantees, their inefficiency poses challenges with the verification of large neural networks.
-To address this issue, we introduce two efficient approaches for estimating minimal NAP specifications without requiring expensive calls to the verification tool.
-Unlike verification-based approaches, these estimation methods are deeply linked to mandatory neurons, a key concept in dissecting the minimal NAP specification problem, as discussed below.
-Definition 4.1 (Mandatory neuron).
-[…]
-$M=\{N_{i,l}\mid P_{i,l}\in\{\mathbf{0},\mathbf{1}\},P\textit{ is minimal}\}$
-Note that $M$ is the union of the set of mandatory neurons from all minimal NAP specifications. It follows that $|M|\geq s$, where $s$ denotes the size of the largest minimal NAP specification.
-The minimal NAP specification problem can be solved trivially if we gain access to $M$.
-Thus, our verification-free approaches are designed to determine mandatory neurons and estimate $M$.
-To better understand these our approaches, we first discuss the properties of mandatory neurons.
-Recall that verifying a robustness query given a NAP specification $P$ is equivalent to showing that $\mathbf{F}(x)\geq 0$ for input $x$ in region $R_{P}$.
-Thus, the necessary conditions of a mandatory neuron $N_{i,l}$ can be written as follows:
-  1. (1)
-    If $N_{i,l}$ is in state $\mathbf{0}$, it implies when $\hat{z}_{i}^{(l)}(x)=0,\mathbf{F}(x)\geq 0$. In addition, $\exists x\text{ s.t. }\hat{z}_{i}^{(l)}(x)>0,\mathbf{F}(x)<0$.
-  2. (2)
-    If $N_{i,l}$ is in state $\mathbf{1}$, it implies $\forall x\text{ s.t. }\hat{z}_{i}^{(l)}(x)>0,\mathbf{F}(x)\geq 0$. In addition, when $\hat{z}_{i}^{(l)}(x)=0,\mathbf{F}(x)<0$.
-[…]
-4.1. Adversarial-guided Prune
-We first introduce Adversarial_Prune to identify mandatory neurons. Intuitively, it attempts to show a neuron $N_{i,l}$ is mandatory by providing an adversarial example $x^{\prime}$ exists.
-When an adversarial example $x^{\prime}$ is found, it immediately indicates that the NAP $\ddot{\mathcal{A}}(N,x^{\prime})$ fails the verification, i.e., $\mathcal{V}(\ddot{\mathcal{A}}(N,x^{\prime}))=0$.
-Moreover, it also implies that any NAP subsumed by $\ddot{\mathcal{A}}(N,x^{\prime})$ fails verification.
-For instance, suppose an adversarial example $x^{\prime}$ is found for a simple one-layer four-neuron neural network and $\ddot{\mathcal{A}}(N,x^{\prime})$ is $\langle\mathbf{1},\mathbf{0},\mathbf{1},\mathbf{0}\rangle$.
-We can infer that NAPs like $\langle\mathbf{1},\mathbf{0},\mathbf{1},\mathbf{*}\rangle$, $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{*}\rangle$, $\langle\mathbf{1},\mathbf{*},\mathbf{*},\mathbf{0}\rangle$, and $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle$ fail the verification.
-This information is particularly useful when determining if a neuron is mandatory.
-For example, if we know that NAP $P:=\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{1}\rangle$ is a specification, i.e., $\mathcal{V}(P)=1$, then we can easily deduce that the fourth neuron $N_{4,1}$ is mandatory.
-This is because that coarsening the fourth neuron would expand $P$ to $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{*}\rangle$, which would include the adversarial example $x^{\prime}$ and thus fail the verification, as illustrated in Figure 5(a).
-It is evident that the neuron where $P$ and $\ddot{\mathcal{A}}(N,x^{\prime})$ disagrees must be mandatory.
-However, when the two NAPs disagree on multiple neurons, things become a little bit different.
-Suppose the NAP specification $P$ is $\langle\mathbf{1},\mathbf{1},\mathbf{*},\mathbf{1}\rangle$, i.e., $\mathcal{V}(\langle\mathbf{1},\mathbf{1},\mathbf{*},\mathbf{1}\rangle)=1$.
-We know $\mathcal{V}(\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle)=0$ by the adversarial example $x^{\prime}$.
-In this case, if we coarsen the second and fourth neurons, $N_{2,1}$ and $N_{4,1}$, the NAP specification will expand to $\langle\mathbf{1},\mathbf{*},\mathbf{*},\mathbf{*}\rangle$, which will cover the $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle$, thus failing the verification.
-In this case, $N_{2,1}$ and $N_{4,1}$ could both be mandatory neurons or either one of them is mandatory, as illustrated in Figures 5(b), 5(c), 5(d).
-So, we simply let $\{N_{2,1},N_{4,1}\}$ be the upper bound of mandatory neurons (learned from $x^{\prime}$).
-Formerly, given a NAP $P$, we say a neuron $N_{i,l}$ is in the upper bound of mandatory neurons $M$ if satisfies the following condition:
-  1. (1)
-    $N_{i,l}$ must be in the binary states, i.e., $P_{i,l}\in\{\mathbf{0},\mathbf{1}\}$
-  2. (2)
-    There exists $x^{\prime}$ such that $\ddot{\mathcal{A}}(N_{i,l},x^{\prime})$ XORs with $P_{i,l}$, i.e., $\exists x^{\prime}\text{ such that }\ddot{\mathcal{A}}(N_{i,l},x^{\prime})\oplus P_{i,l}=1$
-(a) $N_{4,1}$ is mandatory.
-(b) $N_{4,1},N_{2,1}$ are mandatory.
-(c) $N_{4,1}$ is mandatory.
-(d) $N_{2,1}$ is mandatory.
-Figure 5.
-Geometric interpretation of NAPs on mandatory neurons.
-The first subfigure represents the case when the two NAPs disagree on $N_{4,1}$.
-The other three subfigures represent the three cases when the two NAPs disagree on $N_{4,1},N_{2,1}$.
-Regions colored green pass verification, whereas red indicates an adversarial example exists.
-Here, we omit states for $N_{1,1},N_{3,1}$ in those NAPs for simplicity.
-The field of adversarial attacks has been extensively studied, offering a wealth of methods that we can leverage.
-These approaches are usually computationally efficient, making it easy to access a large collection of adversarial examples.
-Therefore, we can compute the upper bound of mandatory neurons efficiently by simply taking the union of upper bounds learned for each adversarial example, as illustrated in Algorithm 3.
-Input: The neural network $N$, the input dataset $X$
-Output: A collection of mandatory neurons
-1
-2Function Adversarial_Prune($N$, $X$)
-3       $Mand\leftarrow\emptyset$ ; $P\leftarrow\widetilde{\mathcal{A}}(N)$
-4       for $x_{j}$ in $X$ do
-5             $x^{\prime}_{j}\leftarrow Adversarial\_Attack(x_{j})$
-6             for $N_{i,l}$ in $N$ do
-7                   if $P_{i,l}\in\{\mathbf{0},\mathbf{1}\}$ and $\ddot{\mathcal{A}}(N_{i,l},x^{\prime}_{j})\oplus P_{i,l}$ then
-                         $Mand\leftarrow Mand\cup\{N_{i,l}\}$
-                          /* Create mandatory neurons for $x^{\prime}_{j}$/
-8
-9
-10      return $Mand$
-11
-Algorithm 3 Adversarial_Prune
-4.2. Gradient-guided Search
-We introduce another approach called Gradient_Search to identify mandatory neurons.
-Similar to Adversarial_Prune, Gradient_Search avoids costly interactions with the verification tool.
-Instead, it leverages gradient estimations to analyze the structure of $\mathbf{F}$.
-Recall our definition of a neural network in Section 2.1, where the output $\mathbf{F}(x)$ can be seen as a function of the post-activation value of any internal neuron $\hat{z}_{i}^{(l)}$.
-We denote this function as $\mathbf{F}(\hat{z}(x))$, omitting $i$ and $l$ for simplicity.
-The function $\mathbf{F}(\hat{z})$ is non-linear and operates within an input range of $\hat{z}\in[0,+\infty)$.
-It is worth noting that the gradient $\frac{\partial\mathbf{F}}{\partial\hat{z}}$ can provide valuable insights into the local structure of $\mathbf{F}$.
-This insight may help determine the necessity condition of mandatory neurons by checking if there exists a $\hat{z}$ such that $\mathbf{F}(\hat{z})<0$ for the corresponding neuron.
-To be more specific, given some sampled data $\hat{z}_{1},\ldots,\hat{z}_{n}$, we can compute their corresponding function values $\mathbf{F}(\hat{z}_{1}),\ldots,\mathbf{F}(\hat{z}_{n})$ along with their gradients $\frac{\partial\mathbf{F}}{\partial\hat{z}}|_{\hat{z}_{1}},\ldots,\frac{\partial\mathbf{F}}{\partial\hat{z}}|_{\hat{z}_{n}}$.
-Then if there exists a $\hat{z}_{j}$ such that $|\mathbf{F}(\hat{z}_{j})|$ is sufficiently small, yet the norm of its gradient $|\frac{\partial\mathbf{F}}{\partial\hat{z}}|_{\hat{z}_{j}}|$ is significantly large, then it is very likely that $\mathbf{F}$ will go below zero nearby, as illustrated in Figure 6.
-Then, if we happen to know that the corresponding neuron $N_{i,l}$ is in state $\mathbf{0}$, it is highly likely that $N_{i,l}$ is mandatory.
-Conversely, if a neuron $N_{i,l}$ is in state $\mathbf{1}$, and we sample an $x$ such that $\hat{z}_{i}^{(l)}=0$ and $\mathbf{F}(x)<0$, we immediately recognize that the neuron is mandatory.
-Algorithm 4 presents pseudocode for estimating whether a neuron is mandatory based on the discussions above.
-Figure 6.
-The gradient provides insights into the local structure of $\mathbf{F}$.
-When the gradient is large and the function value is small, it strongly suggests that $\mathbf{F}$ will go below zero nearby.
-Input: The neural network $N$, the input dataset $X$, and thresholds $\beta$, $\gamma$
-Output: A collection of mandatory neurons
-1
-2Function Gradient_Search($N$, $X$, $\beta$, $\gamma$)
-3       $Mand\leftarrow\emptyset$ ; $P\leftarrow\widetilde{\mathcal{A}}(N)$
-4       for $N_{i,l}$ in $N$ do
-5             for $x_{j}$ in $X$ do
-6                   $\hat{z}_{j}\leftarrow\hat{z}_{i}^{(l)}(x_{j})$
-7                   if $\hat{z}_{j}>0$  then
-8                         if  $|\mathbf{F}(\hat{z}_{j})|<\beta$ and $|\frac{\partial\mathbf{F}}{\partial\hat{z}_{i}^{(l)}}|_{\hat{z}_{j}}|>\gamma$ and $P_{i,l}==\mathbf{0}$ then
-                               $Mand\leftarrow Mand\cup\{N_{i,l}\}$
-                                /* Neuron is in state $\mathbf{0}$, can’t coarsen */
-9
-10
-11                  else
-12                         if $\mathbf{F}(\hat{z}_{j})<0$ and $P_{i,l}==\mathbf{1}$ then
-                               $Mand\leftarrow Mand\cup\{N_{i,l}\}$
-                                /* Neuron is in state $\mathbf{1}$, can’t coarsen */
-13
-14
-15
-16
-17      return $Mand$
-18
-Algorithm 4 Gradient_Search
-Both the estimation method Gradient_Search and Adversarial_Prune can be used together.
-Taking the intersection of their estimated results may provide us with a better overall estimate.
-In addition, the estimated mandatory neurons can serve as an initial starting point in our statistical versions of the Refine and Coarsen approaches introduced in Section 5.
-Our experimental results regarding these two estimation approaches indicate a notable trend: neurons from deeper layers are more likely to be mandatory.
-This observation aligns with commonly-held beliefs that deeper neurons are generally linked to high-level feature representation, and thus play a more important role than shallow neurons in final classification decisions.
-This observation leads us to a valuable heuristic: prioritizing deeper neurons over shallower ones as we iterate through neurons in our implementations of the Refine and Coarsen algorithms.
-5. Statistical Learning Approaches
-In this section, we introduce two new algorithms that address the minimal NAP specification problem.
-While they share the same overarching concept as refine and coarsen, these methods employ sampling and statistical learning principles to efficiently learn a minimal NAP specification.
-Recall that the set of mandatory neurons $M$ is the union of all neurons that appears in any minimal specification and is an upper bound of the largest minimal specification.
-5.1. The Statistical Refine Approach
-Mandatory neurons are crucial for forming NAP specifications, as their binary states play a critical role in determining the neural network’s robustness performance.
-We leverages this property to find mandatory neurons statistically.
-To be more specific, suppose we sample some NAPs $P_{1},P_{2},\ldots,P_{n}$ from the NAP family $\mathcal{P}$.
-For those NAPs that qualify as specifications (i.e., $\mathcal{V}(P)=1$), mandatory neurons should appear more frequently in them than in those NAPs that fail the verification tool (i.e., $\mathcal{V}(P)=0$).
-Based on this insight, we propose an approach called Sample_Refine that relies on non-repetitive sampling to identify mandatory neurons for solving the minimal NAP problem.
-We start with the coarsest NAP and iteratively collect the most probable mandatory neurons.
-In every iteration, we sample $k$ NAPs by refining unvisited neurons with some probability $\theta$.
-The sampled NAPs are then fed to the verification tool, and the neuron that appears most frequently in verifiable NAPs is the most probable mandatory neuron in this iteration.
-The neuron is marked as visited, and the process stops when either we collect $s$ neurons (assuming that $s$ is known), or the current mandatory neurons form a NAP specification $P$, i.e., $\mathcal{V}(P)=1$.
-Finally, we return the learned NAP, obtained by applying $\widetilde{\mathcal{A}}$ to the collected neurons.
-The algorithm 5 provides an overview of the above procedure.
-It is worth noting that Sample_Refine doesn’t guarantee correctness, as we may end up collecting only the $s$ most probable mandatory neurons, which may not be sufficient to form a specification.
-Another concern regarding this algorithm is sampling efficiency, specifically the potential for the number of samples required to grow exponentially with the size of the minimal NAP specification $s$.
-To understand why this is problematic, consider a scenario where the only minimal NAP specification $P$ consists of all mandatory neurons; in this case, all $|M|$ neurons must be selected for $P$ to be learned.
-If $\theta$ is set to a constant value, then the expected number of samples needed to obtain the NAP specification is $(\frac{1}{\theta})^{|M|}$.
-To address this, we set $\theta=\left(\frac{|M|}{|M|+1}\right)^{|M|}$.
-This choice ensures that the sampling efficiency is polynomial in both $|M|$ and $s$, as proven in Theorem 5.1.
-Please refer to the proof in the Appendix B.
-In addition, Theorem 5.1 also shows that with high probability, a mandatory neuron will be found with $O(log|N|)$ calls to $\mathcal{V}$.
 Theorem 3.2.
-With probability $\theta=|(\frac{|M|}{|M|+1})|^{|M|}$, Sample_Refine has $1-\delta$ probability of outputting a minimal NAP specification with $\Theta\left(|M|^{2}(\log|N|+\log(s/\delta))\right)$ examples each iteration and $O(s|M|^{2}(\log|N|+\log(s/\delta)))$ total calls to $\mathcal{V}$.
-5.2. The Statistical Coarsen Approach
-The Coarsen algorithm begins with the most refined NAP and progressively coarsens each neuron until the verification process fails.
+The algorithm Coarsen returns a minimal NAP specification with $\mathcal{O}(|N|)$ calls to $\mathcal{V}$.
+3.3. Statistical Coarsen Approach
+The Coarsen algorithm initiates with the most refined NAP and progressively coarsens each neuron until verification fails.
 Enhancing the algorithm’s performance is possible by coarsening multiple neurons during each iteration.
 However, a fundamental question emerges: How do we determine which set of neurons to coarsen in each round?
-We present Sample_Coarsen to answer this question.
+We present StochCoarsen to answer this question.
 In this approach, we assume that each neuron is independent of the others and select neurons to coarsen in a statistical manner.
@@
 We repeat this process until the NAP size reaches $s$.
-Algorithm 6 provides the pseudocode for Sample_Coarsen.
-Input: The neural network $N$, the probability $\theta$, sample size $k$, and the size $s$
-Output: A minimal NAP specification $P$
-1
-2Function Sample_NAPs($unvisited$, $\theta$)
-       $P\leftarrow\dot{\mathcal{A}}(N)$
-        /* Use the coarsest NAP as a blank template */
-3       for $N_{i,l}$ in $unvisited$ do
-4             $rand\leftarrow\text{random}(0,1)$
-5             if $rand\leq\theta$ then
-                   $P\leftarrow\widetilde{\Delta}(N_{i,l})$
-                    /* Refine unvisited neurons using $\widetilde{\Delta}$ with probability $\theta$/
-6
-7
-8      return $P$
-9
-10
-11Function Sample_Refine($visited$, $\theta$, $k$, $s$)
-12
-13      while $|visited|<s\text{ and }\mathcal{V}(P)==0$  do
-14             $unvisited\leftarrow N\setminus visited$
-             $ctr=dict()$
-              /* Create a counter for each neuron in $unvisited$/
-15             for __\__ in $range(k)$ do
-                   $P\leftarrow Sample\_NAPs(unvisited,\theta)$
-                    /* Sample $k$ NAPs */
-16                   for $N_{i,l}$ in $unvisited$ do
-17                         if $P_{i,l}==\widetilde{\mathcal{A}}(N_{i,l})$ and $\mathcal{V}(P)==1$ then
-                               $ctr[N_{i,l}]$ += 1
-                                /* Count successful NAPs for each neuron */
-18
-19
-20
-            $N_{mad}\leftarrow\operatorname*{argmax}_{N_{i,l}\in unvisited}\{ctr[N_{i,l}]\}$
-              /* Pick potential mandatory neuron */
-21             $visited\leftarrow visited\cup\{N_{mad}\}$
-22
-23      return $\widetilde{\mathcal{A}}(visited)$
-24
-25
-[…]
-27 if $\mathcal{V}(P)==0$ then
-       return $None$
-        /* Return None if the most refined NAP fails verification */
-28
-29 else
-30       if heuristics then
-31             $visited\leftarrow Gradient\_Search(N)\cap Adversarial\_Prune(N)$
-32       else
-             $visited\leftarrow\emptyset$
-              /* Start from the coarsest NAP */
-33
-34      $Sample\_Refine(visited,\theta,k,s)$
-35
-36
-Algorithm 5 Sample_Refine
-Similar to Sample_Refine, Sample_Coarsen also faces the same challenge related to sample efficiency.
-To better illustrate this, suppose a minimal NAP specification $P$ with size $s$ can be found after one iteration of refinement.
-Then, the probability of selecting the exact number of $s$ mandatory neurons in $P$ is $\theta^{s}$.
-Consequently, if $\theta$ is set to a constant value, the expected number of samples needed to find the NAP is $(\frac{1}{\theta})^{s}$.
-Unlike Sample_Refine, we know the refinement will always ensure the resulting NAP can pass verification.
-So, once such a NAP is learned, we can narrow down the estimated mandatory neurons by an expected factor of $\theta$.
-In this way, the expected number of samples and the expected number of iterations are inversely related, yet their product is the total number of calls to $\mathcal{V}$.
-We demonstrate that by setting $\theta=e^{-\frac{1}{s}}$, we not only make the expected number of samples polynomial in $s$, but also minimize the total number of calls to $\mathcal{V}$, as proven in Theorem 5.2.
-Please refer to the proof in Appendix B.
-Theorem 5.2.
-With probability $\theta$=$e^{-\frac{1}{s}}$, Sample_Coarsen learns a minimal NAP specification with $\mathcal{O}(slog|N|)$ calls to $\mathcal{V}$.
+This method can be regarded as optimistic bottom up.
+Algorithm 2 provides the pseudocode for StochCoarsen.
+It’s important to note that the stochastic manner of this algorithm faces the challenge of sample efficiency.
+For instance, if a minimal NAP specification $P$ of size $s$ is identified after one iteration, the probability of selecting the exact $s$ essential neurons in $P$ is $\theta^{s}$.
+Consequently, if we set $\theta$ as a constant, the expected number of samples required to find the NAP becomes $(\frac{1}{\theta})^{s}$.
+StochCoarsen allows us to narrow down the estimated essential neurons by an expected factor of $\theta$ once a valid NAP is learned.
+Thus, the expected number of samples and iterations are inversely related, with their product equating to the total calls to $\mathcal{V}$.
+By setting $\theta=e^{-\frac{1}{s}}$, we can achieve polynomial expected samples in $s$ while minimizing the total number of calls to $\mathcal{V}$, as proven in Theorem 3.3.
+The proof is detailed in Appendix E.
 Input: The neural network $N$, the probability $\theta$, and the size $s$
@@
 1
-2Function Sample_Coarsen($mand\_neurons$, $\theta$, $s$)
+2Function StochCoarsen($mand\_neurons$, $\theta$, $s$)
 3       $P\leftarrow\widetilde{\mathcal{A}}(N)$; $mand\_neurons\leftarrow N$
@@
 20
-21
-Algorithm 6 Sample_Coarsen
+Algorithm 2 StochCoarsen
+Theorem 3.3.
+With probability $\theta$=$e^{-\frac{1}{s}}$, StochCoarsen learns a minimal NAP specification with $\mathcal{O}(slog|N|)$ calls to $\mathcal{V}$.
 Setting the sample probability $\theta$
-Setting $s$ poses a challenge in practice, as we assume that $s$ is always provided in Sample_Refine and Sample_Coarsen.
+Setting $s$ poses a challenge in practice, as we assume that $s$ is always provided in StochCoarsen.
 However, this can be addressed by dynamically updating $\theta$ based on the result of $\mathcal{V}(P)$ (Liang et al., 2011).
-With $\theta$ from theorem 5.2, Sample_Coarsen finds a NAP specification with probability $\left(e^{-1/s}\right)^{s}=e^{-1}$.
-Recall that theorem 5.1 state that Sample_Refine finds a NAP specification with probability $\left(\frac{|M|}{|M|+1}\right)^{|M|}$.
-The probability approaches $e^{-1}$ asymptotically.
+With $\theta$ from theorem 3.3, StochCoarsen finds a NAP specification with probability $\left(e^{-1/s}\right)^{s}=e^{-1}$.
 Thus, we aim to set $\theta$ such that the $Pr(\mathcal{V}(P)=1)=e^{-1}$.
-Intuitively, if a sampled NAP $P$ is a specification, we decrease $\theta$ so less neurons will be refined or more neurons will be coarsened.
+Intuitively, if a sampled NAP $P$ is a specification, we decrease $\theta$ so more neurons will be coarsened.
 Similarly, if $P$ is not a specification, $\theta$ needs to be increased.
@@
 (16) $\lambda\leftarrow\lambda-\eta(\mathcal{V}(P)-e^{-1})$
+4. Optimistic Approach
+While Coarsen and StochCoarsen effectively identify minimal NAP specifications with correctness guarantees, their practicality is limited by the extensive runtime required to perform a large number of expensive verification calls.
+This limitation makes them less suitable for time-critical applications or large-scale neural networks.
+To address these challenges, we propose a optimistic approach for efficiently initializing NAP specifications.
+This approach significantly reduces the number of subsequent verification calls and provide a reliable upper bound on the size of the minimal NAP specification, improving both efficiency and scalability.
+The approach focuses on the concept of essential neurons, key to understanding the minimal NAP specification problem.
+Definition 4.1 (Essential neuron).
+[…]
+$E=\{N_{i,l}\mid P_{i,l}\in\{\mathbf{0},\mathbf{1}\},P\textit{ is minimal}\}$
+Note that $E$ is the union of the set of essential neurons from all minimal NAP specifications. It follows that $|E|\geq s$, where $s$ denotes the size of the largest minimal NAP specification.
+The minimal NAP specification problem can be solved trivially if we gain access to $E$.
+Thus, our optimistic approach is designed to determine essential neurons and estimate $E$.
+To better understand our approach, we first discuss the properties of essential neurons.
+Recall that verifying a robustness query given a NAP specification $P$ is equivalent to showing that $\mathbf{F}(x)\geq 0$ for input $x$ in region $R_{P}$.
+Thus, the necessary conditions of a essential neuron $N_{i,l}$ can be written as follows:
+  1. (1)
+    If $N_{i,l}$ is in state $\mathbf{0}$, it implies when $\hat{z}_{i}^{(l)}(x)=0,\mathbf{F}(x)\geq 0$. In addition, $\exists x\text{ s.t. }\hat{z}_{i}^{(l)}(x)>0,\mathbf{F}(x)<0$.
+  2. (2)
+    If $N_{i,l}$ is in state $\mathbf{1}$, it implies $\forall x\text{ s.t. }\hat{z}_{i}^{(l)}(x)>0,\mathbf{F}(x)\geq 0$. In addition, when $\hat{z}_{i}^{(l)}(x)=0,\mathbf{F}(x)<0$.
+(a) $N_{4,1}$ is essential.
+(b) $N_{4,1},N_{2,1}$ are essential.
+(c) $N_{4,1}$ is essential.
+(d) $N_{2,1}$ is essential.
+Figure 4.
+Geometric interpretation of NAPs on essential neurons.
+The first subfigure represents the case when the two NAPs disagree on $N_{4,1}$.
+The other three subfigures represent the three cases when the two NAPs disagree on $N_{4,1},N_{2,1}$.
+Regions colored green pass verification, whereas red indicates an adversarial example exists.
+Here, we omit states for $N_{1,1},N_{3,1}$ in those NAPs for simplicity.
+[…]
+We introduce OptAdvPrune to identify essential neurons.
+Intuitively, it attempts to show a neuron $N_{i,l}$ is essential by actively falsifying NAP candidates with adversarial examples.
+When an adversarial example $x^{\prime}$ is found, it immediately indicates that the NAP $\ddot{\mathcal{A}}(N,x^{\prime})$ fails the verification, i.e., $\mathcal{V}(\ddot{\mathcal{A}}(N,x^{\prime}))=0$.
+Moreover, it also implies that any NAP subsumed by $\ddot{\mathcal{A}}(N,x^{\prime})$ fails verification.
+For instance, suppose an adversarial example $x^{\prime}$ is found for a simple one-layer four-neuron neural network and $\ddot{\mathcal{A}}(N,x^{\prime})$ is $\langle\mathbf{1},\mathbf{0},\mathbf{1},\mathbf{0}\rangle$.
+We can infer that NAPs like $\langle\mathbf{1},\mathbf{0},\mathbf{1},\mathbf{*}\rangle$, $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{*}\rangle$, $\langle\mathbf{1},\mathbf{*},\mathbf{*},\mathbf{0}\rangle$, and $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle$ fail the verification.
+This information is particularly useful when determining if a neuron is essential.
+For example, if we know that NAP $P:=\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{1}\rangle$ is a specification, i.e., $\mathcal{V}(P)=1$, then we can easily deduce that the fourth neuron $N_{4,1}$ is essential.
+This is because that coarsening the fourth neuron would expand $P$ to $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{*}\rangle$, which would include the adversarial example $x^{\prime}$ and thus fail the verification, as illustrated in Figure 4(a).
+It is evident that the neuron where $P$ and $\ddot{\mathcal{A}}(N,x^{\prime})$ disagrees must be essential.
+Input: The neural network $N$, the input dataset $X$
+Output: A collection of essential neurons
+1
+2Function OptAdvPrune($N$, $X$)
+3       $Mand\leftarrow\emptyset$ ; $P\leftarrow\widetilde{\mathcal{A}}(N)$
+4       for $x_{j}$ in $X$ do
+5             $x^{\prime}_{j}\leftarrow Adversarial\_Attack(x_{j})$
+6             for $N_{i,l}$ in $N$ do
+7                   if $P_{i,l}\in\{\mathbf{0},\mathbf{1}\}$ and $\ddot{\mathcal{A}}(N_{i,l},x^{\prime}_{j})\oplus P_{i,l}$ then
+                         $Mand\leftarrow Mand\cup\{N_{i,l}\}$
+                          /* Create essential neurons for $x^{\prime}_{j}$/
+8
+9
+10      return $Mand$
+11
+Algorithm 3 OptAdvPrune
+However, when the two NAPs disagree on multiple neurons, things become a little bit different.
+Suppose the NAP specification $P$ is $\langle\mathbf{1},\mathbf{1},\mathbf{*},\mathbf{1}\rangle$, i.e., $\mathcal{V}(\langle\mathbf{1},\mathbf{1},\mathbf{*},\mathbf{1}\rangle)=1$.
+We know $\mathcal{V}(\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle)=0$ by the adversarial example $x^{\prime}$.
+In this case, if we coarsen the second and fourth neurons, $N_{2,1}$ and $N_{4,1}$, the NAP specification will expand to $\langle\mathbf{1},\mathbf{*},\mathbf{*},\mathbf{*}\rangle$, which will cover the $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle$, thus failing the verification.
+In this case, $N_{2,1}$ and $N_{4,1}$ could both be essential neurons or either one of them is essential, as illustrated in Figures 4(b), 4(c), 4(d).
+So, we simply let $\{N_{2,1},N_{4,1}\}$ be the upper bound of essential neurons (learned from $x^{\prime}$).
+Formally, given a NAP $P$, we say a neuron $N_{i,l}$ is in the upper bound of essential neurons $E$ if satisfies the following condition:
+  1. (1)
+    $N_{i,l}$ must be in the binary states, i.e., $P_{i,l}\in\{\mathbf{0},\mathbf{1}\}$
+  2. (2)
+    There exists $x^{\prime}$ such that $\ddot{\mathcal{A}}(N_{i,l},x^{\prime})$ XORs with $P_{i,l}$, i.e., $\exists x^{\prime}\text{ such that }\ddot{\mathcal{A}}(N_{i,l},x^{\prime})\oplus P_{i,l}=1$
+The field of adversarial attacks provides a diverse set of computationally efficient methods, enabling access to numerous adversarial examples.
+In this study, we use a collection of different attacks, including the Projected Gradient Descent (PGD) attack (Madry et al., 2018) and the Carlini-Wagner (CW) attack (Carlini and Wagner, 2017).
+By computing the upper bound of essential neurons for each adversarial example and taking their union, we efficiently estimate the overall upper bound, as shown in Algorithm 3.
 5. Volume Estimation of $R_{P}$
@@
 6. Evaluation
-In this section, we conduct a comprehensive evaluation of our methods for learning minimal NAP specifications across a range of benchmarks, spanning from a simple binary classifier to the state-of-the-art image classification model.
-We illustrate that minimal NAP specifications usually involve only a fraction of neurons compared to the most refined NAPs, yet they dramatically extend the verifiable bounds to several magnitudes larger.
+In this section, we conduct a comprehensive evaluation of our algorithms for learning minimal NAP specifications across a range of benchmarks, from decision-critical tasks in cancer diagnosis to state-of-the-art image classification models.
+To illustrate the effectiveness of our approaches, we chose the method proposed in (Geng et al., 2023) as the baseline, denoted as the $\widetilde{\mathcal{A}}$ function.
 Experiment Setup
@@
 For verification, we utilized Marabou (Katz et al., 2019), a dedicated state-of-the-art neural network verifier.
-We configured a timeout of 10 minutes for each call to the verification tool.
+We set a timeout of 5 minutes for each query call to the verification tool.
 If the timeout is exceeded, the current neuron is retained in the minimal NAP specification even if its status cannot be determined.
-[…]
-0 1
-$P$$\mathcal{V}$ train test $P$$\mathcal{V}$ train test
-The $\widetilde{\mathcal{A}}$ function 102 - 78.11 81.40 93 - 87.06 80.28
-Coarsen 31 102 98.22 95.35 32 93 99.65 94.37
-Adversarial_Prune 61 - 81.06 88.37 54 - 87.06 80.28
-Gradient_Search 43 - 82.84 88.37 39 - 93.37 87.32
-Sample_Refine 53 187 92.43 89.37 56 264 89.86 88.51
-Sample_Coarsen 42 47 94.69 92.34 45 41 94.15 91.52
-Wisconsin Breast Cancer with Binary Classifier
+6.1. The Wisconsin Breast Cancer Dataset with Binary Classifier
 We conduct our first experiment using a four-layer neural network as a binary classifier, where each layer consists of 32 neurons.
-This classifier is trained on the Wisconsin Breast Cancer (WBC) dataset (William et al., 1995), representing a decision-critical task where robustness is essential.
+This classifier is trained on the Wisconsin Breast Cancer (WBC) dataset (Wolberg et al., 1995).
 Our trained model achieves a test set accuracy of 95.61%.
-We calculate the most refined NAP specifications $\widetilde{P}^{0}$ and $\widetilde{P}^{1}$ for labels 0 and 1 using the statistical abstraction function $\widetilde{\mathcal{A}}$ with a confidence ratio of $\delta=0.95$.
+We calculate the most refined (baseline) NAP specifications $\widetilde{P}^{0}$ and $\widetilde{P}^{1}$ for labels 0 and 1 using the statistical abstraction function $\widetilde{\mathcal{A}}$ with a confidence ratio of $\delta=0.95$.
 The size of $\widetilde{P}^{0}$ and $\widetilde{P}^{1}$ are 102 and 93, respectively.
 In contrast, the sizes of the minimal NAP specifications learned by the Coarsen algorithm for labels 0 and 1 are significantly reduced to 31 and 32, respectively.
-It is worth mentioning that our estimation approaches provide a fairly accurate estimate of the mandatory neurons, despite computing rather loose upper bounds.
-To be more specific, Adversarial_Prune and Gradient_Search compute 61 and 43 mandatory neurons for label 0, respectively.
-Together, they cover 25 out of the 31 mandatory neurons appearing in the minimal NAP specification for label 0.
-For label 1, Adversarial_Prune and Gradient_Search compute 54 and 39 mandatory neurons, respectively, covering 25 out of the 32 mandatory neurons appearing in the minimal NAP specification for label 1.
-Regarding the statistical approaches, Sample_Refine computes NAP specifications of size 53 and 56 for label 0 and label 1, respectively.
-This is achieved while making 187 and 264 calls, respectively.
-Although these numbers are quite large, it is expected since it has to sample multiple trials in each iteration.
-In contrast, Sample_Coarsen is significantly more efficient.
-It can learn NAP specifications of size 42 and 45 for label 0 and label 1 using only 47 and 41 calls, respectively.
-Recall that one of the main motivations for learning the minimal NAP specifications is their potential to verify larger input regions compared to the refined NAP specifications.
-To support this claim, we compute the percentile of unseen test data that can be verified using these NAPs.
-We choose test data as they are sampled from the input space and can be viewed as a proxy to understand the verifiable bounds of different NAPs.
-We observe that the most refined NAP specifications $\widetilde{P}^{0}$ and $\widetilde{P}^{1}$ can cover 81.40% and 80.28% of the test data for labels 0 and 1, respectively.
-In contrast, this coverage significantly increases to 95.35% and 94.37% of the test data for the minimal NAP specifications of labels 0 and 1, respectively.
-To intuitively understand the extent of change in verifiable regions from the refined NAP specifications to the minimal NAP specifications, we also compare their estimated volume of the corresponding region $R_{P}$.
-The estimated volumetric changes are substantial, with label 0 experiencing an increase 58324.35 times larger and label 1 showing an increase 3053.53 times larger, respectively.
-To comprehend this change in each dimension, Figure 8 illustrates the comparison of verifiable input ranges between the refined NAP specifications and the minimal NAP specifications for labels 0 and 1, referencing the anchor point.
-(a) Comparing verifiable input ranges for label 0
-(b) Comparing verifiable input ranges for label 1
+It is worth mentioning that our optimistic approach provides a fairly accurate estimate of the essential neurons, despite computing rather loose upper bounds.
+To be more specific, OptAdvPrune compute 43 essential neurons for label 0, respectively, covering 25 out of the 31 essential neurons appearing in the minimal NAP specification for label 0 computed by Coarsen.
+For label 1, OptAdvPrune computes 39 essential neurons, covering 25 out of the 32 essential neurons appearing in the minimal NAP specification for label 1 computed by Coarsen.
+The NAP specification learnt from OptAdvPrune are verified, showing it provides an optimistic upper bound to the size of minimal NAP with minimal verification calls yet finding neurons .
+We show it’s efficiency on small models.
+This sheds light to solving minimal NAP problem on large models when using Coarsen and StochCoarsen are impossible given the current limit on verification engine.
+Regarding the statistical approaches, StochCoarsen learnt NAP specifications of size 42 and 45 for label 0 and label 1 using only 47 and 41 calls to verification respectively.
+Recall that one of the main motivations for learning the minimal NAP specifications is their potential to verify larger input regions compared to refined NAP specifications.
+To support this, we compute the percentile of unseen test data that can be verified using these NAPs.
+Test data, sampled from the input space, serve as a proxy to understand the verifiable bounds of different NAPs.
+We find that the most refined NAP specifications $\widetilde{P}^{0}$ and $\widetilde{P}^{1}$ cover 81.40% and 80.28% of test data for labels 0 and 1, respectively.
+In contrast, minimal NAP specifications cover 95.35% and 94.37% of test data for labels 0 and 1, respectively.
+To intuitively understand the change in verifiable regions $R_{P}$ from refined to minimal NAP specifications, we compare their estimated volumes.
+The increase in estimated volume is substantial—on the order of $10^{5}$ times larger for labels 0 and 1.
+Table 1.
+Overview of the size of learned minimal NAP specifications using various approaches on the WBC benchmark.
+The columns represent different labels, while each row corresponds to a different algorithm.
+$P$ denotes the size of the learned NAP, and #$\mathcal{V}$ represents the number of calls to $\mathcal{V}$.
+The train and test columns report the percentile (%) of train and test data covered by $P$.
+We assess the effectiveness of NAP specifications in covering both training and test data, reporting the coverage percentage (%) along with the estimated change in volume, expressed as an order of magnitude (x) relative to a baseline normalized to 1.
+0 1
+$P$$\mathcal{V}$ train test $vol.$ $P$$\mathcal{V}$ train test $vol.$
+$\widetilde{\mathcal{A}}$ function (baseline.) 102 1 78.11 81.40 1 93 1 83.06 80.28 1
+Coarsen 31 102 98.22 95.35 $10^{5}$ 32 93 99.65 94.37 $10^{5}$
+StochCoarsen 42 47 94.69 92.34 $10^{3}$ 45 41 94.15 91.52 $10^{2}$
+OptAdvPrune 61 1 91.06 89.37 $10^{2}$ 54 3 87.06 87.28 $10^{2}$
+6.2. MNIST with Fully Connected Network
+To show that our insights and approaches can be applied to more complicated datasets and networks, we conduct the second set of experiments using the mnistfc_256x4 model (VNNCOMP, 2021), a 4-layer fully connected network with 256 neurons per layer trained on the MNIST dataset.
+Due to space constraints, we present results for randomly selected classes 0, 1, and 4, with the remaining results provided in Appendix A.
+Table 2. Overview of learned minimal NAP specifications on the MNIST benchmark.
+0 1 4
+$|P|$$\#\mathcal{V}$ traintestvol. $|P|$$\#\mathcal{V}$ traintestvol. $|P|$$\#\mathcal{V}$ traintestvol.
+$\widetilde{\mathcal{A}}$ function (baseline.) 751 1 79.50 71.51 1 745 1 75.01 70.11 1 712 1 77.54 75.24 1
+Coarsen 480 751 98.68 98.78 $10^{8}$ 491 745 98.90 98.59 $10^{6}$ 506 712 98.51 97.45 $10^{9}$
+StochCoarsen 532 33 93.19 93.01 $10^{5}$ 559 27 94.12 93.68 $10^{3}$ 562 25 93.89 93.64 $10^{6}$
+OptAdvPrune 618 15 83.13 81.32 $10^{2}$ 630 18 86.02 85.51 $10^{2}$ 699 21 82.66 84.12 $10^{2}$
+We compute the most refined NAP specifications, $\widetilde{P}^{0}$, $\widetilde{P}^{1}$, and $\widetilde{P}^{4}$, for these labels with a confidence level of $\delta=0.99$.
+Their sizes are 751, 745, and 712, respectively, consistent with the baseline results from previous work (Geng et al., 2023).
+In contrast, the minimal NAP specifications learned by the Coarsen algorithm for labels 0, 1, and 4 are significantly reduced to 480, 491, and 506, respectively, despite requiring over 700 calls.
+These specifications cover more than 98% of the training and test data, demonstrating a significant improvement over the baseline.
+For estimated volume, the volumetric changes can be on the order of magnitude of $10^{9}$ larger.
+Notably, StochCoarsen achieves comparable results with only around 30 calls, at the cost of approximately a 5% reduction in coverage on both training and test data compared to Coarsen.
+This highlights the potential of these learned minimal NAPs as robust specifications that generalize well to unseen data from the same distribution.
+In comparison, our empirical analysis reveals that local neighborhood specifications cover zero test images in the entire MNIST dataset with $\epsilon=0.2$, which is the maximum $L_{\infty}$ verifiable bound used in VNNCOMP (Brix et al., 2023), the annual neural network verification competition.
+This further validates our motivation for using NAPs as specifications.
+For instance, for label 0, OptAdvPrune detects 618 essential neurons and correctly identify 445 and 160 of the 480 neurons in the minimal NAP founded by Coarsen.
+When interacting with $\mathcal{V}$, OptAdvPrune computes minimal NAPs with an average of fewer than 20 calls.
+However, it experiences a decline in data coverage, with an average drop of over 5% compared to the baseline.
+In summary, there is no universal solution to a given minimal specification problem.
+Our experiment highlights the distinct strengths and trade-offs of the three algorithms.
+While the results from all have gurantee on correctness, Coarsen finds the minimal NAP among the three, making it the most reliable for applications where rigorous minimality is priority.
+StochCoarsen offers a significant speed advantage on top of Coarsen while maintaining minimality: the learned minimal NAP specification uses approximately 1.1x more neurons but requires only 5% of the verification calls.
+This significant reduction in computational overhead makes StochCoarsen particularly well-suited for scenarios where efficiency is a priority.
+Finally, OptAdvPrune, employing an optimistic approach, is the fastest.
+The initial set of neurons, when they can be successfully verified, provides a tight upper bound for the global minimum.
+When the initial set cannot be verified, it provides a effective starting point for StochCoarsen and Coarsen.
+Moreover, OptAdvPrune can be utilized in a verification-dependent manner, they excel in scenarios that scale beyond our verification capabilities.
+In these situations, they provide accurate estimations of essential neurons, enabling us to examine potential causal links between neurons and the interpretability and robustness of deep neural networks, as we will illustrate in the next experiment.
+Together, these algorithms apply to diverse scenarios, allowing users to balance speed and minimality accuracy, and guarantee of validity based on their specific requirements.
+6.3. ImageNet with Deep Convolutional Neural Network
+In our third experiment, we present our finds on the fully connected layers  [^1] NAP computation for convolutional layers is left for future work.
+of VGG-19 network pretrained on the ImageNet dataset, consisting of 8192 neurons.
+Provided it is currently impossible to verify NAP specifications on models of this size, we focus on analyzing the effect of essential neurons estimated using OptAdvPrune.
+We also limit our analysis to the top five largest classes, each containing approximately 1000 training and 350 test images.
+Our findings indicate that these estimated NAPs cover a significant portion of unseen test data, highlighting their potential as robust certificates for the test set.
+More details can be found in Appendix B.
+NAP Captures Visual Interpretability and Inherent Robustness
+From the perspective of representation learning, neural networks acquire both low- and high-level feature extractors, which they use to make final classification decisions based on hidden features (neuron representations) (Bengio et al., 2013).
+Therefore, the robustness and consistency of a model’s predictions are influenced by the quality of these learned features.
+In essence, achieving an accurate and robust model hinges on learning ”good” hidden representations, which are characterized by better interpretability (Zhang et al., 2021).
+Many studies suggest a close relationship between visual interpretability and robustness, often observed in the learned features and representations (Alvarez Melis and Jaakkola, 2018; Boopathy et al., 2020; Dong et al., 2017).
+Thus, although we cannot yet formally verify the correctness of these estimated NAP specifications, we demonstrate that these NAPs are indeed ”meaningful” through visual interpretability—strong evidence that the estimated essential neurons (NAPs) contribute to the model’s robustness.
 Figure 6.
-Comparison of verifiable input ranges between the refined NAP specifications and the minimal NAP specifications for labels 0 and 1 on the Wisconsin Breast Cancer dataset, with reference to the anchor point.
-The minimal NAP specifications enable verification over a larger input range for all dimensions.
-Table 2. Overview of the size of learned minimal NAP specifications on the MNIST benchmark.
-0 1 4
-$P$$\mathcal{V}$ train test $P$$\mathcal{V}$ train test $P$$\mathcal{V}$ train test
-The $\widetilde{\mathcal{A}}$ function 751 - 79.50 80.51 745 - 86.01 85.11 712 - 77.54 80.24
-Coarsen 480 751 98.68 98.78 491 745 98.90 98.59 506 712 98.51 97.45
-Adversarial_Prune 618 - 83.13 71.32 630 - 79.02 76.51 699 - 82.66 84.12
-Gradient_Search 195 - 92.1 91.80 237 - 89.73 88.39 169 - 87.59 88.51
-Sample_Refine 564 4059 90.03 87.91 570 3663 88.36 86.84 581 4532 82.90 83.46
-Sample_Coarsen 532 33 93.19 93.01 559 27 94.12 93.68 562 25 93.89 93.64
-MNIST with Fully Connected Network
-To show that our insights and approaches can be applied to more complicated datasets and networks, we conduct the second set of experiments using the mnistfc_256x4 model (VNNCOMP, 2021), a 4-layer fully connected network with 256 neurons per layer trained on the MNIST dataset.
-We specifically focus on classes 0, 1, and 4, as the most refined NAPs for other labels either encounter timeouts or fail the verification.
-We compute the most refined NAP specifications $\widetilde{P}^{0}$, $\widetilde{P}^{1}$, and $\widetilde{P}^{4}$ for labels 0, 1, and 4 using the $\widetilde{\mathcal{A}}$ function with a confidence ratio of $\delta=0.99$.
-The size of $\widetilde{P}^{0}$, $\widetilde{P}^{1}$, and $\widetilde{P}^{4}$ are 751, 745 and 712, respectively, which aligns with the results from (Geng et al., 2023a).
-In contrast, the sizes of the minimal NAP specifications learned by the Coarsen algorithm for labels 0, 1, and 4 are significantly reduced to 480, 491, and 506, respectively.
-Moreover, our estimation approaches provide highly accurate estimates of the mandatory neurons present in minimal NAP specifications.
-For instance, considering label 0, Adversarial_Prune and Gradient_Search identify 618 and 195 mandatory neurons, respectively.
-And 445 and 160 out of the 480 mandatory neurons presented in the minimal NAP specification are discovered by Adversarial_Prune and Gradient_Search, respectively.
-[…]
-Moreover, the learned minimal NAP specifications correspond to significantly larger verifiable regions compared to the refined NAP specifications.
-Using the percentile of test data as a metric, the minimal NAP specifications increase the coverage ratio from 80.51% to 98.78%, 85.11% to 98.59%, and 80.24% to 97.45% for labels 0, 1, and 4, respectively.
-From the perspective of estimated volume, the volumetric changes are on the order of $10^{8}$ times larger.
-ImageNet with Deep Convolutional Neural Networks
-Current verification approaches struggle to scale to state-of-the-art neural networks such as deep convolutional neural networks, making them inapplicable for verifying certain critical systems like deep learning-based autonomous driving software.
-Consequently, our verification-dependent approaches also face scalability challenges.
-However, we can still assess the network’s robustness from a neuron-centric perspective using our estimation approaches, an aspect not well-studied in the current literature.
-Thus, for the third experiment, we choose a deep convolutional network, specifically VGG-19 trained on the ImageNet dataset, as the benchmark for estimating mandatory neurons.
-We select neurons from fully connected layers [^1] We leave NAP computation for other layers such as convolutional layers for future work.
-of VGG-19, which contains a total of 8192 neurons, to compute NAPs.
-The original ImageNet includes 1000 classes with millions of images.
-We also narrow our focus to the top five largest classes in ImageNet.
-Each class consists of around 1000 training images and 300 test images.
-Table 3 presents the results of the learned NAP using verification-free approaches.
-To show that these estimated mandatory neurons are indeed meaningful, we employ several approaches from the model interpretability domain such as Grad-CAM (Selvaraju et al., 2017) and saliency maps (Simonyan et al., 2014a).
-We observe a strong correlation between neurons exhibiting high interpretability under certain metrics and our estimated mandatory neurons computed by Adversarial_Prune and Gradient_Search.
-However, we also find that some mandatory neurons do not seem to participate significantly in interpretability, yet they are indeed necessary for the robustness of models.
-We find this interesting and plan to investigate those neurons in future work.
-Although we cannot definitively determine if these NAPs can serve as specifications, these NAPs do cover significant portions of unseen test data.
-For example, the NAP formed by mandatory neurons learned through Gradient_Search can cover $87.01\%$ of unseen test data.
-This aspect makes them a promising basis for computing NAP specifications, especially if verification tools become accessible for such large models in the future.
-Table 3. Overview of the size of learned minimal NAP specifications on the ImageNet benchmark.
-box_turtle spiny_lobster acorn_squash confectionery stone_wall
-$P$ test $P$ test $P$ test $P$ test $P$ test
-The $\widetilde{\mathcal{A}}$ function 1978 39.42 1006 39.61 1003 29.23 878 33.84 971 31.54
-Adversarial_Prune 1863 39.42 955 39.61 823 25.68 845 23.07 865 39.61
-Gradient_Search 611 53.30 572 87.01 301 52.40 256 44.23 260 54.90
+Visual interpretability of learned hidden representations retained by estimated NAPs.
+The first two columns represent the original images and GradCams, respectively.
+The third and fourth columns represent modified GradCams using baseline NAPs of the same class and another class, respectively.
+The last two columns represent modified GradCams using minimal NAPs of the same class and another class, respectively.
+To this end, we make a simple modification to Grad-CAM (Selvaraju et al., 2017), which highlights key regions of an input image by using the gradients of the classification score with respect to the final convolutional feature map.
+Specifically, we mask out the neurons that do not appear in NAP in the fully connected layers.
+We then recalculate the backward gradient flow using this modified computation graph to generate the updated Grad-CAM map.
+Finally, we conduct the following experiments on image samples:
+  1. (1)
+    Calculate the modified Grad-CAM map using the most refined (baseline) NAP and compare it with the original Grad-CAM map.
+  2. (2)
+    Calculate the modified Grad-CAM map using minimal NAPs estimated by OptAdvPrune, and compare it with the original map.
+  3. (3)
+    Calculate the modified Grad-CAM map using NAPs from different classes and compare it with the original map.
+All GradCam maps are further validated with sanity checks (Adebayo et al., 2018).
+Figure 6 presents the experimental results.
+The original GradCams highlight important regions of images corresponding to crucial justifications for classification.
+Notably, both the most refined and estimated minimal NAPs produce nearly identical highlights, despite representing only a small fraction of neurons.
+This shows that the neuron abstractions from our minimal specification preserve the key visual features.
+Further, it suggests that the estimated minimal NAPs capture crucial aspects of VGG-19’s internal decision-making process, supporting the ”NAP robustness property.” Furthermore, Grad-CAMs generated from NAPs of different classes highlight distinct regions, strongly indicating that our estimated NAPs are clearly distinguishable, aligning with the ”non-ambiguity property.”
+These results demonstrate that NAPs offer valuable insights into interpretability.
+A small subset of neurons from NAPs can categorize critical internal dynamics of neural networks, potentially helping us unveil the black-box nature of these systems.
+From a machine-checkable definition perspective, concise NAPs are easier to decode into human-understandable programs than more refined NAPs.
+This underscores the importance of learning minimal NAPs.
+Interpreting NAPs into human-readable formats remains a direction for future research.
+NAP as a Defense Against Adversarial Attacks
+From a practical point of view, we believe that even before 1) formal verification finally scales; and/or 2) NAPs are fully interpretable, NAPs, as they are now, can already serve as an empirical certificate of a prediction or some defense mechanism, as shown in recent work (Lukina et al., 2021).
+In the same spirit, we demonstrate that our estimated essential neurons can serve as defense against adversarial attacks.
+We first select images that meet two criteria: 1) correctly predicted by the model; and 2) covered by the respective NAP.
+On average, each NAP covers approximately 40% of the training data of the corresponding class.
+For each selected image, we generate 100 distinct adversarial examples that are misclassified by the model, using Projected Gradient Descent attack (Madry et al., 2018) and Carlini Wagner attack (Carlini and Wagner, 2017) respectively.
+We then check whether each adversarial image’s activation pattern is rejected by the respective NAPs.
+If so, we conclude that NAPs can empirically serve as specifications; Otherwise, they are proven to be not robust.
+Notably, we find that both the baseline NAPs and the estimated minimal NAPs reject all adversarial examples, indicating their effectiveness in describing a safe region and their potential as certificates.
 7. Related Work
@@
 Neural network verification has attracted much attention due to the increasingly widespread applications of neural networks in safety-critical systems.
-it’s NP-hard nature resulting from the non-convexity introduced by activation functions (Katz et al., 2017) makes it a challenging task.
+it’s NP-hard nature resulting from the non-convexity introduced by activation functions (Katz et al., 2017b) makes it a challenging task.
 Thus most of the existing work on neural network verification focuses on designing scalable verification algorithms.
-For instance, while initially proposed solver-based approaches (Huang et al., 2017b; Ehlers, 2017; Cheng et al., 2017; Tjeng et al., 2019) were limited to verifying small-scale neural networks with fewer than 100 neurons, state-of-the-art approaches (Xu et al., 2021; Wang et al., 2021c; Lu and Kumar, 2019) can verify more complex neural networks.
+For instance, while initially proposed solver-based approaches (Huang et al., 2017b; Ehlers, 2017; Cheng et al., 2017; Tjeng et al., 2019) were limited to verify small neural networks with fewer than 100 neurons, state-of-the-art methods (Xu et al., 2021; Wang et al., 2021c; Lu and Kumar, 2019) can verify more complex neural networks.
 It is worth mentioning that most existing work adopts local neighborhood specifications to verify the robustness properties of neural networks (Shapira et al., 2023).
-Despite being a reliable measure, using specifications that define local neighborhood of reference data points may not cover any test data, let alone generalizing to the verification of unseen test set data.
-Yet evidently, being able to verify the neural network on test data is an important task in real-world applications.
-The development of NAP specification (Geng et al., 2023a) sheds light on this challenge.
-While previous work on NAP specifications proves its effectiveness, it assumes that all neurons contribute equally to the model’s robustness.
-In this paper, we challenge this assumption by demonstrating that only a small fraction of neurons are crucially related to robustness.
-This not only allow us study robustness from the neurons’ perspective but also significantly pushes the known verifiable boundaries.
+Despite being a reliable measure, using local neighborhood specifications around reference data points may not cover any test data, let alone generalizing to the verification of unseen test set data.
+Geng et al.
+(2023) propose the new paradigm of NAP specifications to address this challenge.
+Our work advances the understanding of NAP specifications.
 7.2. Abstract Interpretation
@@
 We observe that neuron values exhibit different patterns in range for different input classes, suggesting the potential existence of more abstraction states.
-We plan to explore this direction in our future work.
+We leave this as future work.
 7.3. Neural Activation Patterns
@@
 One popular line of research is feature visualization (Yosinski et al., 2015; Bäuerle et al., 2022), which investigates which neurons are activated or deactivated given different inputs.
-This is also naturally related to the field of activation maximization (Bau et al., 2017; Simonyan et al., 2014b; Erhan et al., 2009), which studies what kind of inputs could mostly activate certain neurons in the neural network.
+This is also naturally related to the field of activation maximization (Simonyan et al., 2014), which studies what kind of inputs could mostly activate certain neurons in the neural network.
 In this way, certain prediction outcomes may be attributed to the behavior of specific neurons, thereby increasing the interpretability of the underlying models.
-In summary, most of existing works focus on learning statistical correlations between either NAPs and inputs or NAPs and prediction outcomes.
-However, whether such correlations can be trusted or even verified is not fully studied.
-To this end, we propose the concept of mandatory neurons and highlight their importance in the robustness of model predictions.
+Lukina et al.
+(2021) demonstrates that neural activation patterns can be used to monitor neural networks and detect novel or unknown input classes at runtime.
+They provide human-level interpretability of neural network decision-making.
+In summary, most of existing works focus on learning statistical correlations between NAPs and inputs (Bau et al., 2017; Erhan et al., 2009), or between NAPs and prediction outcomes (Lukina et al., 2021).
+However, these correlations raises questions which we address in this paper: whether the correlation can be trusted or even verified.
+We propose the concept of essential neurons and highlight their importance in the robustness of model predictions.
 Such causal links between neurons and prediction outcomes are not only identified but also verified.
 We believe this ”identify then verify” paradigm can be extended to existing research on NAPs to certify our understanding of neural networks.
-And we leave the exploration of this direction for our future work.
+We leave the exploration of this direction for our future work.
 8. Conclusion
-We introduce a new problem — learning the minimal NAP specification, and discuss its importance in neural network verification.
-Finding minimal NAP specification not only enables the verification of larger input regions compared to existing methods but also provides a means of inspecting when and how neural networks can make reliable and robust predictions.
-To solve this problem, we first propose two simple approaches Refine and Coarsen, which leverage off-the-shelf verification tools that find the minimal NAP specification with correctness guarantees.
-We also propose the statistical version of Refine and Coarsen, that combine sampling and statistical learning principles to achieve correctness in an more efficient manner.
-However, these approaches depend on underlying verification tools that are computationally expensive.
-To this end, we propose two approximate approaches, Adversarial_Prune and Gradient_Search.
-These methods utilize adversarial attacks and local gradient computation to efficiently investigate potential causal relationships between specific neurons and the model’s robustness.
-This is a task that current verification frameworks struggle to accomplish.
-Finally, to appreciate the volumetric change from the most refined NAP specification to the minimal NAP specification, we propose a simple method for estimating the volume of the region corresponding to a NAP.
-Our experiments indicate that minimal NAP specifications utilize much smaller fractions of neurons compared to the most refined NAPs.
-Nevertheless, they enable a substantial expansion of the verifiable boundaries by several orders of magnitude.
-Appendix A Proofs of simple approaches
-Theorem A.1 (Property of Refine).
+We introduce a new challenge: learning the minimal NAP specification and highlighting its significance in neural network verification.
+Identifying minimal NAP specifications not only facilitates the verification of larger input regions compared to existing methods but also provides insights into when and how neural networks make reliable and robust predictions.
+To address this problem, we propose three approaches—conservative, statistical, and optimistic—each offering distinct trade-offs between efficiency and performance.
+The first two rely on the verification tool to find minimal NAP specifications.
+The optimistic method efficiently estimates minimal NAPs using adversarial examples, without making calls to the verification tool until the very end.
+Each of these methods offers distinct strengths and trade-offs in terms of minimality and computational speed, making each approach suitable for scenarios with different priorities.
+The learnt minimal NAP specification allows us to inspect potential causal links between neurons and the robustness of state-of-the-art neural networks, a task for which existing work fails to scale.
+Our experimental results suggest that minimal NAP specifications require much smaller fractions of neurons compared to the NAP specifications computed by previous work, yet they can significantly expand the verifiable boundaries to several orders of magnitude larger.
+Appendix A Learned Minimal NAP Specifications on the MNIST Benchmark (Complement)
+Figure 7.
+The comparison between NAP specifications (green region) and $L_{\infty}$ ball specifications (gray region) on the MNIST dataset.
+Image $a$ is the reference image, and image $c$ is the closest among all 6000 training images of digit 1, with an $L_{\infty}$ distance of 0.5294.
+However, $c$ cannot be verified using the $L_{\infty}$ ball specification, as an adversarial example $b$ exists at an $L_{\infty}$ distance of 0.2.
+Note that this is not a limitation of the underlying verification engines but rather an intrinsic limitation of the specifications.
+In contrast, NAP specifications allow verification of unseen test set data like $c$.
+Table 3. Overview of the size of learned minimal NAP specifications on the MNIST benchmark.
+2 3 5 6
+$|P|$$\mathcal{V}$ train test $|P|$$\mathcal{V}$ train test $|P|$$\mathcal{V}$ train test $|P|$$\mathcal{V}$ train test
+Baseline (Geng et al., 2023) 751 1 79.50 80.51 745 1 86.01 85.11 712 1 77.54 80.24 1 712 1 77.54 80.24
+Coarsen 480 751 98.68 98.78 491 745 98.90 98.59 506 712 98.51 97.45 503 708 98.81 98.39
+StochCoarsen 532 33 93.19 93.01 559 27 94.12 93.68 562 25 93.89 93.64 542 29 98.81 98.39
+OptAdvPrune 618 15 83.13 71.32 630 18 79.02 76.51 699 21 82.66 84.12 681 20 85.98 87.63
+Table 4. Overview of the size of learned minimal NAP specifications on the MNIST benchmark.
+7 8 9
+$P$$\mathcal{V}$ train test $P$$\mathcal{V}$ train test $P$$\mathcal{V}$ train test
+The $\widetilde{\mathcal{A}}$ function 751 1 79.50 80.51 745 1 86.01 85.11 712 1 77.54 80.24
+Coarsen 480 751 98.68 98.78 491 745 98.90 98.59 506 712 98.51 97.45
+Adversarial_Prune 618 15 83.13 71.32 630 18 79.02 76.51 699 21 82.66 84.12
+Sample_Coarsen 532 33 93.19 93.01 559 27 94.12 93.68 562 25 93.89 93.64
+Appendix B Learned Minimal NAP Specifications on ImageNet with Deep Convolutional Neural Network (Complement)
+Table 5. Overview of the size of learned minimal NAP specifications on the ImageNet benchmark.
+box_turtle labrador_retriever acorn_squash confectionery stone_wall
+$P$ test $P$ test $P$ test $P$ test $P$ test
+The $\widetilde{\mathcal{A}}$ function 1978 39.42 691 39.61 1003 29.23 878 33.84 971 31.54
+Adversarial_Prune 1863 39.42 661 41.28 823 25.68 845 23.07 865 39.61
+Gradient_Search 611 53.30 572 87.01 301 52.40 256 44.23 260 54.90
+Appendix C The Refine Approaches
+C.1. The Refine Approach
+Conceptually, the Refine approach iteratively increases the number of refined neurons in NAP $P$ until $\mathcal{V}(P)=1$, i.e., $P$ is able to prove the underlying robustness query.
+In other words, we gradually increase the size parameter $k$ and iterate over each NAP $P$ of size $k$ to check if $\mathcal{V}(P)=1$, as illustrated in Algorithm 4.
+To determine if a solution to the problem exists, we first check if the most refined NAP can succeed in verification.
+We proceed to iterative refinement only if $\mathcal{V}(\widetilde{P})=1$.
+However, the algorithm is not efficient and requires $12^{|N|}-$ 1 calls to $\mathcal{V}$ in the worst case, as proven in Theorem C.1.
+Please refer to the proof in Appendix D.
+Therefore Refine is only practical when the search space of the NAP family $\mathcal{P}$ is small.
+Theorem C.1.
+The algorithm Refine returns a minimal NAP specification with $\mathcal{O}(2^{|N|})$ calls to $\mathcal{V}$.
+1
+Input: The neural network $N$
+Output: A minimal NAP specification $P$
+2
+3Function Refine(N)
+4       $P\leftarrow\widetilde{\mathcal{A}}(N)$
+5       if $\mathcal{V}(P)==0$ then
+             return $None$ ;
+              /* Return None if even the most refined NAP fails verification */
+6
+7       else
+8             for $k$ from 1 to $|N|$ do
+                   $P\leftarrow\dot{\mathcal{A}}(N)$
+                    /* Refinement starts from the coarsest NAP $\dot{P}$/
+                   $K\_comb\leftarrow$ $Pick\_K\_Neurons(k,N)$
+                    /* Returns all combinations of size k */
+9                   for $comb\textbf{ in }K\_comb$ do
+10                         for $N_{i,l}\textbf{ in }comb$ do
+                               $P\leftarrow\widetilde{\Delta}(N_{i,l})$
+                                /* Refine neurons in the chosen combination */
+11
+12                        if $|\mathcal{V}(P)|==1$ then
+                               return $P$
+                                /* Found the minimal NAP */
+13
+14
+15
+16
+17
+Algorithm 4 refine
+C.2. The Statistical Refine Approach
+essential neurons are crucial for forming NAP specifications, as their binary states play a critical role in determining the neural network’s robustness performance.
+We leverages this property to find essential neurons statistically.
+To be more specific, suppose we sample some NAPs $P_{1},P_{2},\ldots,P_{n}$ from the NAP family $\mathcal{P}$.
+For those NAPs that qualify as specifications (i.e., $\mathcal{V}(P)=1$), essential neurons should appear more frequently in them than in those NAPs that fail the verification tool (i.e., $\mathcal{V}(P)=0$).
+Based on this insight, we propose an approach called Sample_Refine that relies on non-repetitive sampling to identify essential neurons for solving the minimal NAP problem.
+We start with the coarsest NAP and iteratively collect the most probable essential neurons.
+In every iteration, we sample $k$ NAPs by refining unvisited neurons with some probability $\theta$.
+The sampled NAPs are then fed to the verification tool, and the neuron that appears most frequently in verifiable NAPs is the most probable essential neuron in this iteration.
+The neuron is marked as visited, and the process stops when either we collect $s$ neurons (assuming that $s$ is known), or the current essential neurons form a NAP specification $P$, i.e., $\mathcal{V}(P)=1$.
+Finally, we return the learned NAP, obtained by applying $\widetilde{\mathcal{A}}$ to the collected neurons.
+The algorithm 5 provides an overview of the above procedure.
+It is worth noting that Sample_Refine doesn’t guarantee correctness, as we may end up collecting only the $s$ most probable essential neurons, which may not be sufficient to form a specification.
+Another concern regarding this algorithm is sampling efficiency, specifically the potential for the number of samples required to grow exponentially with the size of the minimal NAP specification $s$.
+To understand why this is problematic, consider a scenario where the only minimal NAP specification $P$ consists of all essential neurons; in this case, all $|M|$ neurons must be selected for $P$ to be learned.
+If $\theta$ is set to a constant value, then the expected number of samples needed to obtain the NAP specification is $(\frac{1}{\theta})^{|M|}$.
+To address this, we set $\theta=\left(\frac{|M|}{|M|+1}\right)^{|M|}$.
+This choice ensures that the sampling efficiency is polynomial in both $|M|$ and $s$, as proven in Theorem C.2.
+Please refer to the proof in the Appendix E.
+In addition, Theorem C.2 also shows that with high probability, a essential neuron will be found with $O(log|N|)$ calls to $\mathcal{V}$.
+Input: The neural network $N$, the probability $\theta$, sample size $k$, and the size $s$
+Output: A minimal NAP specification $P$
+1
+2Function Sample_NAPs($unvisited$, $\theta$)
+       $P\leftarrow\dot{\mathcal{A}}(N)$
+        /* Use the coarsest NAP as a blank template */
+3       for $N_{i,l}$ in $unvisited$ do
+4             $rand\leftarrow\text{random}(0,1)$
+5             if $rand\leq\theta$ then
+                   $P\leftarrow\widetilde{\Delta}(N_{i,l})$
+                    /* Refine unvisited neurons using $\widetilde{\Delta}$ with probability $\theta$/
+6
+7
+8      return $P$
+9
+10
+11Function Sample_Refine($visited$, $\theta$, $k$, $s$)
+12
+13      while $|visited|<s\text{ and }\mathcal{V}(P)==0$  do
+14             $unvisited\leftarrow N\setminus visited$
+             $ctr=dict()$
+              /* Create a counter for each neuron in $unvisited$/
+15             for __\__ in $range(k)$ do
+                   $P\leftarrow Sample\_NAPs(unvisited,\theta)$
+                    /* Sample $k$ NAPs */
+16                   for $N_{i,l}$ in $unvisited$ do
+17                         if $P_{i,l}==\widetilde{\mathcal{A}}(N_{i,l})$ and $\mathcal{V}(P)==1$ then
+                               $ctr[N_{i,l}]$ += 1
+                                /* Count successful NAPs for each neuron */
+18
+19
+20
+            $N_{mad}\leftarrow\operatorname*{argmax}_{N_{i,l}\in unvisited}\{ctr[N_{i,l}]\}$
+              /* Pick potential essential neuron */
+21             $visited\leftarrow visited\cup\{N_{mad}\}$
+22
+23      return $\widetilde{\mathcal{A}}(visited)$
+24
+25
+[…]
+27 if $\mathcal{V}(P)==0$ then
+       return $None$
+        /* Return None if the most refined NAP fails verification */
+28
+29 else
+30       if heuristics then
+31             $visited\leftarrow Gradient\_Search(N)\cap Adversarial\_Prune(N)$
+32       else
+             $visited\leftarrow\emptyset$
+              /* Start from the coarsest NAP */
+33
+34      $Sample\_Refine(visited,\theta,k,s)$
+35
+36
+Algorithm 5 Sample_Refine
+Theorem C.2.
+With probability $\theta=|(\frac{|M|}{|M|+1})|^{|M|}$, Sample_Refine has $1-\delta$ probability of outputting a minimal NAP specification with $\Theta\left(|M|^{2}(\log|N|+\log(s/\delta))\right)$ examples each iteration and $O(s|M|^{2}(\log|N|+\log(s/\delta)))$ total calls to $\mathcal{V}$.
+Appendix D Proofs of simple approaches
+Theorem D.1 (Property of Refine).
 The algorithm Refine returns a minimal NAP specification with $\mathcal{O}(2^{|N|})$ calls to $\mathcal{V}$.
@@
 In total, this number of NAPs we need to check is $\binom{|N|}{1}+\binom{|N|}{2}+\cdots+\binom{|N|}{|N|}=2^{|N|}-1$ according to the binomial theorem, resulting in a runtime complexity of $\mathcal{O}(2^{|N|})$.
-Theorem A.2 (Property of Coarsen).
+Theorem D.2 (Property of Coarsen).
 The algorithm Coarsen returns a minimal NAP specification with $\mathcal{O}(|N|)$ calls to $\mathcal{V}$.
@@
 ∎
-Appendix B Proofs of statistical approaches
+Appendix E Proofs of statistical approaches
 Our proofs of properties of Sample_Refine and Sample_Coarsen mainly follow those in (Liang et al., 2011). Interested readers may refer to it for detailed proofs.
-Theorem B.1 (Property of Sample_Refine).
+Theorem E.1 (Property of Sample_Refine).
 With probability $\theta=|(\frac{|M|}{|M|+1})|^{|M|}$, Sample_Refine has $1-\delta$ probability of outputting a minimal NAP specification with $\Theta\left(|M|^{2}(\log|N|+\log(s/\delta))\right)$ examples each iteration and $O(s|M|^{2}(\log|N|+\log(s/\delta)))$ total calls to $\mathcal{V}$.
@@
 Considering the size of the minimal specification $s$, Sample_Refine will execute $s$ iterations.
-If we sample $k$ times in each iteration, then the probability of selecting a mandatory neuron is at least $1-\frac{\delta}{s}$.
+If we sample $k$ times in each iteration, then the probability of selecting a essential neuron is at least $1-\frac{\delta}{s}$.
 Consequently, by applying a union bound, the algorithm will identify a NAP specification with a probability of at least $1-\delta$.
 Now, let’s delve deeper into one iteration of the process.
-The fundamental concept is that a mandatory neuron $m$ demonstrates a stronger correlation with proving the robustness query ($\mathcal{V}(P)=1$) compared to a non-mandatory one.
+The fundamental concept is that a essential neuron $m$ demonstrates a stronger correlation with proving the robustness query ($\mathcal{V}(P)=1$) compared to a non-essential one.
 This enhanced correlation increases the probability of its selection significantly when $k$ is sufficiently large.
-Let’s revisit the notion of $M$, representing the set of mandatory neurons with a size of $|M|$.
-Now, let’s focus on a specific mandatory neuron $n^{+}\in M$.
-We define $B_{j-}$ as the event indicating $k_{n^{-}}>k_{n^{+}}$, and $B$ as the event where $B_{j-}$ holds for any non-mandatory neuron $n^{-}\in N\setminus M$.
-Importantly, if $B$ fails to occur, then the algorithm will correctly identify a mandatory neuron.
+Let’s revisit the notion of $M$, representing the set of essential neurons with a size of $|M|$.
+Now, let’s focus on a specific essential neuron $n^{+}\in M$.
+We define $B_{j-}$ as the event indicating $k_{n^{-}}>k_{n^{+}}$, and $B$ as the event where $B_{j-}$ holds for any non-essential neuron $n^{-}\in N\setminus M$.
+Importantly, if $B$ fails to occur, then the algorithm will correctly identify a essential neuron.
 Hence, our primary objective is to establish that $Pr(B)\leq\frac{\delta}{s}$.
@@
 (20) $\frac{\delta}{s}\leq|N|e^{\frac{-k\epsilon^{2}}{2}}\text{ implies}k\geq\frac{2(log|N|+log(\frac{s}{\delta}))}{\epsilon^{2}}$
-Our attention now shifts towards deriving a lower bound for $\epsilon$, which intuitively reflects the discrepancy (in terms of correlation with proving the robustness query) between a mandatory neuron and a non-mandatory one.
+Our attention now shifts towards deriving a lower bound for $\epsilon$, which intuitively reflects the discrepancy (in terms of correlation with proving the robustness query) between a essential neuron and a non-essential one.
 It’s noteworthy that $Pr(P_{n}\in{\{\mathbf{1},\mathbf{0}}\})=\theta$ for any $n\in N$.
-Furthermore, since $n^{-}$ is non-mandatory, we can assert that $Pr(\mathcal{V}(P)=1|P_{n^{-}}\in\{\mathbf{1},\mathbf{0}\})=Pr(\mathcal{V}(P)=1)$.
+Furthermore, since $n^{-}$ is non-essential, we can assert that $Pr(\mathcal{V}(P)=1|P_{n^{-}}\in\{\mathbf{1},\mathbf{0}\})=Pr(\mathcal{V}(P)=1)$.
 By leveraging these observations, we can express:
@@
 Now, let $C_{j}={c\in C:n\in c}$ denote the clauses containing $n$.
-We reformulate $Pr(\mathcal{V}(P)=1)$ as the sum of two components: one originating from the mandatory neuron $n^{+}$ and the other from the non-mandatory neuron:
+We reformulate $Pr(\mathcal{V}(P)=1)$ as the sum of two components: one originating from the essential neuron $n^{+}$ and the other from the non-essential neuron:
 (22) $Pr(\mathcal{V}(P)=1)$ $=Pr(\mathcal{V}(P;C_{n^{+}})=1,\mathcal{V}(P;C\setminus C_{n^{+}})=0)$
@@
 The only distinction arises from conditioning on $P_{n^{+}}\in\{\mathbf{1},\mathbf{0}\}$, which introduces an extra factor of $\frac{1}{\theta}$ in the first term because conditioning divides by $Pr(P_{n^{+}})=\theta$.
-The second term remains unchanged since no $c\notin C_{n^{+}}$ is mandatory in $P_{n^{+}}$.
+The second term remains unchanged since no $c\notin C_{n^{+}}$ is essential in $P_{n^{+}}$.
 Substituting these two outcomes back into the equation yields:
 (24) $\epsilon=(1-\theta)Pr(\mathcal{V}(P;C_{n^{+}})=1,\mathcal{V}(P;C\setminus C_{n^{+}})=0)$
-Now, our objective is to establish a lower bound for (22) across all possible $\mathcal{V}$ (equivalently, $C$), where $n^{+}$ is permitted to be mandatory in $C$.
+Now, our objective is to establish a lower bound for (22) across all possible $\mathcal{V}$ (equivalently, $C$), where $n^{+}$ is permitted to be essential in $C$.
 Interestingly, the worst possible $C$ can be obtained by either having $|M|$ disjoint clauses ($C={{n}:n\in M}$) or a single clause ($C={M}$ if $s=|M|$).
-The intuition behind this is that when $C$ consists of $|M|$ clauses, there are numerous possibilities ($|M|-1$ of them) for some $c\notin C_{n^{+}}$ to be true, making it challenging to determine that $n^{+}$ is a mandatory neuron; in such cases, $\epsilon=(1-\theta)\theta(1-\theta)^{|M|-1}$.
+The intuition behind this is that when $C$ consists of $|M|$ clauses, there are numerous possibilities ($|M|-1$ of them) for some $c\notin C_{n^{+}}$ to be true, making it challenging to determine that $n^{+}$ is a essential neuron; in such cases, $\epsilon=(1-\theta)\theta(1-\theta)^{|M|-1}$.
 Conversely, if $C$ comprises a single clause, then letting this clause be true becomes exceedingly challenging; in this scenario, $\epsilon=(1-\theta)\theta^{|M|}$.
@@
 ∎
-Theorem B.2 (Property of Sample_Coarsen).
+Theorem E.2 (Property of Stochastic_Coarsen).
 With probability $\theta$=$e^{-\frac{1}{s}}$, Sample_Coarsen learns a minimal NAP specification with $\mathcal{O}(slog|N|)$ calls to $\mathcal{V}$.
@@
 By assumption, there exists a NAP $P^{S}\preccurlyeq P$ of size $s$ that passes the verification.
-Define $G(P)=\neg(P^{S}\preccurlyeq P)$, which is 0 when $P^{S}$ subsumes $P$, i.e., all mandatory neurons in $P^{S}$ shows up in the sampled $P$.
+Define $G(P)=\neg(P^{S}\preccurlyeq P)$, which is 0 when $P^{S}$ subsumes $P$, i.e., all essential neurons in $P^{S}$ shows up in the sampled $P$.
 This follows that $Pr(G(P)=0)=Pr(P^{S}\preccurlyeq P)=\theta^{s}$.